### v1

# Learning Macroeconomic Policies based on Microfoundations: A Stackelberg Mean Field Game Approach

Abstract
Effective macroeconomic policies play a crucial role in promoting economic growth and social stability. This paper models the optimal macroeconomic policy problem based on the Stackelberg Mean Field Game (SMFG), where the government acts as the leader in policy-making, and large-scale households dynamically respond as followers. This modeling method captures the asymmetric dynamic game between the government and large-scale households, and interpretably evaluates the effects of macroeconomic policies based on microfoundations, which is difficult for existing methods to achieve. We also propose a solution for SMFGs, incorporating pre-training on real data and a model-free Stackelberg mean-field reinforcement learning (SMFRL) algorithm, which operates independently of prior environmental knowledge and transitions. Our experimental results showcase the superiority of the SMFG method over other economic policies in terms of performance, efficiency-equity tradeoff, and SMFG assumption analysis. This paper significantly contributes to the domain of AI for economics by providing a powerful tool for modeling and solving optimal macroeconomic policies.
Macroeconomic policies; Stackelberg mean field game; reinforcement learning; multi-agent; social good
1 Introduction
The formulation of macroeconomic policies is crucial for the sustained development of an economy (Schneider & Frey, 1988; Persson & Tabellini, 1999). Governments can adjust economic production, wealth distribution, social stability, and welfare through economic policies such as interest rates, taxation, and fiscal spending. Thus, how to effectively model and solve optimal macroeconomic policies, and simulate the effect of policy implementation is a very important issue. Existing research methods in economic policy primarily rely on simulations based on macroeconomic models like the Dynamic Stochastic General Equilibrium (DSGE) model (An & Schorfheide, 2007) and the Solow growth model (Brock & Taylor, 2010), or on empirical analysis (Robnik-Šikonja & Kononenko, 2003) assessing policy impacts through posterior analysis of historical data. However, these methods have limitations: (1) Difficulty in modeling dynamic interactions between the government and large-scale households, including feedback from households to policy and governmental adjustments in response to this feedback. (2) Challenge in solving the behavioral strategies of large-scale micro-agents. (3) The real economic and social world is highly complex, making it challenging for macroeconomic models to satisfy assumptions such as state transition and complete information.
Addressing these challenges, this paper employs the Stackelberg Mean Field Game (SMFG) to model optimal macroeconomic policy problems. In this framework, the government, acting as the leader, initially sets policies, while a vast number of households, as followers, respond to these policies. However, considering the large scale of the followers, it is impractical for the leader to consider each follower’s dynamic response. Consequently, followers’ action-state distribution, analogous to the overall market information in economics, is introduced as a bridge between macro-level policy and micro-level individual decisions. The optimal decisions of the leader and followers depend on the followers’ overall distribution rather than on a specific agent. This assumption is crucial for modeling optimal macroeconomic policies as SMFGs and is easily comprehensible in the real world. For example, the government lowers property taxes to stimulate a sluggish housing market. This policy alters the market’s overall state, including housing prices and market demand. Households should consider both the government’s tax policy and the overall market information when making their purchasing decisions.
Furthermore, this paper proposes a solution method for the optimal macroeconomic policy issues based on SMFG: Firstly, behavior cloning of households (followers) based on real data serves as the initialization for the policy network. This pre-training enhances the stability and performance of training and prevents ineffective solutions. Subsequently, a model-free algorithm, Stackelberg Mean-Field Reinforcement Learning (SMFRL), is introduced to solve SMFGs. This method does not rely on prior knowledge of the environment or transition information, offering advantages in solving complex economic issues. We conduct experiments on TaxAI (Mi et al., 2023), an environment for studying optimal tax policies, first to validate the performance of SMFRL against other model-free algorithms such as independent learning, MADDPG, and mean field MARL in solving SMFGs. Secondly, the paper compares the SMFG method with AI Economist, the Saez tax, the 2022 U.S. Federal tax, and free market policy, demonstrating the superiority of SMFG solutions and exploring the efficiency-equity tradeoff of these economic policies. Lastly, our tests on SMFG modeling assumptions reveal that our approach remains effective even if some households’ optimal decisions do not depend on market information. In summary, the contributions of this paper are threefold:
1. This paper is the first to model optimal macroeconomic policy problems using Stackelberg mean field games, linking macro-policies (the leader) with microfoundations (large-scale followers) through the followers’ overall distribution or average actions. It overcomes limitations such as dynamic interactions and feedback, and the dimensionality curse of large-scale households.
2. This paper employs real data for pre-training and proposes a model-free Stackelberg mean field reinforcement learning algorithm to solve SMFGs, with experimental validation of its effectiveness.
3. The paper empirically showcases the superiority of modeling and solving optimal macroeconomic policy problems by SMFGs. We also explore the efficiency-equity tradeoff in various economic policies and SMFG assumptions.
Our code is shown in an anonymous GitHub repository  [^1] https://anonymous.4open.science/r/SMFG_macro_7740.
2 Related work
2.1 Economic Methods
In the field of economics, research on economic policy typically relies on theoretical model studies, empirical analysis, and econometric methods. Classic theoretical models, such as the IS-LM model (Hicks, 1980; Gali, 1992), are applicable for analyzing short-term policy effects but overlook long-term factors and price fluctuations. The aggregate demand-aggregate supply (AD-AS) model (Dutt, 2006; Lee et al., 1997) focuses on the relationship between aggregate demand and supply, integrating short-term and long-term factors, but simplifies macroeconomic dynamics. The Solow growth model (Brock & Taylor, 2010) provides a framework for economic growth but does not consider market imperfections and externalities. New Keynesian models (Blanchard & Galí, 2007; Gabaix, 2020) emphasize the stickiness of prices and wages, suitable for explaining economic fluctuations and policy interventions, but are based on strong assumptions. DSGE models (An & Schorfheide, 2007; Smets & Wouters, 2007), built on micro-foundations for macroeconomic forecasting, offer a consistent framework for policy analysis but are limited in handling nonlinearities and market imperfections. Empirical analysis (Ramesh et al., 2010; Vedung, 2017) primarily evaluates policy effects through historical data but faces limitations in data timeliness and accuracy. Econometric methods (Johnston & DiNardo, 1963; Davidson et al., 2004), such as regression and time series analysis, can quantify policy effects but are constrained by model settings and data quality.
2.2 AI for Economics
Artificial intelligence (AI), such as reinforcement learning, may offer new perspectives for solving complex economic problems (Tilbury, 2022). In terms of macroeconomics, AI Economist (Zheng et al., ) employs curriculum learning on tax policy design, Trott et al. studies the collaboration between central and local governments under COVID-19, as well as monetary policy (Hinterlang & Tänzer, ; Chen et al., 2023), international trade (Sch, 2021), market pricing (Danassis et al., 2023). However, these studies consider the government as a single agent, neglecting the dynamic response of the households to policies. Koster et al. studies democratic AI by human voting models, Yaman et al. examines the impact of social sanction rules on labor division. However, these works are based on simplified settings and remain a gap in real-world economic policy. In terms of microeconomics, Shi (2021); Rui & Shi (2022); Atashbar & Aruhan Shi (2023) study the optimal saving and consumption problems of micro-agents, and other researchers explore rational expectation equilibrium (Kuriksha, 2021; Hill et al., ), multiple equilibria under real-business-cycle (Curry et al., 2022), the emergence of barter behavior (Johanson et al., 2022), optimal asset allocation and savings strategies (Ozhamaratli & Barucca, 2022). In summary, research on macroeconomic problems still lacks a bridge framework of macro policy and microfoundations, which has drawn our attention to the Stackelberg mean field game.
2.3 Stackelberg Mean Field Game
The Stackelberg Mean Field Game models one leader and a large number of homogeneous followers, and current research can be divided into model-based methods and model-free methods. Model-based approaches typically involve complex theoretical analysis and precise mathematical modeling. For instance, Guo et al. (2022) reconceptualizes SMFGs as minimax optimization problems, while Dayanikli & Lauriere (2023); Bergault et al. (2023); Bensoussan et al. (2015); Fu & Horst (2020) initially resolve the followers’ mean field equilibrium using forward-backward stochastic functional differential equations, subsequently deriving the leader’s optimal control. Huang & Yang (2020); Bensoussan et al. (2015) explore linear-quadratic Stackelberg games. These methods provide a theoretically robust foundation but may require simplified assumptions for practical applications, particularly in complex real-world scenarios. Notably, Aurell et al. (2022) extends the mean field game model to capture epidemic dynamics, demonstrating potential applications in public health management. In contrast, model-free methods offer a way to handle complex situations that traditional numerical methods struggle with, especially in data-rich environments. Campbell et al. (2021) employs a modified deep backward stochastic differential equation (BSDE) method for solving the followers’ equilibrium, yet the leader’s choices are limited. In conclusion, we plan to use SMFGs to model macroeconomic policy issues and propose a model-free algorithm to solve this complex task.
Figure 1: The mapping from optimal macroeconomic policy problems (left) to the Stackelberg Mean Field Games (right). The left represents the reality where the macro-level government implements fiscal or monetary policies to optimize GDP, welfare, or equity, and the micro-level consists of large-scale households engaging in economic activities and forming the market. In SMFGs, the government corresponds to the leader, households to the followers, and market information to the followers’ state-action distribution.
3 Macroeconomic Policy Formulation
3.1 Economic Modeling
In reality, an economy comprises at least one government and large-scale households. The diverse economic activities of households, such as production, trade, and consumption, aggregate to form the market. The government can maintain economic growth and social stability (Figure 1 left) via economic policies. The following is a mathematical model for optimizing tax policy:
Households
At timestep $t$, a household holds assets $a_{t}$. Its income $i_{t}$ consists of labor income $W_{t}e_{t}h_{t}$, and capital income $r_{t}a_{t}$. Here, $W_{t}$ and $r_{t}$ are the wage and interest rates determined by the labor and capital markets, respectively, while $e_{t}$ indicates the private education level. Households are obligated to pay taxes as per the government’s tax policy $T_{t}(i_{t},a_{t})$. Each household decides on its consumption $c_{t}\in R^{+}$, working hours $h_{t}\in R^{+}$, and next savings $a_{t+1}\in R^{+}$ to maximize long-term utility until $T_{max}$, subject to the budget constraint:
$\max_{a_{t+1},h_{t}}$ $\quad\mathbb{E}_{0}\sum_{t=0}^{T_{max}}u(c_{t},h_{t})$
[…]
Government
The government aims to foster economic growth, maintain social equity, and enhance societal welfare through three fiscal tools: debt $B_{t}\in R^{+}$, taxation policy $T_{t}(i_{t},a_{t}):R^{+}\times R^{+}\rightarrow R^{+}$, and spending $G_{t}\in R^{+}$.
$\max_{B_{t},T_{t},G_{t}}\quad\mathbb{E}_{0}\sum_{t=0}^{T_{max}}\left[\alpha_{1}Y_{t}+\alpha_{2}S(\mathcal{I}_{t},\mathcal{W}_{t})+\alpha_{3}\sum_{i}^{N}u_{t}^{i}\right],$
where $\alpha_{1},\alpha_{2},\alpha_{3}$ are weight parameters reflecting the government’s preference for different objectives. Economic output (e.g., GDP) can be modeled using the Cobb–Douglas production function $Y_{t}=F(K_{t},\mathcal{L}_{t})$, with production capital $K_{t}$ and labor $\mathcal{L}_{t}$. $S(\mathcal{I}_{t},\mathcal{W}_{t})$ denotes the social equity function, where $\mathcal{I}_{t},\mathcal{W}_{t}$ are the income and wealth distributions among households. The sum of households’ utilities $\sum_{i}^{N}u_{t}^{i}$ represents the overall societal welfare.
3.2 Stackelberg Mean Field Game Modeling
In the aforementioned issue, the government (as the leader) formulates policies to guide the decisions of households (as followers), thereby collectively achieving macroeconomic objectives, such as economic growth. Households provide feedback to government policies, creating an asymmetric game with leader-follower dynamics. However, the dynamic interactions among large-scale households extend beyond the standard Stackelberg games. Therefore, this study proposes modeling the problem as a Stackelberg Mean Field Game (Figure 1), based on the following assumptions:
  1. 1.
    Stackelberg assumption: the government (leader) sets policies first, and households (followers) adjust their behavior based on these policies. Both the leader and followers dynamically optimize their strategies, gradually reaching equilibrium.
  2. 2.
    Mean field assumption: households (followers) are assumed to be homogeneous agents. The optimal policies of the leader and followers depend on the overall distribution or average behavior of the followers, rather than a specific one. For example, both governments and households should consider overall market information when making decisions in reality.
Following this, the optimal macroeconomic policy issues can be rewritten as the following game proceeding: For any given time $t\in\{0,\ldots,T_{max}\}$, the leader selects an action $a_{t}^{l}\in\mathcal{A}^{l}$ based on the leader policy $\pi_{t}^{l}(\cdot|s_{t}^{l})$, given the state $s_{t}^{l}$. Subsequently, the representative follower, located in state $s_{t}^{f}\in\mathcal{S}^{f}$, chooses an action $a_{t}^{f}\in\mathcal{A}^{f}$ according to the policy $\pi_{t}^{f}(\cdot|s_{t}^{f},a^{l}_{t})$. The sequences $\{\pi_{t}^{l}\}_{t=0}^{T-1}$ and $\{\pi_{t}^{f}\}_{t=0}^{T-1}$ are denoted as $\pi^{l}$ and $\pi^{f}$, respectively.
Followers’ Mean Field Game
Given the leader’s policy $\pi^{l}$, the SMFG is simplified into a mean field game for the followers. At time $t$, given the leader’s action $a^{l}_{t}$, the representative follower’s state $s_{t}^{f}$ and action $a_{t}^{f}$ determine the followers’ state-action distribution as $L^{f}_{t}(s_{t}^{f},a_{t}^{f})=\mathbb{P}(s_{t}^{f},a_{t}^{f})$. The representative follower then receives a reward $r^{f}(s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$ and transitions to the next state $s_{t+1}^{f}\sim P(\cdot|s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$. Each follower aims to find the optimal policy $\pi^{f}$ that maximizes his cumulative reward over the time horizon:
$J^{f}(\pi^{l},\pi^{f},L^{f})$ $=\mathbb{E}^{f}\left[\sum_{t=1}^{T}r^{f}(s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})\right]$
$=\mathbb{E}^{f}\left[\sum_{t=1}^{T}u(c_{t},h_{t})\right],$
where $\text{where }c_{t}$ $=W_{t}e_{t}h_{t}+r_{t}a_{t}-T_{t}(i_{t},a_{t})-a_{t+1},$
where $\mathbb{E}^{f}$ denotes the expectation with respect to initial state $s_{0}^{f}\sim\mu_{0}^{f}$, $s_{t+1}^{f}\sim P(\cdot\mid s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$ and $a^{f}_{t}\sim\pi^{f}(\cdot\mid s_{t}^{f},a^{l}_{t})$ for each $t$. For households in economic problems, the follower’s state $s^{f}_{t}$ incorporates asset $a_{t}$ and education level $e_{t}$, and the follower’s actions $a^{f}_{t}$ include next savings $a_{t+1}$ and working hours $h_{t}$. The leader’s action $a^{l}_{t}$ denotes tax function $T_{t}(i_{t},a_{t})$, and followers’ distribution $L^{f}_{t}$ includes wage rate $W_{t}$ and interest rate $r_{t}$. Given the leader’s policy $\pi^{l}$, the best response operator of followers is defined as
${\pi^{f}}^{*}\in\mathrm{BR}^{f}(\pi^{l},L^{f})=\arg\max_{\pi^{\prime}}J^{f}(\pi^{l},\pi^{\prime},L^{f}).$
Stackelberg Leader’s Game
At time $t$, given the leader’s state $s^{l}_{t}$, action $a^{l}_{t}$, and the followers’ state-action distribution $L_{t}^{f}$, the leader will receive a reward $r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})$, and transition to a new state $s_{t+1}^{l}$ according to the transition probabilities $P(\cdot|s_{t}^{l},a_{t}^{l},L^{f}_{t})$. The leader aims to find the best policy $\pi^{l}$ to maximize the total expected rewards:
$J^{l}(\pi^{l},\pi^{f},L^{f})=\mathbb{E}^{l}[\sum_{t=1}^{T}r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})],$
$r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})=\alpha_{1}Y_{t}+\alpha_{2}S(\mathcal{I}_{t},\mathcal{W}_{t})+\alpha_{3}\sum_{i}^{N}u_{t}^{i}$
$=\alpha_{1}F(K_{t},\mathcal{L}_{t})+\alpha_{2}S(\mathcal{I}_{t},\mathcal{W}_{t})+\alpha_{3}U(T_{t},W_{t},r_{t})$
where $\mathbb{E}^{l}$ denotes the expectation under this probability distribution for the leader when $s_{0}^{l}\sim\mu_{0}^{l},s_{t+1}^{l}\sim P(\cdot|s_{t}^{l},a_{t}^{l},L^{f}_{t})$ and $a_{t}^{l}\sim\pi_{t}^{l}(\cdot\mid s_{t}^{l})$ for $t\in\{0,...,T-1\}$. For the government in economic problems, the leader’s state $s^{l}_{t}$ is characterized by production capital $K_{t}$, and the leader’s actions $a^{l}_{t}$ encompass government spending $G_{t}$ and a taxation policy $T_{t}(\cdot)$. The followers’ state-action distribution $L^{f}_{t}$ includes aggregate labor $\mathcal{L}_{t}$, income distribution $\mathcal{I}_{t}$, wealth distribution $\mathcal{W}_{t}$, wage rate $W_{t}$, and interest rate $r_{t}$, these all belong to market information. $U(\cdot)$ represents a measurable function for $T_{t},W_{t},r_{t}$. Therefore, we can define the leader’s best response operator given $\pi^{f},L^{f}$:
${\pi^{l}}^{*}\in\text{BR}^{l}(\pi^{f},L^{f}):=\mathop{\arg\max}_{\pi^{l}}J^{l}(\pi^{l},\pi^{f},L^{f}).$
We then define the transition operator $\mathcal{F}$ as $L^{f}=\mathcal{F}(\pi^{l},\pi^{f},L^{f}_{0},\mu_{0})$ if the evolution of population distribution $L$ satisfies the Mckean-Vlasov equation (1).
$L^{f}_{t+1}(s_{t+1}^{f},a_{t+1}^{f})=$ $\sum_{s_{t}^{l},a_{t}^{l},s_{t}^{f},a_{t}^{f}}L^{f}_{t}(s_{t}^{f},a_{t}^{f})\pi_{t}^{l}(a_{t}^{l}|s_{t}^{l})\mu_{t}^{l}(s_{t}^{l})$ (1)
$\quad P(s_{t+1}^{f}|s_{t}^{f},a_{t}^{f},a_{t}^{l})\pi_{t+1}^{f}(a_{t+1}^{f}|s_{t+1}^{f}).$
$\mu_{t+1}^{l}(s_{t+1}^{l})=$ $\sum_{s_{t}^{l},a_{t}^{l}}\mu_{t}^{l}(s_{t}^{l})\pi_{t}^{l}(a_{t}^{l}|s_{t}^{l})P(s_{t+1}^{l}|s_{t}^{l},a_{t}^{l}).$
Definition 3.1.
The Stackelberg mean field equilibrium is a tuple $({\pi^{l}}^{*},{\pi^{f}}^{*},{L^{f}}^{*})$, which satisfies the following conditions:
${\pi^{l}}^{*}\in\mathrm{BR}^{l}({\pi^{f}}^{*},{L^{f}}^{*}),{\pi^{f}}^{*}\in\mathrm{BR}^{f}({\pi^{l}}^{*},{L^{f}}^{*}),$
${L^{f}}^{*}=\mathcal{F}({\pi^{l}}^{*},{\pi^{f}}^{*},L^{f}_{0},\mu_{0}).$
4 Methods
Solving the optimal macroeconomic policy is very complex in the real world. Reliance solely on mathematical models for simulation and problem-solving inevitably creates a gap with actual scenarios, thus incorporating real historical data is essential. Nonetheless, a singular focus on historical data for the empirical analysis of economic policies inadequately captures the complex and asymmetric dynamics of interactions between the government and households. Therefore, our solution method is divided into two steps: firstly, we pre-train the follower’s policy based on real data using behavior cloning (BC). Secondly, this paper proposes a model-free Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm to solve the Stackelberg Mean Field Games (SMFGs) between the government and large-scale households.
4.1 Pretain by Behavior Cloning
Based on real historical data, we fetch a large number of followers’ state-action pairs $\{s^{f},a^{f}\}$ from a real-data buffer $\mathcal{D}_{real}$ for behavior cloning. For different settings of network structures, we have chosen two types of loss: when the neural network outputs a probability distribution of actions, we use the negative log-likelihood loss (NLL loss); when the neural network outputs action values, we employ the mean square error loss (MSE loss). Our goal is to find the optimal parameters $\theta$ as the follower’s policy network $\pi_{\theta}$ initialization, thereby minimizing the loss to its lowest convergence.
$\min_{\theta}\mathcal{L}_{NLL}$ $=-\mathbb{E}_{s^{f},a^{f}\sim\mathcal{D}}\log{\pi_{\theta}(a^{f}\mid s^{f})},$
$\min_{\theta}\mathcal{L}_{MSE}$ $=\mathbb{E}_{s^{f},a^{f}\sim\mathcal{D}}\left(a^{f}-a\right)^{2}|_{a=\pi_{\theta}(s^{f})}.$
4.2 Stackelberg Mean Field Reinforcement Learning
To solve SMFGs in sequential decision-making, we design a Stackelberg mean field reinforcement learning algorithm based on centralized training with decentralized execution. Consider a game with a leader and $N$ homogeneous follower agents. The leader makes decisions first, and all followers make simultaneous decisions after observing the leader’s actions. The leader aims to find a policy $\pi^{l}:\mathcal{S}^{l}\rightarrow\mathcal{A}^{l}$ to optimize his long-term objective, which is often challenging and requires the cooperation of all followers. Each follower agent seeks to find a policy $\pi^{f}:\mathcal{S}^{f}\times\mathcal{A}^{l}\rightarrow\mathcal{A}^{f}$ to optimize long-term utility after observing the current leader’s action $a^{l}_{t}=\pi^{l}(s^{l}_{t})$. The joint action of the leader and follower agents $\mathbf{a}_{t}=\{a^{l}_{t},a^{f1}_{t},...,a^{fN}_{t}\},a^{fi}_{t}=\pi^{f}(s^{fi}_{t})$ impacts the environment and results in a reward $\mathbf{r}_{t}=\{r^{l}_{t},r^{f1}_{t},...,r^{fN}_{t}\}$ and the next observation $\mathbf{s}_{t+1}=\{s^{l}_{t+1},s^{f1}_{t+1},...,s^{fN}_{t+1}\}$. Here the experience replay buffer $\mathcal{D}$ contains the tuples $(\mathbf{s}_{t},\mathbf{s}_{t+1},\mathbf{a}_{t},\mathbf{r}_{t},L^{f}_{t})$, recording experiences of all agents. $Q_{i}(\mathbf{s}_{t},\mathbf{a}_{t})$ is a standard centralized action-value function that takes as input the joint state $\mathbf{s}_{t}$ and joint action $\mathbf{a}_{t}$ and outputs the Q-value for agent $i$. However, due to numerous follower agents, the dimensions of both the joint state and joint action increase significantly with $N$, making the Q-function infeasible to learn. We equate the centralized action-value function $Q_{i}(\mathbf{s}_{t},\mathbf{a}_{t})$ to $\tilde{Q}_{t}(s_{t}^{l},a_{t}^{l},L_{t}^{f})$ based on mean field theory. The specific theoretical description is shown in the Appendix D.
4.3 Implementation
Furthermore, we consider deterministic policies for both a leader and its followers in a multi-agent system. Specifically, we define the leader’s policy as $\pi_{\theta^{l}}$ parameterized by $\theta^{l}$ (abbreviated as $\pi^{l}$), and a shared policy for the followers denoted as $\pi_{\theta^{f}}$ with parameters $\theta^{f}$ (abbreviated as $\pi^{f}$).
For the leader agent
The policy network $\pi_{\theta^{l}}$ of the leader agent, i.e. the actor, is trained by the sampled policy gradient (Silver et al., 2014):
$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{-},L^{f}_{t})|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}\right]$
In the subsequent step, we employ a neural network with parameters $\phi^{l}$ to approximate the leader agent’s action-value function, denoted as $\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$ or $\tilde{Q}^{l}$ for brevity. It is updated by minimizing a loss function:
$\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)^{2}\right],$
$y^{l}_{t}=r_{t}^{l}+\gamma\tilde{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L^{f}_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})},$ (2)
where $y^{l}_{t}$ represents the target Q value, computed using the parameters $\phi^{l}_{-}$, the next population distribution $L^{f}_{t+1}$ is based on the Mckean-Vlasov equation (1), and $\gamma$ is discount factor. Differentiating the loss function $\mathcal{L}(\phi^{l})$ yields the gradient utilized for training:
$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=$
$\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{l}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right]$
For the representative follower agent
The policy network $\pi_{\theta^{f}}$ is trained by the policy gradient:
$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L^{f}_{t}\sim\mathcal{D}}\Big{[}$
$\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{-},a^{l}_{t},L^{f})|_{a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})}\Big{]},$
Similar to Mean Field Approximation (Yang et al., 2018), the standard centralized action-value function $Q(\mathbf{s}_{t},\mathbf{a}_{t})$ is approximated as $\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})$ with parameters $\phi^{f}$, which is updated by minimizing the MSE loss function:
$\mathcal{L}(\phi^{f})=$ $\mathbb{E}_{{s}_{t}^{f},{a}_{t}^{f},a^{l}_{t},L^{f}_{t},s_{t+1}^{f}\sim\mathcal{D}}\left(y^{f}_{t}-\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)^{2}$
$y^{f}_{t}=$ $r^{f}_{t}+\gamma\tilde{Q}_{\phi^{f}_{-}}(s^{f}_{t+1},a^{f}_{t+1},a^{l}_{t+1},L^{f}_{t+1}),$
where the next leader’s action $a^{l}_{t+1}=\mu_{\phi^{l}_{-}}(s^{l}_{t+1})$, the next population distribution $L^{f}_{t+1}=L(\mathbf{s}_{t+1}^{f},\mathbf{a}_{t+1}^{f})$, the next joint followers’ action $\mathbf{a}_{t+1}^{f}=\{a^{f1}_{t+1},...,a^{fN}_{t+1}\}$, $a^{fi}_{t+1}=\mu(s^{fi}_{t+1},a^{l}_{t+1})$. Differentiating the loss function $\mathcal{L}(\phi^{f})$ yields the gradient utilized for training:
$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\mathcal{D}}\Big{[}\left(y^{f}_{t}-\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)$
$\nabla_{\phi^{f}}\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\Big{]}$
The full algorithm (1) is described in the Appendix.
5 Experiment
Figure 2: The training curves for 6 algorithms on 4 macroeconomic indicators, comparing settings without behavior cloning as pre-train (N=100 without BC) and with behavior cloning (N=100-BC & N=1000-BC).
Table 1: Test results for 6 algorithms on 5 indicators under 3 settings (N=100 without BC / N=100-BC / N=1000-BC).
Algorithms Per Capita GDP Social Welfare Income Gini Wealth Gini Years
Rule-based 1.41e+05 / 3.66e+11 / 1.41e+05 69.27 / 79.23 / 334.79 0.89 / 0.52 / 0.90 0.92 / 0.53 / 0.93 1.00 / 217.45 / 1.00
DDPG 1.41e+05 / 2.03e+12 / 4.92e+11 70.91 / 94.50 / 527.09 0.88 / 0.46 / 0.75 0.92 / 0.48 / 0.79 1.00 / 299.85 / 100.68
MADDPG 4.93e+04 / 6.38e+12 / 6.82e+12 55.74 / 93.89 / 954.88 0.91 / 0.57 / 0.56 0.92 / 0.58 / 0.62 1.00 / 268.53 / 278.50
IDDPG 1.21e+05 / 7.41e+12 / 2.79e+12 83.09 / 98.16 / 512.19 0.88 / 0.53 / 0.77 0.92 / 0.55 / 0.81 1.00 / 300.00 / 100.68
MF-MARL 8.66e+07 / 5.44e+12 / 1.13e+05 82.02 / 98.21/ 440.00 0.82 / 0.50 / 0.90 0.83 / 0.52 / 0.93 75.75 / 300.00 / 1.00
SMFRL 9.59e+12 / 1.01e+13 / 9.68e+12 96.87 / 96.90 / 975.15 0.52 / 0.51 / 0.52 0.51 / 0.53 / 0.51 300.00 / 299.89 / 300.00
In this section, we conduct two important experiments in the TaxAI environment: (1) We compare the performance of the SMFRL algorithm with 5 baselines in solving SMFGs; (2) We compare the effects of the SMFG method with 5 economic policies, and discuss the efficiency-equity tradeoff and SMFG assumptions.
5.1 Environment
TaxAI (Mi et al., 2023) is a reinforcement learning simulator used for studying optimal tax policy. The environmental settings include the government promoting economic growth (i.e., GDP) through adjusting tax policy and government spending, and households optimizing labor supply and savings ratio for long-term personal utility. It is one of the few simulators for researching optimal macroeconomic policies, hence we conduct experiments on it.
5.2 Algorithmic Analysis for SMFGs
Baselines
We compare popular reinforcement learning algorithms including DDPG (Lillicrap et al., 2015) (only for leader), IDDPG (de Witt et al., 2020), MADDPG (Lowe et al., 2017), Mean Field MARL (Yang et al., 2018), and rule-based policy with our algorithm SMFRL. For more details see Appendix B.1 and Table 2.
Results
We compared SMFRL with 5 baselines in three different experimental setups: without behavior cloning as pre-training and with 100 followers (marked as N=100 without BC); with BC-based pre-training for follower agents at N=100 and N=1000 (N=100-BC; N=1000-BC). Figure 2 illustrates the training curves of 4 key macroeconomic indicators under these three settings. Each row corresponds to one setting, and each column to a macroeconomic indicator, including per capita GDP, social welfare, income Gini, and wealth Gini. A rise in per capita GDP indicates economic growth, an increase in social welfare implies happier households and a lower Gini index indicates a fairer society. Each subplot’s Y-axis represents the indicators’ values, and the X-axis represents the training steps. Table 1 displays the test results of the 6 algorithms across 5 indicators, with each column corresponding to an experimental setting.
Figure 2 and Table 1 present two experimental findings: (1) Using BC as a pre-training method for the follower’s policy enhances the algorithms’ stability and performance. Comparing settings with and without BC (the first two rows), our method, SMFRL, shows similar convergence outcomes; however, the performance of other algorithms significantly improves across all four indicators with BC-based pre-training. Furthermore, the training curves of each algorithm are more stable. (2) The SMFRL algorithm substantially outperforms other algorithms in solving SMFGs, both in large-scale followers and without pre-training scenarios. In the setting of N=100-BC, SMFRL achieved a significantly higher per capita GDP compared to other algorithms, while its social welfare and Gini index are similar to others, essentially reaching the upper limit. Besides, in N=100 without BC and N=1000-BC, SMFRL consistently obtains the most optimal solutions across all indicators.
Figure 3: The training curves for 6 tax policies on 5 macroeconomic indicators (N=100 & N=1000 with BC)
5.3 Comparison of Macroeconomic Policies
Baselines
We compare free market policy, Saez tax (Saez, 2001), U.S. Federal tax  [^2] https://data.oecd.org/, AI Economist (Zheng et al., ), and AI Economist-BC. Saez tax will provide suggestions for real-world tax reform. For more details see Appendix B.1.
Policies Results
We compare the performance of 6 policies across four economic indicators under two settings: with N=100 and N=1000 households. Figure 3 displays the training curves and Table 3 shows the test results in Appendix. Both Figure 3 and Table 3 indicate that the SMFG method significantly surpasses other policies in the task of optimizing GDP, and achieves the highest social welfare. When N=100, the Saez tax achieves the lowest income and wealth Gini coefficients, suggesting greater fairness. However, at N=1000, SMFG performs optimally across all economic indicators, while the effectiveness of other policies noticeably diminishes as the number of households increases. The Saez tax also reduces the Gini index, but not as effectively or stably as the SMFG.
Figure 4: Comparative performance of various policies under multi-objective assessment (Efficiency-Equity).
Figure 5: Comparison of micro and macroeconomic indicators at varying proportions of SMFG households, demonstrating the positive impact on the individual and overall economy of SMFG households.
Efficiency-Equity Tradeoff
In economics, the Efficiency-Equity Tradeoff is a highly debated issue. We find that our SMFG method is optimal in balancing efficiency-equity, except in cases of extreme concern for social fairness. In our study, we depict the economic efficiency (Per capita GDP) on the Y-axis and equity (wealth Gini) on the X-axis of Figure 3(a) for various policies. Different policies are represented by circles of different colors, with their sizes proportional to social welfare. Different circles of the same color correspond to different seeds. Figure 3 (a) shows that the wealth Gini indices for SMFG and AI Economist-BC are similar, but SMFG has a higher GDP, suggesting its superiority over AI Economist-BC. SMFG significantly outperforms the free market policy and AI Economist due to its higher GDPs and lower wealth Ginis. However, comparing SMFG with the Saez tax and the U.S. Federal tax policy in terms of both economic efficiency (GDP) and social equity (Gini) is challenging. Therefore, we introduce Figure 3 (b) to demonstrate the performance of different policies under various weights in a multi-objective assessment.
In Figure 3 (b), the Y-axis shows the weighted values of the multi-objective function $Y=log(\text{per capita GDP})+\alpha(\text{wealth Gini})$, and the X-axis represents the weight of the wealth Gini index. For each weight $\alpha$, we compute the multi-objective weighted values for those policies, represented as circles of different colors. Due to the logarithmic treatment of GDP in (b), when $\alpha=10$, the overall objective focuses solely on social fairness; when $\alpha=0$, the overall objective is concerned only with efficiency. Our findings in (b) reveal that only when $\alpha\geq 8$, which indicates a substantial emphasis on social equity, does the Saez tax outperform SMFG. However, SMFG consistently proves to be the most effective policy under a wide range of preference settings.
SMFG Assumption Analysis
The above experiments have already demonstrated the superiority of the SMFG method. However, the modeling and solving of SMFG rely on an important assumption: the optimal decisions of governments and households depend on the overall state information. In reality, it is natural for the government to focus on market information in decision-making, but it cannot be guaranteed that all households will consider market information in their decisions. To discuss the applicability of the SMFG method, we introduce two types of households within the same economy: those based on real data (Real-data Households) and those that focus on market information (SMFG households). We simulate the economic operations at various proportions of SMFG households: 0%, 25%, 50%, 75%, and 100%. Figure 5 displays the microeconomic indicators and macroeconomic GDP, with each subplot illustrating the average values of different economic indicators for different households (bars in left Y-axis) and their aggregate values (dots in right Y-axis). GDP, being a macroeconomic indicator, did not have corresponding values plotted for households.
The results indicate that, across all proportions, the microeconomic indicators of SMFG households, such as utility, wealth, and income, are consistently higher than Real-data households. As the proportion of SMFG households increases, there is an upward trend in economic indicators. This suggests that the presence of SMFG households positively impacts both individual and overall economic development, and our SMFG method also works when some households do not meet the assumption of focusing on market information. Therefore, governments should encourage households to pay more attention to market information in decision-making.
6 Conclusion
This paper innovatively models the optimal macroeconomic policy problem based on the Stackelberg mean field game, and proposes a solution method involving pre-training based on real data and a model-free Stackelberg mean field reinforcement learning algorithm. Our method addresses the limitations of existing approaches, including dynamic interactions and feedback between governments and households, interactions among large populations of households, and solving complex tasks without relying on environmental transitions. Experimental results have showcased the performance of the SMFG method. In conclusion, this paper contributes an effective approach to modeling and solving optimal macroeconomic policies in the field of AI for economics and AI for social good.
Impact Statements
This paper presents work whose goal is to advance the fields of AI for social good and AI for economics. Our work aims to offer suggestions and references for governments and the people, yet it must not be rashly applied to the real world. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Economic Model Details
Economic activities among households aggregate into labor markets, capital markets, goods markets, etc. In the labor market, households are the providers of labor, with the aggregate supply $S(W_{t})=\sum_{i}^{N}e_{t}^{i}h_{t}^{i}$, and firms are the demanders of labor, with the aggregate demand $D(W_{t})=\mathcal{L}_{t}$. When supply equals demand in the labor market, there exists an equilibrium price $W_{t}^{*}$ that satisfies:
$S(W^{*}_{t})=D(W^{*}_{t}),\mathcal{L}_{t}=\sum^{N}_{i}e_{t}^{i}h_{t}^{i}.$
In the capital market, financial intermediaries play a crucial role, lending the total deposits of households $A_{t+1}=\sum_{i}^{N}a_{t+1}$ to firms as production capital $K_{t+1}$, and purchasing government bonds $B_{t+1}$ at the interest rate $r_{t}$. The capital market clears when supply equals demand:
$K_{t+1}+B_{t+1}-A_{t+1}=\left(r_{t}+1\right)(K_{t}+B_{t}-A_{t})$
In the goods market, firms produce and supply goods, while all households, the government, and physical capital investments $X_{t}$ demand them. The goods market clears when:
$Y_{t}=C_{t}+G_{t}+X_{t}$
where $C_{t}=\sum_{i}^{N}c^{i}_{t}$ represents the total consumption of consumers, and $G_{t}$ is government spending. The supply, demand, and price represent the states of the market.
Algorithm 1 Stackelberg Mean Field Reinforcement Learning (SMFRL)
  Initialize $\tilde{Q}_{\phi^{l}},\tilde{Q}_{\phi^{l}_{-}},\tilde{Q}_{\phi^{f}},\tilde{Q}_{\phi^{f}_{-}},\pi_{\theta^{l}},\pi_{\theta^{l}_{-}},\pi_{\theta^{f}},\pi_{\theta^{f}_{-}}$.
  for epoch = 1 to M do
     Recieve initial observation $\mathbf{s}_{t}=\{s^{l}_{t},s^{f1}_{t},...,s^{fN}_{t}\}$.
     for  $t=1$ to max-epoch-length do
        For leader agent, select action $a^{l}_{t}=\pi_{\theta^{l}}(s^{l}_{t})+\mathcal{N}_{t}$; for each follower agent $i$, select action $a^{fi}=\pi_{\theta^{f}}(s^{fi}_{t},a^{l}_{t})+\mathcal{N}_{t}$ w.r.t. the current policy and exploration.
        Execute actions $\mathbf{a}_{t}=\{a^{l}_{t},a^{f1}_{t},...,a^{fN}_{t}\}$ and observe reward $\mathbf{r}_{t}=\{r_{t}^{l},r^{f1}_{t},...,r^{fN}_{t}\}$ and next observation $\mathbf{s}_{t+1}$.
        Store $(\mathbf{s}_{t},\mathbf{s}_{t+1},\mathbf{a}_{t},\mathbf{r}_{t},L^{f}_{t})$ in the replay buffer $\mathcal{D}$.
        $\mathbf{s}_{t}\leftarrow\mathbf{s}_{t+1}$.
        for j =1 to update-cycles do
           Sample a random minibatch of $N_{b}$ samples from $\mathcal{D}$.
           For the leader agent, set
$y^{l}_{t}=r_{t}^{l}+\gamma\tilde{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L^{f}_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})}.$
           Update leader’s critic net by minimizing the loss:
$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{l}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right]$
           Update leader’s actor by sampled policy gradient:
$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{-},L^{f}_{t})|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}\right].$
           For the representative follower agent, set
$y^{f}_{t}=r^{f}_{t}+\gamma\tilde{Q}_{\phi^{f}_{-}}(s^{f}_{t+1},a^{f}_{t+1},a^{l}_{t+1},L^{f}_{t+1}).$
           Update follower’s critic by minimizing the loss:
$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\mathcal{D}}\Big{[}\left(y^{f}_{t}-\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{f}}\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\Big{]}.$
           Update follower’s actor by policy gradient:
$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L^{f}_{t}\sim\mathcal{D}}\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{-},a^{l}_{t},L^{f})|_{a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})}\Big{]}.$
        end for
        Update target network parameters for each agent with learning rates $\tau_{\phi}$ and $\tau_{\theta}$:
$\begin{gathered}\phi_{-}^{l}\leftarrow\tau_{\phi}\phi^{l}+\left(1-\tau_{\phi}\right)\phi_{-}^{l}\\ \theta_{-}^{l}\leftarrow\tau_{\theta}\theta^{l}+\left(1-\tau_{\theta}\right)\theta_{-}^{l}\\ \phi_{-}^{f}\leftarrow\tau_{\phi}\phi^{f}+\left(1-\tau_{\phi}\right)\phi_{-}^{f}\\ \theta_{-}^{f}\leftarrow\tau_{\theta}\theta^{f}+\left(1-\tau_{\theta}\right)\theta_{-}^{f}\end{gathered}$
     end for
  end for
Appendix B Additional Results
B.1 Baselines
Baselines for solving SMFGs
Current research on model-based methods for SMFG is limited by strong assumptions and simplifications, rendering them ineffective for complex problem-solving. However, there is a lack of model-free algorithms for solving SMFG in continuous decision spaces. Therefore, this study opts to compare with popular reinforcement learning algorithms, including DDPG, MADDPG, IDDPG, and mean field MARL. Table 2 provides detailed information on the algorithms used by the leader and follower agent in these baselines.
Table 2: The details of the algorithms used by the leader and follower agents in the baselines for solving SMFG.
Baselines Follower’s Algorithm Leader’s Algorithm
Rule-based Random/Behavior cloning Rule-based/Free market
DDPG Random/Behavior cloning DDPG
MADDPG MADDPG MADDPG
IDDPG IDDPG IDDPG
MF-MARL Mean Field MARL DDPG
SMFRL SMFRL SMFRL
Baselines for economic policies
We compare 5 economic policies with our method SMFG:
(1) Free Market: A market without policy intervention.
(2) Saez Tax (Saez, 2001): The Saex tax policy is often considered a suggestion for specific tax reforms in the real world. Details are shown in Appendix C.
(3) U.S. Federal Tax: Real data from OECD  [^3] https://data.oecd.org/ in 2022 for this policy.
(4) AI Economist (Zheng et al., ): This is a two-level MARL method based on Proximal Policy Optimization (PPO). In the first phase, households’ policies are trained from scratch in a free-market (no-tax) environment. In the second phase, households continue to learn under an RL social planner.
(5) AI Economist-BC: For fairness in comparison, we evaluated the AI Economist method with behavior cloning as pre-training to determine its effectiveness.
B.2 Efficiency-Equity Tradeoff
In Figure 6, we also present the efficiency (per capita GDP) and equity (wealth Gini) corresponding to various algorithms for solving SMFGs. These methods result in similar Gini indices, with SMFRL achieving the highest per capita GDP. Apart from a strong emphasis on societal fairness, SMFRL emerges as the algorithm that can best strike a balance between efficiency and equity.
Figure 6: Comparison of macroeconomic indicators under different algorithms for solving SMFGs.
Table 3: Test results for 6 tax policies on 5 macroeconomic indicators ($N=100$ $N=1000$).
Algorithms $N=100$ $N=1000$
Per Capita Social Income Wealth Years Per Capita Social Income Wealth Years
GDP Welfare Gini Gini GDP Welfare Gini Gini
Free Market 1.37e+05 32.97 0.89 0.92 1.10 1.41e+05 334.79 0.90 0.93 1.00
Saez Tax 2.34e+12 73.82 0.21 0.38 300.00 6.35e+11 498.88 0.68 0.73 100.578
US Federal Tax 4.88e+11 94.19 0.40 0.40 289.55 1.41e+05 351.17 0.89 0.93 1.00
AI Economist-BC 4.24e+12 97.24 0.54 0.52 299.55 N/A N/A N/A N/A N/A
AI Economist 1.26e+05 72.81 0.88 0.91 1.00 N/A N/A N/A N/A N/A
SMFG 1.01e+13 96.90 0.51 0.53 299.89 9.68e+12 975.15 0.52 0.51 300.00
B.3 Behavior Cloning Experiments
This experiment conducts behavior cloning on networks for four different household policies: Multilayer Perceptron (MLP), AI economist’s network (MLP+LSTM+MLP), MF-MARL, and MADDPG network. The first two, as their network outputs are probability distributions, use negative log-likelihood loss (Figure 7 left); the latter two’s networks employ deterministic policies, hence they use mean square error loss against real data (Figure 7 right). The loss convergence curve of behavior cloning is shown in Figure 7. It can be observed that the AI economist’s network, due to its complexity, struggles to converge to near -1 like MLP. The losses corresponding to MFRL and MADDPG can converge to below 0.1.
Figure 7: The behavior cloning loss for 4 networks in two loss types.
Appendix C Saez tax
The Saex tax policy is often considered a suggestion for specific tax reforms in the real world. The specific calculation method is as follows (Saez, 2001). The Saez tax utilizes income distribution $f(z)$ and cumulative distribution $F(z)$ to get the tax rates. The marginal tax rates denoted as $\tau(z)$, are expressed as a function of pretax income $z$, incorporating elements such as the income-dependent social welfare weight $G(z)$ and the local Pareto parameter $\alpha(z)$.
$\tau(z)=\frac{1-G(z)}{1-G(z)+\alpha(z)e(z)}$
To further elaborate, the marginal average income at a given income level $z$, normalized by the fraction of incomes above $z$, is denoted as $\alpha(z)$.
$\alpha(z)=\frac{zf(z)}{1-F(z)}$
The reverse cumulative Pareto weight over incomes above $z$ is represented by $G(z)$.
$G(z)=\frac{1}{1-F(z)}\int_{z^{\prime}=z}^{\infty}p\left(z^{\prime}\right)g\left(z^{\prime}\right)\mathrm{d}z^{\prime}$
From the above calculation formula, we can calculate $G(z)$ and $\alpha(z)$ by income distribution. We obtain the data of income and marginal tax rate through the interaction between the agent and environment and store them in the buffer. It is worth noting that the amount of buffer is fixed.
To simplify the environment, we discretize the continuous income distribution, by dividing income into several brackets and calculating a marginal tax rate $\tau(z)$ for each income range. Within each tax bracket, we determine the tax rate for that bracket by averaging the income ranges in that bracket. In other words, income levels falling within the income range are calculated as the average of that range. In particular, when calculating the top bracket rate, it is not convenient to calculate the average because its upper limit is infinite. So here $G(z)$ represents the total social welfare weight of incomes in the top bracket, when calculating $\alpha(z)$, we take the average income of the top income bracket as the average of the interval.
Elasticity $e(z)$ shows the sensitivity of the agent’s income z to changes in tax rates. Estimating elasticity is very difficult in the process of calculating tax rates, here we estimate the elasticity $e(z)$ using a regression method through income and marginal tax rates under varying fixed flat-tax systems, which produces an estimate equal to approximately 1.
$e(z)=\frac{1-\tau(z)}{z}\frac{dz}{d(1-\tau(z))}$
$\log(Z)=\hat{e}\cdot\log(1-\tau)+\log\left(\widehat{Z}^{0}\right)$
where $Z=\sum_{i}z_{i}$ when tax rates is $\tau$.
Appendix D Theoretical Proof
Proposition D.1.
The action value functions $Q_{i}$ and $\tilde{Q}_{t}$ are equivalent, i.e. $Q_{i}(\mathbf{s}_{t},\mathbf{a}_{t})=\tilde{Q}_{t}(s_{t}^{l},a_{t}^{l},L_{t}^{f})$, $\forall\mathbf{s}_{t}\in\mathcal{S}^{l}\times\{\mathcal{S}^{f}\}^{N}$, $\forall\mathbf{a}_{t}\in\mathcal{A}^{l}\times\{\mathcal{A}^{f}\}^{N}$, $\forall L_{t}^{f}\in\mathbb{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$.
Proof.
We commence by defining a partial order relation $\prec$ on $\mathcal{S}^{f}\times\mathcal{A}^{f}$. For any given vector $\mathbf{v}$, we construct a sorted counterpart $\tilde{\mathbf{v}}$, which maintains the same elements as $\mathbf{v}$ but orders them in ascending sequence based on $\prec$. For example, if $\mathbf{v}=\{(s_{1},a_{1}),(s_{2},a_{2}),(s_{3},a_{3})\}$, then $\tilde{\mathbf{v}}$ is arranged as $\{(s_{(1)},a_{(1)}),(s_{(2)},a_{(2)}),(s_{(3)},a_{(3)})\}$, where $(s_{(1)},a_{(1)})\prec(s_{(2)},a_{(2)})\prec(s_{(3)},a_{(3)})$). Exploiting the homogeneity of the follower agents, we can equate $Q(\mathbf{s}_{t},\mathbf{a}_{t})$ with $Q(s^{l}_{t},a^{l}_{t},\tilde{\mathbf{v}})$, where $\mathbf{v}$ encapsulates the joint state-action pairs of the followers, and $\tilde{\mathbf{v}}$ represents its sorted form. We further define $\mathcal{\mathbf{v}}$ as the collection of all such sorted vectors, and $\mathcal{L}$ as the ensemble of mean-field distributions. The empirical distribution of the followers’ state-action pairs, denoted as $L:\mathcal{V}\rightarrow\Delta(\mathcal{V})$, is expressed as $L^{f}_{t}(s,a)=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}_{\{(\tilde{s}_{i},\tilde{a}_{i})=(\tilde{s},\tilde{a})\}}$. The bijectiveness of the mapping $L$, ensured by the sorting process, leads to the conclusion that $Q_{t}(\mathbf{s}_{t},\mathbf{a}_{t})=Q_{t}(s^{l}_{t},a^{l}_{t},\mathbf{v})=Q_{t}(s^{l}_{t},a^{l}_{t},\tilde{\mathbf{v}})=\tilde{Q}_{t}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$. ∎

### v2

@@
-Learning Macroeconomic Policies based on Microfoundations: A Stackelberg Mean Field Game Approach
+Learning Macroeconomic Policies based on Microfoundations: A Dynamic Stackelberg Mean Field Game Approach
 Abstract
-Effective macroeconomic policies play a crucial role in promoting economic growth and social stability.
-This paper models the optimal macroeconomic policy problem based on the Stackelberg Mean Field Game (SMFG), where the government acts as the leader in policy-making, and large-scale households dynamically respond as followers.
-This modeling method captures the asymmetric dynamic game between the government and large-scale households, and interpretably evaluates the effects of macroeconomic policies based on microfoundations, which is difficult for existing methods to achieve.
-We also propose a solution for SMFGs, incorporating pre-training on real data and a model-free Stackelberg mean-field reinforcement learning (SMFRL) algorithm, which operates independently of prior environmental knowledge and transitions.
-Our experimental results showcase the superiority of the SMFG method over other economic policies in terms of performance, efficiency-equity tradeoff, and SMFG assumption analysis.
-This paper significantly contributes to the domain of AI for economics by providing a powerful tool for modeling and solving optimal macroeconomic policies.
-Macroeconomic policies; Stackelberg mean field game; reinforcement learning; multi-agent; social good
+The Lucas critique emphasizes the importance of considering the impact of policy changes on the expectations of micro-level agents in macroeconomic policymaking.
+However, the inherently self-interested nature of large-scale micro-agents, who pursue long-term benefits, complicates the formulation of optimal macroeconomic policies.
+This paper proposes a novel general framework named Dynamic Stackelberg Mean Field Games (Dynamic SMFG) to model such policymaking within sequential decision-making processes, with the government as the leader and households as dynamic followers.
+Dynamic SMFGs capture the dynamic interactions among large-scale households and their response to macroeconomic policy changes.
+To solve dynamic SMFGs, we propose the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm, which leverages the population distribution of followers to represent high-dimensional joint state and action spaces.
+In experiments, our method surpasses macroeconomic policies in the real world, existing AI-based and economic methods.
+It allows the leader to approach the social optimum with the highest performance, while large-scale followers converge toward their best response to the leader’s policy.
+Besides, we demonstrate that our approach retains effectiveness even when some households do not adopt the SMFG policy.
+In summary, this paper contributes to the field of AI for economics by offering an effective tool for modeling and solving macroeconomic policy-making issues.
 1 Introduction
-The formulation of macroeconomic policies is crucial for the sustained development of an economy (Schneider & Frey, 1988; Persson & Tabellini, 1999).
-Governments can adjust economic production, wealth distribution, social stability, and welfare through economic policies such as interest rates, taxation, and fiscal spending.
-Thus, how to effectively model and solve optimal macroeconomic policies, and simulate the effect of policy implementation is a very important issue.
-Existing research methods in economic policy primarily rely on simulations based on macroeconomic models like the Dynamic Stochastic General Equilibrium (DSGE) model (An & Schorfheide, 2007) and the Solow growth model (Brock & Taylor, 2010), or on empirical analysis (Robnik-Šikonja & Kononenko, 2003) assessing policy impacts through posterior analysis of historical data.
-However, these methods have limitations: (1) Difficulty in modeling dynamic interactions between the government and large-scale households, including feedback from households to policy and governmental adjustments in response to this feedback.
-(2) Challenge in solving the behavioral strategies of large-scale micro-agents.
-(3) The real economic and social world is highly complex, making it challenging for macroeconomic models to satisfy assumptions such as state transition and complete information.
-Addressing these challenges, this paper employs the Stackelberg Mean Field Game (SMFG) to model optimal macroeconomic policy problems.
-In this framework, the government, acting as the leader, initially sets policies, while a vast number of households, as followers, respond to these policies.
-However, considering the large scale of the followers, it is impractical for the leader to consider each follower’s dynamic response.
-Consequently, followers’ action-state distribution, analogous to the overall market information in economics, is introduced as a bridge between macro-level policy and micro-level individual decisions.
-The optimal decisions of the leader and followers depend on the followers’ overall distribution rather than on a specific agent.
-This assumption is crucial for modeling optimal macroeconomic policies as SMFGs and is easily comprehensible in the real world.
-For example, the government lowers property taxes to stimulate a sluggish housing market.
-This policy alters the market’s overall state, including housing prices and market demand.
-Households should consider both the government’s tax policy and the overall market information when making their purchasing decisions.
-Furthermore, this paper proposes a solution method for the optimal macroeconomic policy issues based on SMFG: Firstly, behavior cloning of households (followers) based on real data serves as the initialization for the policy network.
-This pre-training enhances the stability and performance of training and prevents ineffective solutions.
-Subsequently, a model-free algorithm, Stackelberg Mean-Field Reinforcement Learning (SMFRL), is introduced to solve SMFGs.
-This method does not rely on prior knowledge of the environment or transition information, offering advantages in solving complex economic issues.
-We conduct experiments on TaxAI (Mi et al., 2023), an environment for studying optimal tax policies, first to validate the performance of SMFRL against other model-free algorithms such as independent learning, MADDPG, and mean field MARL in solving SMFGs.
-Secondly, the paper compares the SMFG method with AI Economist, the Saez tax, the 2022 U.S.
-Federal tax, and free market policy, demonstrating the superiority of SMFG solutions and exploring the efficiency-equity tradeoff of these economic policies.
-Lastly, our tests on SMFG modeling assumptions reveal that our approach remains effective even if some households’ optimal decisions do not depend on market information.
+Macroeconomic policymaking is fundamental to the sustained development of an economy [52, 46].
+Governments influence economic production, wealth distribution, social stability, and welfare through economic policies such as interest rates, taxation, and fiscal spending [49].
+Therefore, accurately modeling and solving optimal macroeconomic policies and simulating the effects of their implementation are critical issues.
+According to the Lucas critique [39], predicting the impact of macroeconomic policies effectively requires modeling individual micro-level behavior—termed "microfoundations"—to forecast personal responses to policy changes and aggregate these into macroeconomic effects.
+Nonetheless, several challenges persist: dynamic interactions among micro-agents, between micro-agents and macro-authorities, the inherently selfish behavior of agents focused on optimizing long-term returns, and the curse of dimensionality due to the large-scale nature of agents such as households and firms.
+This paper introduces a novel framework termed Dynamic Stackelberg Mean Field Game (Dynamic SMFG) designed to address complex macroeconomic policy-making problems.
+Within this dynamic setting, the government, serving as the leader, adjusts policies iteratively, while a large population of households, acting as followers, respond accordingly.
+This process involves sequential decision-making, with both the leader and the followers aiming to optimize their cumulative long-term benefits.
+Given the leader’s policy, the SMFG reduces to a mean field game, which is widely used for modeling interactions among large-scale homogenous agents [33].
+In Dynamic SMFGs, followers aim to maximize their long-term benefits by providing the best response to the leader’s policy, while the leader’s goal is to find an optimal policy that effectively guides the massive followers toward achieving comprehensive macroeconomic objectives.
+Existing approaches often simplify Dynamic SMFGs to single-step decisions [45] or assume myopic followers [63], which do not adequately capture the complexities of macroeconomic policy-making.
+Therefore, we propose a data-driven reinforcement learning algorithm, Stackelberg Mean-Field Reinforcement Learning (SMFRL), to solve Dynamic SMFGs without requiring transition information or prior knowledge.
+The SMFRL algorithm utilizes a mean field approximation [28] to reduce the high-dimensional joint state and actions of followers to a manageable population distribution, allowing for the effective learning of a central critic Q function based on the overall followers’ distribution.
+Experiments conducted on TaxAI [41], an environment designed for studying optimal tax policies, first validate the performance of SMFRL in solving Dynamic SMFGs.
+The results indicate that SMFRL enables followers to closely achieve their best response while allowing the leader to approach the social optimum with the highest performance compared with baselines.
+Furthermore, the paper compares the Dynamic SMFG method with macroeconomic policies such as AI Economist [62], the Saez tax [50], the 2022 U.S.
+Federal tax, and free market policy [5], demonstrating the superiority of the Dynamic SMFG framework in macroeconomic policymaking.
+Considering real-world applications, our experiments show that even if a portion of households do not adopt the follower policy learned by the SMFG method, it does not adversely affect the utility, wealth, and income of households that do adopt the SMFG policy, which are in fact higher.
+However, the GDP of the economy may decrease with an increase in the number of households not supporting the SMFG policy, as they typically contribute less labor.
 In summary, the contributions of this paper are threefold:
 1.
-This paper is the first to model optimal macroeconomic policy problems using Stackelberg mean field games, linking macro-policies (the leader) with microfoundations (large-scale followers) through the followers’ overall distribution or average actions.
-It overcomes limitations such as dynamic interactions and feedback, and the dimensionality curse of large-scale households.
-2. This paper employs real data for pre-training and proposes a model-free Stackelberg mean field reinforcement learning algorithm to solve SMFGs, with experimental validation of its effectiveness.
-3.
-The paper empirically showcases the superiority of modeling and solving optimal macroeconomic policy problems by SMFGs.
-We also explore the efficiency-equity tradeoff in various economic policies and SMFG assumptions.
-Our code is shown in an anonymous GitHub repository  [^1] https://anonymous.4open.science/r/SMFG_macro_7740.
+It innovatively proposes a general framework based on Dynamic Stackelberg Mean Field Games to model macroeconomic policy-making, considering key issues such as microfoundations, dynamic interactions and feedback, rational expectations, and randomness.
+2.
+It introduces a data-driven Stackelberg Mean Field Reinforcement Learning algorithm to solve dynamic SMFGs without environmental transition information and prior knowledge, validated experimentally for its effectiveness.
+3. It empirically demonstrates the superiority of the SMFG method in modeling and solving optimal macroeconomic policy problems and discusses its robustness in real-world scenarios.
+The source code for the SMFG method is available in an anonymous GitHub repository  [^1] https://anonymous.4open.science/r/SMFG_macro_7740.
 2 Related work
@@
 In the field of economics, research on economic policy typically relies on theoretical model studies, empirical analysis, and econometric methods.
-Classic theoretical models, such as the IS-LM model (Hicks, 1980; Gali, 1992), are applicable for analyzing short-term policy effects but overlook long-term factors and price fluctuations.
-The aggregate demand-aggregate supply (AD-AS) model (Dutt, 2006; Lee et al., 1997) focuses on the relationship between aggregate demand and supply, integrating short-term and long-term factors, but simplifies macroeconomic dynamics.
-The Solow growth model (Brock & Taylor, 2010) provides a framework for economic growth but does not consider market imperfections and externalities.
-New Keynesian models (Blanchard & Galí, 2007; Gabaix, 2020) emphasize the stickiness of prices and wages, suitable for explaining economic fluctuations and policy interventions, but are based on strong assumptions.
-DSGE models (An & Schorfheide, 2007; Smets & Wouters, 2007), built on micro-foundations for macroeconomic forecasting, offer a consistent framework for policy analysis but are limited in handling nonlinearities and market imperfections.
-Empirical analysis (Ramesh et al., 2010; Vedung, 2017) primarily evaluates policy effects through historical data but faces limitations in data timeliness and accuracy.
-Econometric methods (Johnston & DiNardo, 1963; Davidson et al., 2004), such as regression and time series analysis, can quantify policy effects but are constrained by model settings and data quality.
+Classic theoretical models, such as the IS-LM model [24, 22], are applicable for analyzing short-term policy effects but overlook long-term factors and price fluctuations.
+The aggregate demand-aggregate supply (AD-AS) model [19, 34] focuses on the relationship between aggregate demand and supply, integrating short-term and long-term factors, but simplifies macroeconomic dynamics.
+The Solow growth model [10] provides a framework for economic growth but does not consider market imperfections and externalities.
+New Keynesian models [9, 21] emphasize the stickiness of prices and wages, suitable for explaining economic fluctuations and policy interventions, but are based on strong assumptions.
+DSGE models [1, 55], built on micro-foundations for macroeconomic forecasting, offer a consistent framework for policy analysis but are limited in handling nonlinearities and market imperfections.
+Empirical analysis [47, 59] primarily evaluates policy effects through historical data but faces limitations in data timeliness and accuracy.
+Econometric methods [30, 15], such as regression and time series analysis, can quantify policy effects but are constrained by model settings and data quality.
 2.2 AI for Economics
-Artificial intelligence (AI), such as reinforcement learning, may offer new perspectives for solving complex economic problems (Tilbury, 2022).
-In terms of macroeconomics, AI Economist (Zheng et al., ) employs curriculum learning on tax policy design, Trott et al.
-studies the collaboration between central and local governments under COVID-19, as well as monetary policy (Hinterlang & Tänzer, ; Chen et al., 2023), international trade (Sch, 2021), market pricing (Danassis et al., 2023).
+Artificial intelligence (AI), such as reinforcement learning, may offer new perspectives for solving complex economic problems [56].
+In terms of macroeconomics, AI Economist [62] employs curriculum learning on tax policy design, [58] studies the collaboration between central and local governments under COVID-19, as well as monetary policy [26, 12], international trade [51], market pricing [14].
 However, these studies consider the government as a single agent, neglecting the dynamic response of the households to policies.
-Koster et al.
-studies democratic AI by human voting models, Yaman et al.
-examines the impact of social sanction rules on labor division.
+[31] studies democratic AI by human voting models, [60] examines the impact of social sanction rules on labor division.
 However, these works are based on simplified settings and remain a gap in real-world economic policy.
-In terms of microeconomics, Shi (2021); Rui & Shi (2022); Atashbar & Aruhan Shi (2023) study the optimal saving and consumption problems of micro-agents, and other researchers explore rational expectation equilibrium (Kuriksha, 2021; Hill et al., ), multiple equilibria under real-business-cycle (Curry et al., 2022), the emergence of barter behavior (Johanson et al., 2022), optimal asset allocation and savings strategies (Ozhamaratli & Barucca, 2022).
+In terms of microeconomics, [53, 48, 3] study the optimal saving and consumption problems of micro-agents, and other researchers explore rational expectation equilibrium [32, 25], multiple equilibria under real-business-cycle [13], the emergence of barter behavior [29], optimal asset allocation and savings strategies [44].
 In summary, research on macroeconomic problems still lacks a bridge framework of macro policy and microfoundations, which has drawn our attention to the Stackelberg mean field game.
+Figure 1: The mapping from optimal macroeconomic policy problems (left) to the Dynamic Stackelberg Mean Field Games (right).
+The left represents the reality where the macro-level government implements fiscal or monetary policies to optimize GDP, welfare, or equity, and the micro-level consists of large-scale households engaging in economic activities.
+In dynamic SMFGs, the government acts as the leader, and households as followers, respond to the leader’s policy.
 2.3 Stackelberg Mean Field Game
-The Stackelberg Mean Field Game models one leader and a large number of homogeneous followers, and current research can be divided into model-based methods and model-free methods.
-Model-based approaches typically involve complex theoretical analysis and precise mathematical modeling.
-For instance, Guo et al.
-(2022) reconceptualizes SMFGs as minimax optimization problems, while Dayanikli & Lauriere (2023); Bergault et al.
-(2023); Bensoussan et al.
-(2015); Fu & Horst (2020) initially resolve the followers’ mean field equilibrium using forward-backward stochastic functional differential equations, subsequently deriving the leader’s optimal control.
-Huang & Yang (2020); Bensoussan et al.
-(2015) explore linear-quadratic Stackelberg games.
-These methods provide a theoretically robust foundation but may require simplified assumptions for practical applications, particularly in complex real-world scenarios.
-Notably, Aurell et al.
-(2022) extends the mean field game model to capture epidemic dynamics, demonstrating potential applications in public health management.
-In contrast, model-free methods offer a way to handle complex situations that traditional numerical methods struggle with, especially in data-rich environments.
-Campbell et al.
-(2021) employs a modified deep backward stochastic differential equation (BSDE) method for solving the followers’ equilibrium, yet the leader’s choices are limited.
-In conclusion, we plan to use SMFGs to model macroeconomic policy issues and propose a model-free algorithm to solve this complex task.
-Figure 1: The mapping from optimal macroeconomic policy problems (left) to the Stackelberg Mean Field Games (right).
-The left represents the reality where the macro-level government implements fiscal or monetary policies to optimize GDP, welfare, or equity, and the micro-level consists of large-scale households engaging in economic activities and forming the market.
-In SMFGs, the government corresponds to the leader, households to the followers, and market information to the followers’ state-action distribution.
-3 Macroeconomic Policy Formulation
-3.1 Economic Modeling
-In reality, an economy comprises at least one government and large-scale households.
-The diverse economic activities of households, such as production, trade, and consumption, aggregate to form the market.
-The government can maintain economic growth and social stability (Figure 1 left) via economic policies.
-The following is a mathematical model for optimizing tax policy:
-Households
-At timestep $t$, a household holds assets $a_{t}$.
-Its income $i_{t}$ consists of labor income $W_{t}e_{t}h_{t}$, and capital income $r_{t}a_{t}$.
-Here, $W_{t}$ and $r_{t}$ are the wage and interest rates determined by the labor and capital markets, respectively, while $e_{t}$ indicates the private education level.
-Households are obligated to pay taxes as per the government’s tax policy $T_{t}(i_{t},a_{t})$.
-Each household decides on its consumption $c_{t}\in R^{+}$, working hours $h_{t}\in R^{+}$, and next savings $a_{t+1}\in R^{+}$ to maximize long-term utility until $T_{max}$, subject to the budget constraint:
-$\max_{a_{t+1},h_{t}}$ $\quad\mathbb{E}_{0}\sum_{t=0}^{T_{max}}u(c_{t},h_{t})$
+Stackelberg Mean Field Games (SMFG) model interactions between the leader and numerous homogeneous followers, employing both model-based and model-free approaches.
+Model-based methods, such as those presented in [23], treat SMFGs as minimax optimization problems, streamlining the leader’s strategy to a single optimal action rather than a comprehensive policy.
+Studies by [7, 20, 16, 8] utilize forward-backward stochastic functional differential equations to determine the followers’ mean field equilibrium before computing the leader’s optimal strategy.
+This includes addressing terminal state constraints [20] and simplifying scenarios via linear-quadratic formulations [6, 37, 43, 18, 27].
+Such methods, though theoretically robust, often rely on simplifying real-world complexities and require extensive knowledge of environmental transitions.
+Extending SMFG to model epidemic dynamics [4] showcases its broader applicability in public health management.
+Conversely, model-free approaches leverage complex, data-rich environments without the need for prior environmental knowledge.
+[11] employs a modified deep backward stochastic differential equation framework to establish followers’ equilibrium, despite constraints on leader decisions.
+[45] investigates applications in privacy protection through a single-step decision model.
+[42] delves into locally optimal defensive strategies given predefined attacker trajectories.
+[35] learns followers’ transition functions from empirical data, and facilitates SMFG solutions through the empirical Fokker-Planck equation.
+However, in real-world settings, the sequential decision-making between a leader and numerous followers to optimize long-term gains is common but inadequately studied.
+This paper proposes a data-driven reinforcement learning algorithm for solving the aforementioned problem, termed the Dynamic Stackelberg Mean Field Game, without reliance on environmental transition information or prior knowledge.
+3 Stackelberg Mean Field Game Modeling
+3.1 Macroeconomic Policy-Making
+Macroeconomic policy-making involves formulating effective policies to regulate and guide an economy toward achieving sustainable growth, low unemployment, and stable prices.
+In addressing these challenges, several key elements must be considered: ① Microfoundations: According to Lucas critique, basing macroeconomic models on microeconomic behaviors is crucial.
+This ensures that the actions of individual agents, such as households and firms, aggregate to produce macroeconomic outcomes.
+② Dynamic Interactions: Extensive interactions occur both between the macro-agent (the government) and a multitude of micro-agents (households), and within groups of households.
+These interactions can lead to complex economic dynamics and endogenous shocks.
+③ Rational Expectations: Agents are typically modeled as rational and self-interested, making decisions based on available information and their expectations about future policies and economic conditions.
+④ Stochastic Elements: Incorporating randomness and shocks, such as technological changes, policy shifts, and external economic shocks, is essential for understanding variability in economic outcomes.
+⑤ Feedback Mechanisms: Feedback loops exist within the economy where macro-policy impacts micro-agents, and the micro-agents’ behaviors, in turn, influence policy decisions.
+Given the aforementioned key points, Dynamic Stackelberg Mean Field Games (Dynamic SMFGs) are ideal for macroeconomic policy-making due to:
+1.
+Leader-Follower Dynamics (②⑤): The government, acting as the leader, sets policies first.
+Households, as followers, adjust their behaviors based on these policies, which captures the strategic interactions between policy-setters and economic agents.
+2.
+Mean Field Interaction (①⑤): The collective behavior of numerous micro-agents is efficiently modeled as a "mean field", which simplifies the analysis of large-scale interactions and their overall impact on the macroeconomy.
+3. Dynamic Strategy (②③④): The dynamic nature of these games reflects the evolving economic conditions and policy-making, enabling the model to adapt to new information and economic shocks.
+3.2 Dynamic Stackelberg Mean Field Games
+Following this, we consider a dynamic SMFG over discrete periods $t\in\{0,\ldots,T\}$ for macroeconomic policy-making issues: At each time $t$, the leader first selects an action $a_{t}^{l}\in\mathcal{A}^{l}$ based on the leader’s state $s_{t}^{l}\in\mathcal{S}^{l}$ and policy $\pi^{l}:\mathcal{S}^{l}\rightarrow\mathcal{A}^{l}$.
+Subsequently, followers simultaneously determine their actions based on the leader’s action $a^{l}_{t}$ and their private states $s^{f}_{t}\in\mathcal{S}^{f}$.
+The representative follower’s action $a_{t}^{f}\in\mathcal{A}^{f}$ is derived from a shared policy $\pi^{f}:\mathcal{S}^{f}\times\mathcal{A}^{l}\rightarrow\mathcal{A}^{f}$.
+The sequences $\{\pi_{t}^{l}\}_{t=0}^{T}$ and $\{\pi_{t}^{f}\}_{t=0}^{T}$ are denoted as $\pi^{l}$ and $\pi^{f}$, respectively.
+Followers’ Mean Field Game
+Given the leader’s policy $\pi^{l}$, the dynamic SMFG is simplified into a mean field game for the followers.
+At time $t\in\{0,...,T-1\}$, given the leader’s action $a^{l}_{t}$, the representative follower’s state $s^{f}_{t}$ and action $a^{f}_{t}$, and the population state-action distribution of all followers denoted as $L^{f}_{t}\in\mathcal{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$, the representative follower receives a reward $r^{f}(s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$ and move to the next state $s_{t+1}^{f}\sim\mathbb{P}(\cdot|s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$.
+Each follower aims to find the optimal policy $\pi^{f}$ that maximizes his cumulative reward over the time horizon:
+$J^{f}(\pi^{l},\pi^{f},L^{f})=\mathbb{E}^{f}\left[\sum_{t=0}^{T}r^{f}(s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})\right]$ (1)
+where $\mathbb{E}^{f}$ denotes the expectation with respect to initial state $s_{0}^{f}\in\mu_{0}^{f}$, $(s_{t+1}^{l},s_{t+1}^{f})\sim P(\cdot\mid s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$ for $t\in\{0,...,T-1\}$, and $a^{l}_{t}\sim\pi^{l}(\cdot\mid s_{t}^{l}),a^{f}_{t}\sim\pi^{f}(\cdot\mid s_{t}^{f},a^{l}_{t})$ for $t\in\{0,...,T\}$.
+Definition 3.1.
+[Followers’ Best Response for Leader’s Policy] Given any leader’s policy $\pi^{l}\in\Pi^{l}$ and the followers’ state-action distribution $L^{f}\in\mathcal{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$, the followers’ best response policy ${\pi^{f}}^{*}(\pi^{l},L^{f})$ is defined as
+${\pi^{f}}^{*}(\pi^{l},L^{f})\in\arg\max_{\pi^{\prime}}J^{f}(\pi^{l},\pi^{\prime},L^{f}).$
+Leader’s Game
+At time $t\in\{0,...,T-1\}$, given the leader’s state $s^{l}_{t}$, action $a^{l}_{t}$, and the followers’ state-action distribution $L_{t}^{f}$, the leader will receive a reward $r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})$, and move to a new state $s_{t+1}^{l}$ according to the transition probabilities $\mathbb{P}(\cdot|s_{t}^{l},a_{t}^{l},L^{f}_{t})$.
+The leader aims to find the optimal policy $\pi^{l}$ to maximize the total expected reward:
+$J^{l}(\pi^{l},\pi^{f},L^{f})=\mathbb{E}^{l}\left[\sum_{t=1}^{T}r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})\right],$ (2)
+where $\mathbb{E}^{l}$ denotes the expectation under this probability distribution for the leader when $s_{0}^{l}\sim\mu_{0}^{l},s_{t+1}^{l}\sim\mathbb{P}(\cdot|s_{t}^{l},a_{t}^{l},L^{f}_{t})$ for $t\in\{0,...,T-1\}$, and $a_{t}^{l}\sim\pi_{t}^{l}(\cdot\mid s_{t}^{l})$ for $t\in\{0,...,T\}$.
+Definition 3.2 (Leader’s Optimal Policy in Dynamic Stackelberg Mean Field Games).
+Considering the followers’ best response $(\pi^{f},L^{f})$ for the leader’s policy, which satisfies Definition 3.1, the leader’s optimal policy ${\pi^{l}}^{*}$ in dynamic Stackelberg mean field games is defined as follows:
+${\pi^{l}}^{*}$ $\in\mathop{\arg\max}_{{\pi^{l}}^{\prime}}J^{l}({\pi^{l}}^{\prime},\pi^{f},L^{f}),$
 […]
-Government
-The government aims to foster economic growth, maintain social equity, and enhance societal welfare through three fiscal tools: debt $B_{t}\in R^{+}$, taxation policy $T_{t}(i_{t},a_{t}):R^{+}\times R^{+}\rightarrow R^{+}$, and spending $G_{t}\in R^{+}$.
-$\max_{B_{t},T_{t},G_{t}}\quad\mathbb{E}_{0}\sum_{t=0}^{T_{max}}\left[\alpha_{1}Y_{t}+\alpha_{2}S(\mathcal{I}_{t},\mathcal{W}_{t})+\alpha_{3}\sum_{i}^{N}u_{t}^{i}\right],$
-where $\alpha_{1},\alpha_{2},\alpha_{3}$ are weight parameters reflecting the government’s preference for different objectives.
-Economic output (e.g., GDP) can be modeled using the Cobb–Douglas production function $Y_{t}=F(K_{t},\mathcal{L}_{t})$, with production capital $K_{t}$ and labor $\mathcal{L}_{t}$.
-$S(\mathcal{I}_{t},\mathcal{W}_{t})$ denotes the social equity function, where $\mathcal{I}_{t},\mathcal{W}_{t}$ are the income and wealth distributions among households.
-The sum of households’ utilities $\sum_{i}^{N}u_{t}^{i}$ represents the overall societal welfare.
-3.2 Stackelberg Mean Field Game Modeling
-In the aforementioned issue, the government (as the leader) formulates policies to guide the decisions of households (as followers), thereby collectively achieving macroeconomic objectives, such as economic growth.
-Households provide feedback to government policies, creating an asymmetric game with leader-follower dynamics.
-However, the dynamic interactions among large-scale households extend beyond the standard Stackelberg games.
-Therefore, this study proposes modeling the problem as a Stackelberg Mean Field Game (Figure 1), based on the following assumptions:
-  1. 1.
-    Stackelberg assumption: the government (leader) sets policies first, and households (followers) adjust their behavior based on these policies.
-Both the leader and followers dynamically optimize their strategies, gradually reaching equilibrium.
-  2. 2.
-    Mean field assumption: households (followers) are assumed to be homogeneous agents.
-The optimal policies of the leader and followers depend on the overall distribution or average behavior of the followers, rather than a specific one.
-For example, both governments and households should consider overall market information when making decisions in reality.
-Following this, the optimal macroeconomic policy issues can be rewritten as the following game proceeding: For any given time $t\in\{0,\ldots,T_{max}\}$, the leader selects an action $a_{t}^{l}\in\mathcal{A}^{l}$ based on the leader policy $\pi_{t}^{l}(\cdot|s_{t}^{l})$, given the state $s_{t}^{l}$.
-Subsequently, the representative follower, located in state $s_{t}^{f}\in\mathcal{S}^{f}$, chooses an action $a_{t}^{f}\in\mathcal{A}^{f}$ according to the policy $\pi_{t}^{f}(\cdot|s_{t}^{f},a^{l}_{t})$.
-The sequences $\{\pi_{t}^{l}\}_{t=0}^{T-1}$ and $\{\pi_{t}^{f}\}_{t=0}^{T-1}$ are denoted as $\pi^{l}$ and $\pi^{f}$, respectively.
-Followers’ Mean Field Game
-Given the leader’s policy $\pi^{l}$, the SMFG is simplified into a mean field game for the followers.
-At time $t$, given the leader’s action $a^{l}_{t}$, the representative follower’s state $s_{t}^{f}$ and action $a_{t}^{f}$ determine the followers’ state-action distribution as $L^{f}_{t}(s_{t}^{f},a_{t}^{f})=\mathbb{P}(s_{t}^{f},a_{t}^{f})$.
-The representative follower then receives a reward $r^{f}(s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$ and transitions to the next state $s_{t+1}^{f}\sim P(\cdot|s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$.
-Each follower aims to find the optimal policy $\pi^{f}$ that maximizes his cumulative reward over the time horizon:
-$J^{f}(\pi^{l},\pi^{f},L^{f})$ $=\mathbb{E}^{f}\left[\sum_{t=1}^{T}r^{f}(s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})\right]$
-$=\mathbb{E}^{f}\left[\sum_{t=1}^{T}u(c_{t},h_{t})\right],$
-where $\text{where }c_{t}$ $=W_{t}e_{t}h_{t}+r_{t}a_{t}-T_{t}(i_{t},a_{t})-a_{t+1},$
-where $\mathbb{E}^{f}$ denotes the expectation with respect to initial state $s_{0}^{f}\sim\mu_{0}^{f}$, $s_{t+1}^{f}\sim P(\cdot\mid s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$ and $a^{f}_{t}\sim\pi^{f}(\cdot\mid s_{t}^{f},a^{l}_{t})$ for each $t$.
-For households in economic problems, the follower’s state $s^{f}_{t}$ incorporates asset $a_{t}$ and education level $e_{t}$, and the follower’s actions $a^{f}_{t}$ include next savings $a_{t+1}$ and working hours $h_{t}$.
-The leader’s action $a^{l}_{t}$ denotes tax function $T_{t}(i_{t},a_{t})$, and followers’ distribution $L^{f}_{t}$ includes wage rate $W_{t}$ and interest rate $r_{t}$.
-Given the leader’s policy $\pi^{l}$, the best response operator of followers is defined as
-${\pi^{f}}^{*}\in\mathrm{BR}^{f}(\pi^{l},L^{f})=\arg\max_{\pi^{\prime}}J^{f}(\pi^{l},\pi^{\prime},L^{f}).$
-Stackelberg Leader’s Game
-At time $t$, given the leader’s state $s^{l}_{t}$, action $a^{l}_{t}$, and the followers’ state-action distribution $L_{t}^{f}$, the leader will receive a reward $r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})$, and transition to a new state $s_{t+1}^{l}$ according to the transition probabilities $P(\cdot|s_{t}^{l},a_{t}^{l},L^{f}_{t})$.
-The leader aims to find the best policy $\pi^{l}$ to maximize the total expected rewards:
-$J^{l}(\pi^{l},\pi^{f},L^{f})=\mathbb{E}^{l}[\sum_{t=1}^{T}r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})],$
-$r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})=\alpha_{1}Y_{t}+\alpha_{2}S(\mathcal{I}_{t},\mathcal{W}_{t})+\alpha_{3}\sum_{i}^{N}u_{t}^{i}$
-$=\alpha_{1}F(K_{t},\mathcal{L}_{t})+\alpha_{2}S(\mathcal{I}_{t},\mathcal{W}_{t})+\alpha_{3}U(T_{t},W_{t},r_{t})$
-where $\mathbb{E}^{l}$ denotes the expectation under this probability distribution for the leader when $s_{0}^{l}\sim\mu_{0}^{l},s_{t+1}^{l}\sim P(\cdot|s_{t}^{l},a_{t}^{l},L^{f}_{t})$ and $a_{t}^{l}\sim\pi_{t}^{l}(\cdot\mid s_{t}^{l})$ for $t\in\{0,...,T-1\}$.
-For the government in economic problems, the leader’s state $s^{l}_{t}$ is characterized by production capital $K_{t}$, and the leader’s actions $a^{l}_{t}$ encompass government spending $G_{t}$ and a taxation policy $T_{t}(\cdot)$.
-The followers’ state-action distribution $L^{f}_{t}$ includes aggregate labor $\mathcal{L}_{t}$, income distribution $\mathcal{I}_{t}$, wealth distribution $\mathcal{W}_{t}$, wage rate $W_{t}$, and interest rate $r_{t}$, these all belong to market information.
-$U(\cdot)$ represents a measurable function for $T_{t},W_{t},r_{t}$.
-Therefore, we can define the leader’s best response operator given $\pi^{f},L^{f}$:
-${\pi^{l}}^{*}\in\text{BR}^{l}(\pi^{f},L^{f}):=\mathop{\arg\max}_{\pi^{l}}J^{l}(\pi^{l},\pi^{f},L^{f}).$
-We then define the transition operator $\mathcal{F}$ as $L^{f}=\mathcal{F}(\pi^{l},\pi^{f},L^{f}_{0},\mu_{0})$ if the evolution of population distribution $L$ satisfies the Mckean-Vlasov equation (1).
-$L^{f}_{t+1}(s_{t+1}^{f},a_{t+1}^{f})=$ $\sum_{s_{t}^{l},a_{t}^{l},s_{t}^{f},a_{t}^{f}}L^{f}_{t}(s_{t}^{f},a_{t}^{f})\pi_{t}^{l}(a_{t}^{l}|s_{t}^{l})\mu_{t}^{l}(s_{t}^{l})$ (1)
-$\quad P(s_{t+1}^{f}|s_{t}^{f},a_{t}^{f},a_{t}^{l})\pi_{t+1}^{f}(a_{t+1}^{f}|s_{t+1}^{f}).$
-$\mu_{t+1}^{l}(s_{t+1}^{l})=$ $\sum_{s_{t}^{l},a_{t}^{l}}\mu_{t}^{l}(s_{t}^{l})\pi_{t}^{l}(a_{t}^{l}|s_{t}^{l})P(s_{t+1}^{l}|s_{t}^{l},a_{t}^{l}).$
-Definition 3.1.
-The Stackelberg mean field equilibrium is a tuple $({\pi^{l}}^{*},{\pi^{f}}^{*},{L^{f}}^{*})$, which satisfies the following conditions:
-${\pi^{l}}^{*}\in\mathrm{BR}^{l}({\pi^{f}}^{*},{L^{f}}^{*}),{\pi^{f}}^{*}\in\mathrm{BR}^{f}({\pi^{l}}^{*},{L^{f}}^{*}),$
-${L^{f}}^{*}=\mathcal{F}({\pi^{l}}^{*},{\pi^{f}}^{*},L^{f}_{0},\mu_{0}).$
-4 Methods
-Solving the optimal macroeconomic policy is very complex in the real world.
-Reliance solely on mathematical models for simulation and problem-solving inevitably creates a gap with actual scenarios, thus incorporating real historical data is essential.
-Nonetheless, a singular focus on historical data for the empirical analysis of economic policies inadequately captures the complex and asymmetric dynamics of interactions between the government and households.
-Therefore, our solution method is divided into two steps: firstly, we pre-train the follower’s policy based on real data using behavior cloning (BC).
-Secondly, this paper proposes a model-free Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm to solve the Stackelberg Mean Field Games (SMFGs) between the government and large-scale households.
-4.1 Pretain by Behavior Cloning
-Based on real historical data, we fetch a large number of followers’ state-action pairs $\{s^{f},a^{f}\}$ from a real-data buffer $\mathcal{D}_{real}$ for behavior cloning.
-For different settings of network structures, we have chosen two types of loss: when the neural network outputs a probability distribution of actions, we use the negative log-likelihood loss (NLL loss); when the neural network outputs action values, we employ the mean square error loss (MSE loss).
-Our goal is to find the optimal parameters $\theta$ as the follower’s policy network $\pi_{\theta}$ initialization, thereby minimizing the loss to its lowest convergence.
-$\min_{\theta}\mathcal{L}_{NLL}$ $=-\mathbb{E}_{s^{f},a^{f}\sim\mathcal{D}}\log{\pi_{\theta}(a^{f}\mid s^{f})},$
-$\min_{\theta}\mathcal{L}_{MSE}$ $=\mathbb{E}_{s^{f},a^{f}\sim\mathcal{D}}\left(a^{f}-a\right)^{2}|_{a=\pi_{\theta}(s^{f})}.$
-4.2 Stackelberg Mean Field Reinforcement Learning
-To solve SMFGs in sequential decision-making, we design a Stackelberg mean field reinforcement learning algorithm based on centralized training with decentralized execution.
-Consider a game with a leader and $N$ homogeneous follower agents.
-The leader makes decisions first, and all followers make simultaneous decisions after observing the leader’s actions.
-The leader aims to find a policy $\pi^{l}:\mathcal{S}^{l}\rightarrow\mathcal{A}^{l}$ to optimize his long-term objective, which is often challenging and requires the cooperation of all followers.
-Each follower agent seeks to find a policy $\pi^{f}:\mathcal{S}^{f}\times\mathcal{A}^{l}\rightarrow\mathcal{A}^{f}$ to optimize long-term utility after observing the current leader’s action $a^{l}_{t}=\pi^{l}(s^{l}_{t})$.
-The joint action of the leader and follower agents $\mathbf{a}_{t}=\{a^{l}_{t},a^{f1}_{t},...,a^{fN}_{t}\},a^{fi}_{t}=\pi^{f}(s^{fi}_{t})$ impacts the environment and results in a reward $\mathbf{r}_{t}=\{r^{l}_{t},r^{f1}_{t},...,r^{fN}_{t}\}$ and the next observation $\mathbf{s}_{t+1}=\{s^{l}_{t+1},s^{f1}_{t+1},...,s^{fN}_{t+1}\}$.
-Here the experience replay buffer $\mathcal{D}$ contains the tuples $(\mathbf{s}_{t},\mathbf{s}_{t+1},\mathbf{a}_{t},\mathbf{r}_{t},L^{f}_{t})$, recording experiences of all agents.
-$Q_{i}(\mathbf{s}_{t},\mathbf{a}_{t})$ is a standard centralized action-value function that takes as input the joint state $\mathbf{s}_{t}$ and joint action $\mathbf{a}_{t}$ and outputs the Q-value for agent $i$.
-However, due to numerous follower agents, the dimensions of both the joint state and joint action increase significantly with $N$, making the Q-function infeasible to learn.
-We equate the centralized action-value function $Q_{i}(\mathbf{s}_{t},\mathbf{a}_{t})$ to $\tilde{Q}_{t}(s_{t}^{l},a_{t}^{l},L_{t}^{f})$ based on mean field theory.
-The specific theoretical description is shown in the Appendix D.
-4.3 Implementation
-Furthermore, we consider deterministic policies for both a leader and its followers in a multi-agent system.
+where the followers’ state-action distribution $L^{f}$ satisfies the following Mckean-Vlasov equation:
+$L^{f}_{t+1}(s_{t+1}^{f},a_{t+1}^{f})=\sum_{s_{t}^{l},a_{t}^{l},s_{t}^{f},a_{t}^{f}}L^{f}_{t}(s_{t}^{f},a_{t}^{f})\pi_{t}^{l}(a_{t}^{l}\mid s_{t}^{l})\mu_{t}^{l}(s_{t}^{l})P(s_{t+1}^{f}\mid s_{t}^{f},a_{t}^{f},a_{t}^{l})\pi_{t+1}^{f}(a_{t+1}^{f}\mid s_{t+1}^{f}),$
+$\mu_{t+1}^{l}(s_{t+1}^{l})=\sum_{s_{t}^{l},a_{t}^{l}}\mu_{t}^{l}(s_{t}^{l})\pi_{t}^{l}(a_{t}^{l}\mid s_{t}^{l})P(s_{t+1}^{l}\mid s_{t}^{l},a_{t}^{l}).$ (3)
+In conclusion, to study the issue of macroeconomic policy-making, we model it as a dynamic Stackelberg mean field game.
+This modeling approach aims to learn the optimal policy of the leader, namely the social planner, while considering the followers’ best response over discrete timesteps (Definition 3.2).
+3.3 Metrics for Dynamic SMFGs
+We will utilize the following three metrics to evaluate the effectiveness of the leader and follower policies: (1) The leader’s payoff, which indicates the performance of the leader’s policy in optimizing the leader’s objective; (2) Exploitability, which measures the deviation of the followers’ policy from the best response to the leader’s policy; (3) Social welfare, which assesses the deviation of the current state from the social optimum.
+Leader’s Payoff
+We define the leader’s payoff using the long-term expected rewards of the leader’s policy $\pi^{l}$ over discrete timesteps, as detailed in Equation 2.
+Exploitability
+Exploitability is a widely used metric for evaluating convergence and measuring the distance between the representative follower’s policy $\pi^{f}$ and the best response ${\pi^{f}}^{*}$ for the leader’s policy $\pi^{l}$.
+Formally, it quantifies the extent to which a single follower can improve their payoff by deviating from the population’s behavior to adopt a different policy.
+Given the leader’s policy $\pi^{l}$ and the followers’ state-action distribution $L^{f}$, the exploitability $\mathcal{E}(\pi^{f};\pi^{l})$ of the follower’s policy $\pi^{f}$ is defined by comparing the value functions induced by $\pi^{f}$ and its best response $\pi^{f*}$:
+$\mathcal{E}(\pi^{f};\pi^{l})=J^{f}(\pi^{l},{\pi^{f}}^{*},L^{f})-J^{f}(\pi^{l},{\pi^{f}},L^{f}).$
+Social Optimum and Social Welfare
+In economic theory, the Social Optimum describes a state in which the allocation of resources achieves maximum efficiency, as measured by social welfare [2, 40].
+Given the leader’s policy $\pi^{l}$ and the representative follower’s policy $\pi^{f}$ among large-scale followers, social welfare $\mathcal{SW}(\pi^{l},\pi^{f})$ is approximately calculated as the sum of the utility functions of the $N$ followers in Equation 1 when $N\to\infty$:
+$\mathcal{SW}(\pi^{l},\pi^{f})=\sum^{N}J^{f}\left(\pi^{l},\pi^{f},L^{f}\right).$
+4 Stackelberg Mean Field Reinforcement Learning
+However, solving for the leader’s optimal policy in dynamic SMFGs, as defined in Definition 3.2, presents several challenges, such as vast state and action spaces, strategic dependencies and information asymmetry between the leader and followers, and potentially conflicting objectives including the leader’s payoff, followers’ utilities, and social welfare.
+These complexities make devising a model-based algorithm with theoretical proof extremely difficult.
+Previous works have attempted simplifications using the myopic followers setting [63] or linear-quadratic MFGs [6], but this approach fails to align with macroeconomic policy-making issues.
+Consequently, we have designed the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm to address the above limitations.
+This is a model-free approach that operates independently of environmental transition information.
+Figure 2: The architecture of SMFRL algorithm.
+Considering the dynamic SMFGs with a leader and $N$ homogeneous follower agents, at any time $t\in\{0,\ldots,T-1\}$, the joint action of the leader and follower agents $\mathbf{a}_{t}=\{a^{l}_{t},a^{f1}_{t},...,a^{fN}_{t}\},a^{fi}_{t}=\pi^{f}(s^{fi}_{t})$ impacts the environment and results in a reward $\mathbf{r}_{t}=\{r^{l}_{t},r^{f1}_{t},...,r^{fN}_{t}\}$ and the next state $\mathbf{s}_{t+1}=\{s^{l}_{t+1},s^{f1}_{t+1},...,s^{fN}_{t+1}\}$.
+The experience replay buffer $\mathcal{D}$ contains the tuples $(\mathbf{s}_{t},\mathbf{s}_{t+1},\mathbf{a}_{t},\mathbf{r}_{t},L^{f}_{t})$, recording the experiences of all agents.
 Specifically, we define the leader’s policy as $\pi_{\theta^{l}}$ parameterized by $\theta^{l}$ (abbreviated as $\pi^{l}$), and a shared policy for the followers denoted as $\pi_{\theta^{f}}$ with parameters $\theta^{f}$ (abbreviated as $\pi^{f}$).
-For the leader agent
-The policy network $\pi_{\theta^{l}}$ of the leader agent, i.e. the actor, is trained by the sampled policy gradient (Silver et al., 2014):
-$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{-},L^{f}_{t})|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}\right]$
-In the subsequent step, we employ a neural network with parameters $\phi^{l}$ to approximate the leader agent’s action-value function, denoted as $\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$ or $\tilde{Q}^{l}$ for brevity.
-It is updated by minimizing a loss function:
-$\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)^{2}\right],$
-$y^{l}_{t}=r_{t}^{l}+\gamma\tilde{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L^{f}_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})},$ (2)
-where $y^{l}_{t}$ represents the target Q value, computed using the parameters $\phi^{l}_{-}$, the next population distribution $L^{f}_{t+1}$ is based on the Mckean-Vlasov equation (1), and $\gamma$ is discount factor.
+The SMFRL algorithm incorporates several settings (shown in Figure 2) to address the aforementioned challenges:
+Mean Field Approximation
+To mitigate the problem of non-stationarity, we consider a central critic Q function $Q(\mathbf{s}_{t},\mathbf{a}_{t}):\mathcal{S}^{l}\times\{\mathcal{S}^{f}\}^{N}\times\mathcal{A}^{l}\times\{\mathcal{A}^{f}\}^{N}\to\mathbb{R}$ for each agent.
+However, due to the high dimensionality of the joint state and action, the standard Q network is infeasible to learn.
+Thus, the SMFRL algorithm employs a mean field approximation, characterizing the followers’ state-action pairs $(\mathbf{s}^{f}_{t},\mathbf{a}^{f}_{t})$ by using the followers’ population distribution $L^{f}_{t}$, which can be presented through various metrics such as mean, variance, probability density functions, skewness, and kurtosis.
+The theoretical proof is detailed in Appendix F.
+This approach yields $Q^{l}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$ and $Q^{f}(s^{f}_{t},a^{l}_{t},a^{f}_{t},L^{f}_{t})$ for evaluating the leader and follower agents respectively.
+The mean field approximation effectively reduces the dimensionality of the inputs to the Q function, thereby facilitating the learning of the critic Q network and ensuring the Q function can evaluate leader and follower policies based on the overall follower distribution.
+Leader’s Update for Optimal Policy
+In SMFRL, the leader’s policy $\pi_{\theta^{l}}$, is optimized using a deterministic policy gradient approach [54]. The policy gradient for the leader can be estimated as:
+$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{-},L^{f}_{t})|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}\right],$
+where ${Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$, the leader’s action-value function, is approximated by a neural network parameterized by $\phi^{l}$.
+This network is periodically updated to minimize the loss:
+$\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)^{2}\right],$
+where the target value $y^{l}_{t}$ is given by:
+$y^{l}_{t}=r_{t}^{l}+\gamma{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L^{f}_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})},$
+and $\phi^{l}_{-}$ represents the parameters of the target Q network, $L^{f}_{t+1}$ is determined based on Mckean-Vlasov equation (3), and $\gamma$ is the discount factor.
 Differentiating the loss function $\mathcal{L}(\phi^{l})$ yields the gradient utilized for training:
-$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=$
-$\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{l}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right]$
-For the representative follower agent
-The policy network $\pi_{\theta^{f}}$ is trained by the policy gradient:
-$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L^{f}_{t}\sim\mathcal{D}}\Big{[}$
-$\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{-},a^{l}_{t},L^{f})|_{a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})}\Big{]},$
-Similar to Mean Field Approximation (Yang et al., 2018), the standard centralized action-value function $Q(\mathbf{s}_{t},\mathbf{a}_{t})$ is approximated as $\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})$ with parameters $\phi^{f}$, which is updated by minimizing the MSE loss function:
-$\mathcal{L}(\phi^{f})=$ $\mathbb{E}_{{s}_{t}^{f},{a}_{t}^{f},a^{l}_{t},L^{f}_{t},s_{t+1}^{f}\sim\mathcal{D}}\left(y^{f}_{t}-\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)^{2}$
-$y^{f}_{t}=$ $r^{f}_{t}+\gamma\tilde{Q}_{\phi^{f}_{-}}(s^{f}_{t+1},a^{f}_{t+1},a^{l}_{t+1},L^{f}_{t+1}),$
-where the next leader’s action $a^{l}_{t+1}=\mu_{\phi^{l}_{-}}(s^{l}_{t+1})$, the next population distribution $L^{f}_{t+1}=L(\mathbf{s}_{t+1}^{f},\mathbf{a}_{t+1}^{f})$, the next joint followers’ action $\mathbf{a}_{t+1}^{f}=\{a^{f1}_{t+1},...,a^{fN}_{t+1}\}$, $a^{fi}_{t+1}=\mu(s^{fi}_{t+1},a^{l}_{t+1})$.
-Differentiating the loss function $\mathcal{L}(\phi^{f})$ yields the gradient utilized for training:
-$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\mathcal{D}}\Big{[}\left(y^{f}_{t}-\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)$
-$\nabla_{\phi^{f}}\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\Big{]}$
-The full algorithm (1) is described in the Appendix.
+$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{l}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right].$
+Followers’ Update for Best Response
+To enable follower agents to effectively learn the best response policy $\pi^{f*}$ to the leader’s policy $\pi^{l}$, we employ multiple updates for followers’ shared policy and Q networks in the SMFRL algorithm.
+Each update cycle refines the policy network $\pi_{\theta^{f}}$ for the followers using the policy gradient method:
+$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L^{f}_{t}\sim\mathcal{D}}\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{-},a^{l}_{t},L^{f}_{t})|_{a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})}\Big{]},$
+where the action-value function ${Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})$, parameterized by $\phi^{f}$, is periodically updated by minimizing the mean squared error (MSE) loss:
+$\mathcal{L}(\phi^{f})=$ $\mathbb{E}_{{s}_{t}^{f},{a}_{t}^{f},a^{l}_{t},L^{f}_{t},s_{t+1}^{f}\sim\mathcal{D}}\left[\left(y^{f}_{t}-{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)^{2}\right],$ (4)
+$y^{f}_{t}=$ $r^{f}_{t}+\gamma{Q}_{\phi^{f}_{-}}(s^{f}_{t+1},a^{f}_{t+1},a^{l}_{t+1},L^{f}_{t+1}),$
+with the next leader’s action $a^{l}_{t+1}=\mu_{\phi^{l}_{-}}(s^{l}_{t+1})$, and the next population distribution $L^{f}_{t+1}=L(\mathbf{s}_{t+1}^{f},\mathbf{a}_{t+1}^{f})$.
+The joint actions of followers $\mathbf{a}_{t+1}^{f}=\{a^{f1}_{t+1},\ldots,a^{fN}_{t+1}\}$ are determined with each action $a^{fi}_{t+1}=\mu(s^{fi}_{t+1},a^{l}_{t+1})$.
+The gradient of the loss function $\mathcal{L}(\phi^{f})$ is derived as:
+$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\mathcal{D}}\Big{[}\left(y^{f}_{t}-{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{f}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\Big{]}.$
+The pseudocode for the SMFRL algorithm (1) is presented in Appendix A.
+Table 1: Evaluation of SMFRL and five baselines on 3 metrics for dynamic Stackelberg mean field games with different followers’ numbers.
+Metrics Followers’ Optimal SMFRL MF-MARL MADDPG IDDPG DDPG U.S. Policy
+Number Value +Real +Real
+Leader’s Payoff N=100 \ 3387.36 -76.82 -282.62 -701.61 2840.07 1718.08
+N=1000 \ 3411.16 -668.03 121.18 -767.53 2874.19 -796.4
+Exploitability N=100 0. 0.000107 0.002190 0.007908 0.001333 0.009885 0.032490
+N=1000 0. 0.000158 0.003481 0.035613 0.000565 0.003002 0.010522
+Social Welfare N=100 100 96.87 82.02 55.74 83.09 94.5 79.23
+N=1000 1000 968.94 834.93 645.38 874.53 943.82 334.79
 5 Experiment
-Figure 2: The training curves for 6 algorithms on 4 macroeconomic indicators, comparing settings without behavior cloning as pre-train (N=100 without BC) and with behavior cloning (N=100-BC & N=1000-BC).
-Table 1: Test results for 6 algorithms on 5 indicators under 3 settings (N=100 without BC / N=100-BC / N=1000-BC).
-Algorithms Per Capita GDP Social Welfare Income Gini Wealth Gini Years
-Rule-based 1.41e+05 / 3.66e+11 / 1.41e+05 69.27 / 79.23 / 334.79 0.89 / 0.52 / 0.90 0.92 / 0.53 / 0.93 1.00 / 217.45 / 1.00
-DDPG 1.41e+05 / 2.03e+12 / 4.92e+11 70.91 / 94.50 / 527.09 0.88 / 0.46 / 0.75 0.92 / 0.48 / 0.79 1.00 / 299.85 / 100.68
-MADDPG 4.93e+04 / 6.38e+12 / 6.82e+12 55.74 / 93.89 / 954.88 0.91 / 0.57 / 0.56 0.92 / 0.58 / 0.62 1.00 / 268.53 / 278.50
-IDDPG 1.21e+05 / 7.41e+12 / 2.79e+12 83.09 / 98.16 / 512.19 0.88 / 0.53 / 0.77 0.92 / 0.55 / 0.81 1.00 / 300.00 / 100.68
-MF-MARL 8.66e+07 / 5.44e+12 / 1.13e+05 82.02 / 98.21/ 440.00 0.82 / 0.50 / 0.90 0.83 / 0.52 / 0.93 75.75 / 300.00 / 1.00
-SMFRL 9.59e+12 / 1.01e+13 / 9.68e+12 96.87 / 96.90 / 975.15 0.52 / 0.51 / 0.52 0.51 / 0.53 / 0.51 300.00 / 299.89 / 300.00
-In this section, we conduct two important experiments in the TaxAI environment: (1) We compare the performance of the SMFRL algorithm with 5 baselines in solving SMFGs; (2) We compare the effects of the SMFG method with 5 economic policies, and discuss the efficiency-equity tradeoff and SMFG assumptions.
-5.1 Environment
-TaxAI (Mi et al., 2023) is a reinforcement learning simulator used for studying optimal tax policy.
-The environmental settings include the government promoting economic growth (i.e., GDP) through adjusting tax policy and government spending, and households optimizing labor supply and savings ratio for long-term personal utility.
-It is one of the few simulators for researching optimal macroeconomic policies, hence we conduct experiments on it.
-5.2 Algorithmic Analysis for SMFGs
-Baselines
-We compare popular reinforcement learning algorithms including DDPG (Lillicrap et al., 2015) (only for leader), IDDPG (de Witt et al., 2020), MADDPG (Lowe et al., 2017), Mean Field MARL (Yang et al., 2018), and rule-based policy with our algorithm SMFRL.
-For more details see Appendix B.1 and Table 2.
-Results
-We compared SMFRL with 5 baselines in three different experimental setups: without behavior cloning as pre-training and with 100 followers (marked as N=100 without BC); with BC-based pre-training for follower agents at N=100 and N=1000 (N=100-BC; N=1000-BC).
+In this section, we conduct experiments within the TaxAI environment to answer the following three questions: (1) How effective is our SMFRL algorithm in solving dynamic SMFGs?
+(2) For macroeconomic policy-making issues, how does our method compare to real-world strategies, economic models, and existing AI-based methods?
+(3) Given the likely real-world scenario where some households do not adopt the SMFG household policy learned by our method, does our approach remain effective?
+TaxAI Environment
+TaxAI [41] is one of the rare reinforcement learning environments designed to study optimal macroeconomic policies.
+It simulates dynamic interactions between large-scale households and the government, making it particularly well-suited for experimenting with SMFGs.
+TaxAI models the optimal tax policy problems where the government optimizes tax policies to promote economic growth (i.e.
+GDP) and social equity (i.e.
+Gini).
+In response, households optimize their labor supply and savings strategies to maximize their long-term utility.
+5.1 Solving Dynamic SMFGs
+Metrics and Baselines
+We compare several classic multi-agent reinforcement learning (MARL) algorithms based on three metrics: leader’s payoff, exploitability, and social welfare, as defined in Section 3.3.
+Optimal Value: The ideal value of exploitability is 00, indicating that the follower’s policy achieves the best response to the leader’s policy.
+The upper bound of Social Welfare is $N$, given that household utility in TaxAI cannot exceed 1.
+For Leader’s Payoff, no maximum has been established.
+MF-MARL [61]: Utilizes the DDPG [36] algorithm for the leader and a mean-field MARL approach for the followers.
+MADDPG [38] and IDDPG [17]: Both feature shared actor and critic networks among followers.
+DDPG+Real: DDPG for the leader and behavior cloning [57] from real data for followers.
+U.S.
+policy+Real: U.S.
+federal tax policy for the leader with behavior cloning for the followers.
+Results Analysis
+Table 1 presents the performance metrics for SMFRL and five baselines, alongside their optimal values, averaged over 10 evaluations of models trained with at least five seeds.
+SMFRL notably outperforms the baselines in achieving higher leader’s payoff, demonstrating its effectiveness in optimizing the leader’s objective.
+The DDPG+Real method, using a real-data trained policy, attains a substantial payoff but exhibits a significant gap from the households’ optimal response, as indicated by its higher exploitability.
+In contrast, SMFRL significantly reduces exploitability for both $N=100$ and $N=1000$, suggesting that its policy for households closely approximates the best response to the leader’s strategy, which is crucial for solving Stackelberg mean field games.
+Moreover, SMFRL closely approaches the optimal social welfare value, indicating that its solutions are the most socially optimal among the considered baselines.
+In summary, the SMFRL algorithm effectively learns policies that maximize the leader’s objective while ensuring that followers attain the best response to the leader’s policy, closely approaching the social optimum.
+The training curves are presented in Figure 4, and the test results for macroeconomic indicators are detailed in Table 1 in Appendix D.
+5.2 Macroeconomic Policy-Making
+Table 2: Test results 6 tax policies across five macroeconomic indicators at two numbers of followers, $N=100$ and $N=1000$.
+Households’ Policies Macroeconomic Indicators
+Number Per Capita GDP Social Welfare Income Gini Wealth Gini Years
+$\mathbf{N=100}$ Free Market 1.37e+05 32.97 0.89 0.92 1.10
+Saez Tax 2.34e+12 73.82 0.21 0.38 300.00
+US Federal Tax 4.88e+11 94.19 0.40 0.40 289.55
+AI Economist-BC 4.24e+12 97.24 0.54 0.52 299.55
+AI Economist 1.26e+05 72.81 0.88 0.91 1.00
+SMFG 9.59e+12 96.87 0.52 0.51 300.0
+$\mathbf{N=1000}$ Free Market 1.41e+05 334.79 0.90 0.93 1.00
+Saez Tax 6.35e+11 498.88 0.68 0.73 100.578
+US Federal Tax 1.41e+05 351.17 0.89 0.93 1.00
+AI Economist-BC N/A N/A N/A N/A N/A
+AI Economist N/A N/A N/A N/A N/A
+SMFG 1.10e+13 968.94 0.54 0.53 300.00
+Metrics and Baselines
+In addressing macroeconomic policymaking, we utilize classical macroeconomic indicators from the real world to evaluate the performance of different policies.
+Per Capita GDP indicates the level of economic development, while Income Gini and Wealth Gini measure the inequality in households’ income and wealth, respectively; a lower Gini index suggests greater social equality.
+The Years metric represents the sustainable duration of an economy, capped at 300 years; economies may terminate earlier due to excessive inequality or inadequate productivity to meet the populace’s needs.
+We compare several methods for solving macroeconomic policies: the Free Market [5] policy entails no government intervention; the Saez Tax [50] represents a well-known economic approach, often advocated for specific tax reforms in practice.
+The U.S.
+Federal Tax is the real policy from the OECD data  [^2] https://data.oecd.org/.
+In these approaches, households employ strategies based on behavior cloning from real data.
+Additionally, the AI Economist [62] represents an AI-based policy using a two-level Proximal Policy Optimization (PPO) framework.
+The AI Economist-BC method leverages behavior cloning policy on real-world data for pretraining.
+Results Analysis
+Table 2 presents the average values of macroeconomic indicators across multiple test rounds for different algorithms under varying numbers of households.
+The results reveal that the Saez Tax method achieves the lowest income Gini and wealth Gini when the number of households is small ($N=100$).
+However, our SMFG method yields higher per capita GDP and social welfare.
+At $N=1000$, the advantages of the SMFG algorithm become even more pronounced compared to other methods.
+It secures the longest duration (Years), the highest GDP, and the lowest income Gini and wealth Gini.
+The social welfare derived from the SMFG approach closely approaches the optimal value of 1000.
+This demonstrates that our method can learn the macroeconomic policy that simultaneously optimizes economic growth, social equity, and social welfare, surpassing existing economic approaches like the Saez Tax and AI-based policies (i.e.
+AI Economist).
+We provide the corresponding training curves in Figure 5 in Appendix D.
+5.3 Effectiveness Analysis in Diverse Households
+Figure 3: The left subfigure illustrates a scenario with heterogeneous households adopting various policies.
+The right subfigure presents test results for households’ decisions (consumption, labor), microeconomic indicators (utility, wealth, income), and macroeconomic indicators (GDP) when there are different proportions of SMFG households ($N=100$).
+Problem Description
+The above experiments have already demonstrated the effectiveness of our method in addressing macroeconomic policy-making problems.
+However, when extending SMFG approach to real-world scenarios, where governments and households implement leader and follower policies trained via the SMFG method (referred to as SMFG government and SMFG households), there remain partial households that might not adopt the SMFG follower policy.
+In economic terms, this scenario can be interpreted as the government introducing new policies and providing recommendations to households, some of whom may choose not to follow these recommendations.
+When faced with heterogeneous households that adopt varying strategies, the question arises: is our SMFG method still effective?
+In experiments, to model households that do not adopt the SMFG policy, referred to as Real-data Households, we utilize behavior cloning policy trained by real data.
+Related experimental details are shown in Appendix D.4.
+We set different proportions of SMFG households (0%, 25%, 50%, 75%, and 100%) and evaluate our SMFG method using microeconomic and macroeconomic indicators.
+In Figure 3, each subplot displays average values for two types of households (bars on the left Y-axis) and their aggregate values (dots on the right Y-axis).
+GDP, a macroeconomic indicator, is shown only in aggregate, not for individual households.
+Result Analysis
+From Figure 3, we identify several key findings: (1) SMFG Households outperform Real-data Households across microeconomic indicators such as wealth, income, and utility, which should encourage more households to adopt the SMFG policy; (2) The presence of heterogeneous households does not prevent SMFG Households from achieving near-optimal utility, capped at 1 in the TaxAI environment; (3) The lower labor contribution from Real-data Households leads to a decrease in Per Capita GDP as their proportion increases.
+Thus, the more households adopt the SMFG follower policy, the more effective the SMFG method becomes.
+6 Conclusion
+This paper innovatively models macroeconomic policymaking as a dynamic Stackelberg mean field game, with the government acting as the leader and large-scale households as followers.
+We introduce a data-driven Stackelberg mean field reinforcement learning algorithm that effectively solves SMFGs.
+When applied to macroeconomic policy formulation, our method significantly outperforms existing economic approaches and AI-based policies, especially in scenarios involving large numbers of households.
+Furthermore, we demonstrate the robustness of our approach, showing that it retains effectiveness even when some households do not adopt the SMFG policy.
+In conclusion, this paper contributes an effective approach to modeling and solving optimal macroeconomic policies in the field of AI for economics and AI for social good.
+Appendix A SMFRL algorithm pseudocode
+Algorithm 1 Stackelberg Mean Field Reinforcement Learning (SMFRL)
+  Initialize ${Q}_{\phi^{l}},{Q}_{\phi^{l}_{-}},{Q}_{\phi^{f}},{Q}_{\phi^{f}_{-}},\pi_{\theta^{l}},\pi_{\theta^{l}_{-}},\pi_{\theta^{f}},\pi_{\theta^{f}_{-}}$.
+  for epoch = 1 to M do
+     Receive initial state $\mathbf{s}_{t}=\{s^{l}_{t},s^{f1}_{t},...,s^{fN}_{t}\}$.
+     for  $t=1$ to max-epoch-length do
+        For the leader agent, select action $a^{l}_{t}=\pi_{\theta^{l}}(s^{l}_{t})+\mathcal{N}_{t}$; for each follower agent $i$, select action $a^{fi}=\pi_{\theta^{f}}(s^{fi}_{t},a^{l}_{t})+\mathcal{N}_{t}$ w.r.t.
+the current policy and exploration.
+        Execute actions $\mathbf{a}_{t}=\{a^{l}_{t},a^{f1}_{t},...,a^{fN}_{t}\}$ and observe reward $\mathbf{r}_{t}=\{r_{t}^{l},r^{f1}_{t},...,r^{fN}_{t}\}$ and next state $\mathbf{s}_{t+1}$.
+        Store $(\mathbf{s}_{t},\mathbf{s}_{t+1},\mathbf{a}_{t},\mathbf{r}_{t},L^{f}_{t})$ in the replay buffer $\mathcal{D}$.
+        $\mathbf{s}_{t}\leftarrow\mathbf{s}_{t+1}$.
+        for j =1 to update-cycles do
+           Sample a random minibatch of $N_{b}$ samples from $\mathcal{D}$.
+           For each follower agent, perform inner updates:
+           for k = 1 to inner-update-cycles do
+              Set
+$y^{f}_{t}=r^{f}_{t}+\gamma{Q}_{\phi^{f}_{-}}(s^{f}_{t+1},a^{f}_{t+1},a^{l}_{t+1},L^{f}_{t+1}).$
+              Update follower’s critic by minimizing the loss:
+$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\mathcal{D}}\Big{[}\left(y^{f}_{t}-{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{f}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\Big{]}.$
+              Update follower’s actor by policy gradient:
+$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L^{f}_{t}\sim\mathcal{D}}\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{-},a^{l}_{t},L^{f})|_{a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})}\Big{]}.$
+           end for
+           After follower updates, update the leader’s networks:
+           Set
+$y^{l}_{t}=r_{t}^{l}+\gamma{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L^{f}_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})}.$
+           Update leader’s critic net by minimizing the loss:
+$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{l}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right]$
+           Update leader’s actor by sampled policy gradient:
+$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{-},L^{f}_{t})|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}\right].$
+        end for
+        Update target network parameters:
+$\begin{gathered}\phi_{-}^{l}\leftarrow\tau_{\phi}\phi^{l}+\left(1-\tau_{\phi}\right)\phi_{-}^{l}\\ \theta_{-}^{l}\leftarrow\tau_{\theta}\theta^{l}+\left(1-\tau_{\theta}\right)\theta_{-}^{l}\\ \phi_{-}^{f}\leftarrow\tau_{\phi}\phi^{f}+\left(1-\tau_{\phi}\right)\phi_{-}^{f}\\ \theta_{-}^{f}\leftarrow\tau_{\theta}\theta^{f}+\left(1-\tau_{\theta}\right)\theta_{-}^{f}\end{gathered}$
+     end for
+  end for
+Appendix B Assumptions and Limitations
+Assumptions
+This paper models the problem of macroeconomic policy-making as a Dynamic Stackelberg Mean Field Game, based on the following assumptions: (1) Homogeneous followers: We assume that a large-scale group of households is homogeneous.
+They can use different characteristics as observations to influence decisions, but there are commonalities in human behavioral strategies.
+(2) Rational Expectations: We assume that both macro and micro agents engage in rational decision-making, adjusting their future expectations based on observed information.
+However, in reality, the level of rationality varies among different households.
+Most households exhibit bounded rationality, and their expectations and preferences differ accordingly.
+(3) Experimental environment: We validate our approach through experiments in the TaxAI environment, based on the assumption that results within this environment can provide insights applicable to real-world scenarios.
+Addressing and potentially relaxing these assumptions will be a primary focus of our future research.
+Limitations
+The limitations of our SMFG method will be thoroughly investigated in future work: (1) We plan to consider dynamic games involving multiple leaders and large-scale followers to explore policy coordination across various macroeconomic sectors.
+(2) We will continue to develop theoretical proofs for the equilibrium solutions in Stackelberg mean field games.
+Currently, our approach is empirically demonstrated by showing that followers converge toward their best responses and that the leader achieves higher performance compared to other baselines.
+(3) We intend to examine dynamic games between a leader and a large, heterogeneous group of followers, including scenarios where followers dynamically alter their strategies, to determine the optimal leader policy.
+Addressing these limitations will provide further insights applicable to real-world scenarios.
+Appendix C Impact Statements
+This paper presents work whose goal is to advance the fields of AI for social science and AI for economics.
+Our work aims to offer suggestions and references for governments and the people, yet it must not be rashly applied to the real world.
+There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
+Appendix D Additional Results
+Compute Resources
+All experiments are run on 2 workstations: A 64-bit server with dual AMD EPYC 7742 64-Core Processors @2.25 GHz, 256 cores, 512 threads, 503GB RAM, and 2 NVIDIA A100-PCIE-40GB GPU.
+A 64-bit workstation with Intel Core i9-10920X CPU @ 3.50GHz, 24 cores, 48 threads, 125 GB RAM, and 2 NVIDIA RTX2080 Ti GPUs.
+The following Table 3 shows the approximate training times for several algorithms.
+Table 3: the approximate training times for several algorithms for solving dynamic SMFGs.
+Algorithm N=100 (hours) N=1000 (hours)
+SMFRL 4 14
+MADDPG 3 16
+MF-MARL 4 9
+IDDPG 2 6
+DDPG+Real 1 5
+Rule-based+Real 0.25 2
+D.1 Training Curves for Dynamic SMFGs
+Due to the challenging nature of the task, the baselines we compared did not perform well in solving dynamic SMFGs.
+Therefore, we utilized behavior cloning as a pre-training strategy for the follower agents.
+Therefore, we compared SMFRL with 5 baselines in 4 different experimental setups: without behavior cloning as pre-training for follower agents at N=100 and N=1000 (marked as N=100 without BC and N=1000 without BC); with BC-based pre-training for follower agents at N=100 and N=1000 (N=100-BC; N=1000-BC).
 Figure 4 illustrates the training curves of 4 key macroeconomic indicators under these three settings.
+The solid line represents the average value of the metrics across the 5 random seeds, while the shaded area represents the standard deviation.
 Each row corresponds to one setting, and each column to a macroeconomic indicator, including per capita GDP, social welfare, income Gini, and wealth Gini.
@@
 Besides, in N=100 without BC and N=1000-BC, SMFRL consistently obtains the most optimal solutions across all indicators.
-Figure 3: The training curves for 6 tax policies on 5 macroeconomic indicators (N=100 & N=1000 with BC)
-5.3 Comparison of Macroeconomic Policies
-Baselines
-We compare free market policy, Saez tax (Saez, 2001), U.S.
-Federal tax  [^2] https://data.oecd.org/, AI Economist (Zheng et al., ), and AI Economist-BC.
-Saez tax will provide suggestions for real-world tax reform.
-For more details see Appendix B.1.
-Policies Results
+Table 4: Performance metrics of six algorithms across different settings on key economic indicators.
+Settings Algorithm GDP per Capita Social Welfare Income Gini Wealth Gini Years
+N=100 without BC Rule-based+Random 1.41e+05 69.27 0.89 0.92 1.00
+DDPG+Random 1.41e+05 70.91 0.88 0.92 1.00
+MADDPG 4.93e+04 55.74 0.91 0.92 1.00
+IDDPG 1.21e+05 83.09 0.88 0.92 1.00
+MF-MARL 8.66e+07 82.02 0.82 0.83 75.75
+SMFRL 9.59e+12 96.87 0.52 0.51 300.00
+N=1000 without BC Rule-based+Random 1.42e+05 705.85 0.90 0.93 1.00
+DDPG+Random 1.42e+05 654.37 0.90 0.93 1.00
+MADDPG 3.37e+09 645.38 0.83 0.85 63.58
+IDDPG 1.33e+05 874.53 0.89 0.92 1.00
+MF-MARL 1.31e+05 834.93 0.88 0.92 1.50
+SMFRL 1.10e+13 968.94 0.54 0.53 300.00
+N=100 with BC Rule-based+Real 3.66e+11 79.23 0.52 0.53 217.45
+DDPG+Real 2.03e+12 94.50 0.46 0.48 299.85
+MADDPG 6.38e+12 93.89 0.57 0.58 268.53
+IDDPG 7.41e+12 98.16 0.53 0.55 300.00
+MF-MARL 5.44e+12 98.21 0.50 0.52 300.00
+SMFRL 1.01e+13 96.90 0.51 0.53 299.89
+N=1000 with BC Rule-based+Real 1.41e+05 334.79 0.90 0.93 1.00
+DDPG+Real 4.92e+11 527.09 0.75 0.79 100.68
+MADDPG 6.82e+12 954.88 0.56 0.62 278.50
+IDDPG 2.79e+12 512.19 0.77 0.81 100.68
+MF-MARL 1.13e+05 440.00 0.90 0.93 1.00
+SMFRL 9.68e+12 975.15 0.52 0.51 300.00
+Figure 4: The training curves for 6 algorithms on 4 macroeconomic indicators, comparing settings without behavior cloning as pre-train (N=100 & N=1000) and with behavior cloning (N=100-BC & N=1000-BC).
+D.2 Training Curves for Various Tax Policies
 We compare the performance of 6 policies across four economic indicators under two settings: with N=100 and N=1000 households.
@@
 The Saez tax also reduces the Gini index, but not as effectively or stably as the SMFG.
+Figure 5: The training curves for 6 tax policies on 5 macroeconomic indicators (N=100 & N=1000 with BC)
+D.3 Efficiency-Equity Tradeoff of Policies
 Figure 6: Comparative performance of various policies under multi-objective assessment (Efficiency-Equity).
-Figure 5: Comparison of micro and macroeconomic indicators at varying proportions of SMFG households, demonstrating the positive impact on the individual and overall economy of SMFG households.
-Efficiency-Equity Tradeoff
 In economics, the Efficiency-Equity Tradeoff is a highly debated issue.
@@
 However, SMFG consistently proves to be the most effective policy under a wide range of preference settings.
-SMFG Assumption Analysis
-The above experiments have already demonstrated the superiority of the SMFG method.
-However, the modeling and solving of SMFG rely on an important assumption: the optimal decisions of governments and households depend on the overall state information.
-In reality, it is natural for the government to focus on market information in decision-making, but it cannot be guaranteed that all households will consider market information in their decisions.
-To discuss the applicability of the SMFG method, we introduce two types of households within the same economy: those based on real data (Real-data Households) and those that focus on market information (SMFG households).
-We simulate the economic operations at various proportions of SMFG households: 0%, 25%, 50%, 75%, and 100%.
-Figure 5 displays the microeconomic indicators and macroeconomic GDP, with each subplot illustrating the average values of different economic indicators for different households (bars in left Y-axis) and their aggregate values (dots in right Y-axis).
-GDP, being a macroeconomic indicator, did not have corresponding values plotted for households.
-The results indicate that, across all proportions, the microeconomic indicators of SMFG households, such as utility, wealth, and income, are consistently higher than Real-data households.
-As the proportion of SMFG households increases, there is an upward trend in economic indicators.
-This suggests that the presence of SMFG households positively impacts both individual and overall economic development, and our SMFG method also works when some households do not meet the assumption of focusing on market information.
-Therefore, governments should encourage households to pay more attention to market information in decision-making.
-6 Conclusion
-This paper innovatively models the optimal macroeconomic policy problem based on the Stackelberg mean field game, and proposes a solution method involving pre-training based on real data and a model-free Stackelberg mean field reinforcement learning algorithm.
-Our method addresses the limitations of existing approaches, including dynamic interactions and feedback between governments and households, interactions among large populations of households, and solving complex tasks without relying on environmental transitions.
-Experimental results have showcased the performance of the SMFG method.
-In conclusion, this paper contributes an effective approach to modeling and solving optimal macroeconomic policies in the field of AI for economics and AI for social good.
-Impact Statements
-This paper presents work whose goal is to advance the fields of AI for social good and AI for economics.
-Our work aims to offer suggestions and references for governments and the people, yet it must not be rashly applied to the real world.
-There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
-Appendix A Economic Model Details
-Economic activities among households aggregate into labor markets, capital markets, goods markets, etc.
-In the labor market, households are the providers of labor, with the aggregate supply $S(W_{t})=\sum_{i}^{N}e_{t}^{i}h_{t}^{i}$, and firms are the demanders of labor, with the aggregate demand $D(W_{t})=\mathcal{L}_{t}$.
-When supply equals demand in the labor market, there exists an equilibrium price $W_{t}^{*}$ that satisfies:
-$S(W^{*}_{t})=D(W^{*}_{t}),\mathcal{L}_{t}=\sum^{N}_{i}e_{t}^{i}h_{t}^{i}.$
-In the capital market, financial intermediaries play a crucial role, lending the total deposits of households $A_{t+1}=\sum_{i}^{N}a_{t+1}$ to firms as production capital $K_{t+1}$, and purchasing government bonds $B_{t+1}$ at the interest rate $r_{t}$.
-The capital market clears when supply equals demand:
-$K_{t+1}+B_{t+1}-A_{t+1}=\left(r_{t}+1\right)(K_{t}+B_{t}-A_{t})$
-In the goods market, firms produce and supply goods, while all households, the government, and physical capital investments $X_{t}$ demand them. The goods market clears when:
-$Y_{t}=C_{t}+G_{t}+X_{t}$
-where $C_{t}=\sum_{i}^{N}c^{i}_{t}$ represents the total consumption of consumers, and $G_{t}$ is government spending. The supply, demand, and price represent the states of the market.
-Algorithm 1 Stackelberg Mean Field Reinforcement Learning (SMFRL)
-  Initialize $\tilde{Q}_{\phi^{l}},\tilde{Q}_{\phi^{l}_{-}},\tilde{Q}_{\phi^{f}},\tilde{Q}_{\phi^{f}_{-}},\pi_{\theta^{l}},\pi_{\theta^{l}_{-}},\pi_{\theta^{f}},\pi_{\theta^{f}_{-}}$.
-  for epoch = 1 to M do
-     Recieve initial observation $\mathbf{s}_{t}=\{s^{l}_{t},s^{f1}_{t},...,s^{fN}_{t}\}$.
-     for  $t=1$ to max-epoch-length do
-        For leader agent, select action $a^{l}_{t}=\pi_{\theta^{l}}(s^{l}_{t})+\mathcal{N}_{t}$; for each follower agent $i$, select action $a^{fi}=\pi_{\theta^{f}}(s^{fi}_{t},a^{l}_{t})+\mathcal{N}_{t}$ w.r.t.
-the current policy and exploration.
-        Execute actions $\mathbf{a}_{t}=\{a^{l}_{t},a^{f1}_{t},...,a^{fN}_{t}\}$ and observe reward $\mathbf{r}_{t}=\{r_{t}^{l},r^{f1}_{t},...,r^{fN}_{t}\}$ and next observation $\mathbf{s}_{t+1}$.
-        Store $(\mathbf{s}_{t},\mathbf{s}_{t+1},\mathbf{a}_{t},\mathbf{r}_{t},L^{f}_{t})$ in the replay buffer $\mathcal{D}$.
-        $\mathbf{s}_{t}\leftarrow\mathbf{s}_{t+1}$.
-        for j =1 to update-cycles do
-           Sample a random minibatch of $N_{b}$ samples from $\mathcal{D}$.
-           For the leader agent, set
-$y^{l}_{t}=r_{t}^{l}+\gamma\tilde{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L^{f}_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})}.$
-           Update leader’s critic net by minimizing the loss:
-$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{l}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right]$
-           Update leader’s actor by sampled policy gradient:
-$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{-},L^{f}_{t})|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}\right].$
-           For the representative follower agent, set
-$y^{f}_{t}=r^{f}_{t}+\gamma\tilde{Q}_{\phi^{f}_{-}}(s^{f}_{t+1},a^{f}_{t+1},a^{l}_{t+1},L^{f}_{t+1}).$
-           Update follower’s critic by minimizing the loss:
-$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\mathcal{D}}\Big{[}\left(y^{f}_{t}-\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{f}}\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\Big{]}.$
-           Update follower’s actor by policy gradient:
-$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L^{f}_{t}\sim\mathcal{D}}\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{-},a^{l}_{t},L^{f})|_{a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})}\Big{]}.$
-        end for
-        Update target network parameters for each agent with learning rates $\tau_{\phi}$ and $\tau_{\theta}$:
-$\begin{gathered}\phi_{-}^{l}\leftarrow\tau_{\phi}\phi^{l}+\left(1-\tau_{\phi}\right)\phi_{-}^{l}\\ \theta_{-}^{l}\leftarrow\tau_{\theta}\theta^{l}+\left(1-\tau_{\theta}\right)\theta_{-}^{l}\\ \phi_{-}^{f}\leftarrow\tau_{\phi}\phi^{f}+\left(1-\tau_{\phi}\right)\phi_{-}^{f}\\ \theta_{-}^{f}\leftarrow\tau_{\theta}\theta^{f}+\left(1-\tau_{\theta}\right)\theta_{-}^{f}\end{gathered}$
-     end for
-  end for
-Appendix B Additional Results
-B.1 Baselines
-Baselines for solving SMFGs
-Current research on model-based methods for SMFG is limited by strong assumptions and simplifications, rendering them ineffective for complex problem-solving.
-However, there is a lack of model-free algorithms for solving SMFG in continuous decision spaces.
-Therefore, this study opts to compare with popular reinforcement learning algorithms, including DDPG, MADDPG, IDDPG, and mean field MARL.
-Table 2 provides detailed information on the algorithms used by the leader and follower agent in these baselines.
-Table 2: The details of the algorithms used by the leader and follower agents in the baselines for solving SMFG.
-Baselines Follower’s Algorithm Leader’s Algorithm
-Rule-based Random/Behavior cloning Rule-based/Free market
-DDPG Random/Behavior cloning DDPG
-MADDPG MADDPG MADDPG
-IDDPG IDDPG IDDPG
-MF-MARL Mean Field MARL DDPG
-SMFRL SMFRL SMFRL
-Baselines for economic policies
-We compare 5 economic policies with our method SMFG:
-(1) Free Market: A market without policy intervention.
-(2) Saez Tax (Saez, 2001): The Saex tax policy is often considered a suggestion for specific tax reforms in the real world. Details are shown in Appendix C.
-(3) U.S. Federal Tax: Real data from OECD  [^3] https://data.oecd.org/ in 2022 for this policy.
-(4) AI Economist (Zheng et al., ): This is a two-level MARL method based on Proximal Policy Optimization (PPO).
-In the first phase, households’ policies are trained from scratch in a free-market (no-tax) environment.
-In the second phase, households continue to learn under an RL social planner.
-(5) AI Economist-BC: For fairness in comparison, we evaluated the AI Economist method with behavior cloning as pre-training to determine its effectiveness.
-B.2 Efficiency-Equity Tradeoff
-In Figure 6, we also present the efficiency (per capita GDP) and equity (wealth Gini) corresponding to various algorithms for solving SMFGs.
-These methods result in similar Gini indices, with SMFRL achieving the highest per capita GDP.
-Apart from a strong emphasis on societal fairness, SMFRL emerges as the algorithm that can best strike a balance between efficiency and equity.
-Figure 6: Comparison of macroeconomic indicators under different algorithms for solving SMFGs.
-Table 3: Test results for 6 tax policies on 5 macroeconomic indicators ($N=100$ $N=1000$).
-Algorithms $N=100$ $N=1000$
-Per Capita Social Income Wealth Years Per Capita Social Income Wealth Years
-GDP Welfare Gini Gini GDP Welfare Gini Gini
-Free Market 1.37e+05 32.97 0.89 0.92 1.10 1.41e+05 334.79 0.90 0.93 1.00
-Saez Tax 2.34e+12 73.82 0.21 0.38 300.00 6.35e+11 498.88 0.68 0.73 100.578
-US Federal Tax 4.88e+11 94.19 0.40 0.40 289.55 1.41e+05 351.17 0.89 0.93 1.00
-AI Economist-BC 4.24e+12 97.24 0.54 0.52 299.55 N/A N/A N/A N/A N/A
-AI Economist 1.26e+05 72.81 0.88 0.91 1.00 N/A N/A N/A N/A N/A
-SMFG 1.01e+13 96.90 0.51 0.53 299.89 9.68e+12 975.15 0.52 0.51 300.00
-B.3 Behavior Cloning Experiments
+D.4 Behavior Cloning Experiments
+We conduct behavior cloning based on real data to simulate the behavior strategies of households in realistic scenarios, which are then used in Experiment 5.3 to compare with SMFG Households.
+We use the statistical data from the 2022 Survey of Consumer Finances (SCF)  [^3] https://www.federalreserve.gov/econres/scfindex.htm.
+Details about the behavior cloning experiments are provided below.
 This experiment conducts behavior cloning on networks for four different household policies: Multilayer Perceptron (MLP), AI economist’s network (MLP+LSTM+MLP), MF-MARL, and MADDPG network.
-The first two, as their network outputs are probability distributions, use negative log-likelihood loss (Figure 7 left); the latter two’s networks employ deterministic policies, hence they use mean square error loss against real data (Figure 7 right).
+The first two, as their network outputs, are probability distributions, use negative log-likelihood loss (Figure 7 left); the latter two’s networks employ deterministic policies, hence they use mean square error loss against real data (Figure 7 right).
 The loss convergence curve of behavior cloning is shown in Figure 7.
@@
 Figure 7: The behavior cloning loss for 4 networks in two loss types.
-Appendix C Saez tax
+Appendix E Saez tax
 The Saex tax policy is often considered a suggestion for specific tax reforms in the real world.
-The specific calculation method is as follows (Saez, 2001).
+The specific calculation method is as follows [50].
 The Saez tax utilizes income distribution $f(z)$ and cumulative distribution $F(z)$ to get the tax rates.
@@
 where $Z=\sum_{i}z_{i}$ when tax rates is $\tau$.
-Appendix D Theoretical Proof
-Proposition D.1.
-The action value functions $Q_{i}$ and $\tilde{Q}_{t}$ are equivalent, i.e.
-$Q_{i}(\mathbf{s}_{t},\mathbf{a}_{t})=\tilde{Q}_{t}(s_{t}^{l},a_{t}^{l},L_{t}^{f})$, $\forall\mathbf{s}_{t}\in\mathcal{S}^{l}\times\{\mathcal{S}^{f}\}^{N}$, $\forall\mathbf{a}_{t}\in\mathcal{A}^{l}\times\{\mathcal{A}^{f}\}^{N}$, $\forall L_{t}^{f}\in\mathbb{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$.
+Appendix F Theoretical Proof
+Proposition F.1.
+The action value functions $Q_{i}$ and ${Q}_{t}$ are equivalent, i.e.
+$Q_{i}(\mathbf{s}_{t},\mathbf{a}_{t})={Q}_{t}(s_{t}^{l},a_{t}^{l},L_{t}^{f})$, $\forall\mathbf{s}_{t}\in\mathcal{S}^{l}\times\{\mathcal{S}^{f}\}^{N}$, $\forall\mathbf{a}_{t}\in\mathcal{A}^{l}\times\{\mathcal{A}^{f}\}^{N}$, $\forall L_{t}^{f}\in\mathbb{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$.
 Proof.
@@
 The empirical distribution of the followers’ state-action pairs, denoted as $L:\mathcal{V}\rightarrow\Delta(\mathcal{V})$, is expressed as $L^{f}_{t}(s,a)=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}_{\{(\tilde{s}_{i},\tilde{a}_{i})=(\tilde{s},\tilde{a})\}}$.
-The bijectiveness of the mapping $L$, ensured by the sorting process, leads to the conclusion that $Q_{t}(\mathbf{s}_{t},\mathbf{a}_{t})=Q_{t}(s^{l}_{t},a^{l}_{t},\mathbf{v})=Q_{t}(s^{l}_{t},a^{l}_{t},\tilde{\mathbf{v}})=\tilde{Q}_{t}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$.
+The bijectiveness of the mapping $L$, ensured by the sorting process, leads to the conclusion that $Q_{t}(\mathbf{s}_{t},\mathbf{a}_{t})=Q_{t}(s^{l}_{t},a^{l}_{t},\mathbf{v})=Q_{t}(s^{l}_{t},a^{l}_{t},\tilde{\mathbf{v}})={Q}_{t}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$.
 ∎
+Appendix G Economic Model Details
+Economic activities among households aggregate into labor markets, capital markets, goods markets, etc.
+In the labor market, households are the providers of labor, with the aggregate supply $S(W_{t})=\sum_{i}^{N}e_{t}^{i}h_{t}^{i}$, and firms are the demanders of labor, with the aggregate demand $D(W_{t})=\mathcal{L}_{t}$.
+When supply equals demand in the labor market, there exists an equilibrium price $W_{t}^{*}$ that satisfies:
+$S(W^{*}_{t})=D(W^{*}_{t}),\mathcal{L}_{t}=\sum^{N}_{i}e_{t}^{i}h_{t}^{i}.$
+In the capital market, financial intermediaries play a crucial role, lending the total deposits of households $A_{t+1}=\sum_{i}^{N}a_{t+1}$ to firms as production capital $K_{t+1}$, and purchasing government bonds $B_{t+1}$ at the interest rate $r_{t}$.
+The capital market clears when supply equals demand:
+$K_{t+1}+B_{t+1}-A_{t+1}=\left(r_{t}+1\right)(K_{t}+B_{t}-A_{t})$
+In the goods market, firms produce and supply goods, while all households, the government, and physical capital investments $X_{t}$ demand them. The goods market clears when:
+$Y_{t}=C_{t}+G_{t}+X_{t}$
+where $C_{t}=\sum_{i}^{N}c^{i}_{t}$ represents the total consumption of consumers, and $G_{t}$ is government spending. The supply, demand, and price represent the states of the market.

### v3

@@
-Learning Macroeconomic Policies based on Microfoundations: A Dynamic Stackelberg Mean Field Game Approach
+Learning Macroeconomic Policies based on Microfoundations: A Stackelberg Mean Field Game Approach
 Abstract
-The Lucas critique emphasizes the importance of considering the impact of policy changes on the expectations of micro-level agents in macroeconomic policymaking.
-However, the inherently self-interested nature of large-scale micro-agents, who pursue long-term benefits, complicates the formulation of optimal macroeconomic policies.
-This paper proposes a novel general framework named Dynamic Stackelberg Mean Field Games (Dynamic SMFG) to model such policymaking within sequential decision-making processes, with the government as the leader and households as dynamic followers.
-Dynamic SMFGs capture the dynamic interactions among large-scale households and their response to macroeconomic policy changes.
-To solve dynamic SMFGs, we propose the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm, which leverages the population distribution of followers to represent high-dimensional joint state and action spaces.
-In experiments, our method surpasses macroeconomic policies in the real world, existing AI-based and economic methods.
-It allows the leader to approach the social optimum with the highest performance, while large-scale followers converge toward their best response to the leader’s policy.
-Besides, we demonstrate that our approach retains effectiveness even when some households do not adopt the SMFG policy.
-In summary, this paper contributes to the field of AI for economics by offering an effective tool for modeling and solving macroeconomic policy-making issues.
+The Lucas critique emphasizes the importance of considering microfoundations—how micro-agents (i.e., households) respond to policy changes—in macroeconomic policymaking.
+However, due to the vast scale and complex dynamics among micro-agents, predicting microfoundations is challenging.
+Consequently, this paper introduces a Stackelberg Mean Field Game (SMFG) approach that models macroeconomic policymaking based on microfoundations, with the government as the leader and micro-agents as dynamic followers.
+This approach treats large-scale micro-agents as a population, to optimize macroeconomic policies by learning the dynamic response of this micro-population.
+Our experimental results indicate that the SMFG approach outperforms real-world macroeconomic policies, existing AI-based and economic methods, enabling the learned macroeconomic policy to achieve the highest performance while guiding large-scale micro-agents toward maximal social welfare.
+Additionally, when extended to real-world scenarios, households that do not adopt the SMFG policy experience lower utility and wealth than adopters, thereby increasing the attractiveness of our policy.
+In summary, this paper contributes to the field of AI for economics by offering an effective tool for modeling and solving macroeconomic policymaking issues.
+Figure 1: The mapping from macroeconomic policymaking (left) to Dynamic Stackelberg Mean Field Games (right).
+On the left, macro-level government actions, including fiscal or monetary policies, aim to optimize GDP, welfare, or equity.
+The microfoundations reflect the economic behaviors of large-scale micro-agents responding to shifts in macroeconomic policy.
+In Dynamic SMFGs, the government acts as the leader, and micro-agents as followers, respond to the leader’s policy.
 1 Introduction
-Macroeconomic policymaking is fundamental to the sustained development of an economy [52, 46].
-Governments influence economic production, wealth distribution, social stability, and welfare through economic policies such as interest rates, taxation, and fiscal spending [49].
-Therefore, accurately modeling and solving optimal macroeconomic policies and simulating the effects of their implementation are critical issues.
-According to the Lucas critique [39], predicting the impact of macroeconomic policies effectively requires modeling individual micro-level behavior—termed "microfoundations"—to forecast personal responses to policy changes and aggregate these into macroeconomic effects.
-Nonetheless, several challenges persist: dynamic interactions among micro-agents, between micro-agents and macro-authorities, the inherently selfish behavior of agents focused on optimizing long-term returns, and the curse of dimensionality due to the large-scale nature of agents such as households and firms.
-This paper introduces a novel framework termed Dynamic Stackelberg Mean Field Game (Dynamic SMFG) designed to address complex macroeconomic policy-making problems.
-Within this dynamic setting, the government, serving as the leader, adjusts policies iteratively, while a large population of households, acting as followers, respond accordingly.
-This process involves sequential decision-making, with both the leader and the followers aiming to optimize their cumulative long-term benefits.
-Given the leader’s policy, the SMFG reduces to a mean field game, which is widely used for modeling interactions among large-scale homogenous agents [33].
-In Dynamic SMFGs, followers aim to maximize their long-term benefits by providing the best response to the leader’s policy, while the leader’s goal is to find an optimal policy that effectively guides the massive followers toward achieving comprehensive macroeconomic objectives.
-Existing approaches often simplify Dynamic SMFGs to single-step decisions [45] or assume myopic followers [63], which do not adequately capture the complexities of macroeconomic policy-making.
-Therefore, we propose a data-driven reinforcement learning algorithm, Stackelberg Mean-Field Reinforcement Learning (SMFRL), to solve Dynamic SMFGs without requiring transition information or prior knowledge.
-The SMFRL algorithm utilizes a mean field approximation [28] to reduce the high-dimensional joint state and actions of followers to a manageable population distribution, allowing for the effective learning of a central critic Q function based on the overall followers’ distribution.
-Experiments conducted on TaxAI [41], an environment designed for studying optimal tax policies, first validate the performance of SMFRL in solving Dynamic SMFGs.
-The results indicate that SMFRL enables followers to closely achieve their best response while allowing the leader to approach the social optimum with the highest performance compared with baselines.
-Furthermore, the paper compares the Dynamic SMFG method with macroeconomic policies such as AI Economist [62], the Saez tax [50], the 2022 U.S.
-Federal tax, and free market policy [5], demonstrating the superiority of the Dynamic SMFG framework in macroeconomic policymaking.
-Considering real-world applications, our experiments show that even if a portion of households do not adopt the follower policy learned by the SMFG method, it does not adversely affect the utility, wealth, and income of households that do adopt the SMFG policy, which are in fact higher.
-However, the GDP of the economy may decrease with an increase in the number of households not supporting the SMFG policy, as they typically contribute less labor.
-In summary, the contributions of this paper are threefold:
+Macroeconomic policymaking is fundamental to the sustained development of an economy (Schneider and Frey 1988; Persson and Tabellini 1999).
+Governments influence economic production, wealth distribution, social stability, and welfare through economic policies such as interest rates, taxation, and fiscal spending (Sachs and Warner 1995).
+Therefore, accurately modeling and solving optimal macroeconomic policies, and evaluating the effects of their implementation are critical issues.
+According to the Lucas critique (Lucas Jr 1976), predicting the impact of macroeconomic policies effectively requires modeling individual micro-level behaviors—termed “microfoundations”—to forecast micro-level responses to policy shifts and aggregate these into macroeconomic effects.
+Nonetheless, the vast number of micro-agents (i.e., households) and the complex dynamic interactions among them render predictions based on microfoundations exceedingly challenging.
+To address the challenges outlined, this paper introduces a Stackelberg Mean Field Game (SMFG) approach: initially proposing a Dynamic Stackelberg Mean Field Game (Dynamic SMFG) framework to model macroeconomic policymaking based on microfoundations (see Figure 1).
+Within this framework, we design a model-free reinforcement learning algorithm, Stackelberg Mean-Field Reinforcement Learning (SMFRL), aimed at optimizing macroeconomic policies by learning the dynamic responses of a large-scale population of micro-agents.
+In the Dynamic SMFG framework, the government acts as the leader, iteratively adjusting macroeconomic policies, while a vast array of micro-agents, serving as followers, react accordingly.
+This process entails sequential decision-making, with both the leader and the followers striving to optimize their cumulative long-term benefits.
+Given the leader’s policy, the SMFG is reduced to a mean field game, treating the large-scale micro-agents as a population to study their dynamic response to policy shifts.
+This dynamic SMFG framework responds to the Lucas critique.
+Existing models often simplify Dynamic SMFGs into single-step decisions (Pawlick and Zhu 2017), repeated games (Alcantara-Jimenez and Clempner 2020), or scenarios with myopic followers (Zhong et al.
+2021), which fail to capture the complexities of macroeconomic policymaking adequately.
+Therefore, we introduce the SMFRL algorithm to optimize long-term utility for both the macro-agent and numerous micro-agents under the dynamic SMFG framework.
+In the SMFRL algorithm, we utilize mean field approximation to reduce the dimensionality of the micro-agents’ high-dimensional joint states and actions and employ a leader-follower update to learn the micro-population’s best response to macroeconomic policy adjustments, as well as to optimize macroeconomic policies considering dynamic feedback from the micro-population.
+This learning method enhances the stability and convergence of policies without requiring transition information or prior knowledge of the environment.
+Our experiments, conducted on the TaxAI (Mi et al.
+2023) environment, firstly demonstrate SMFG approach’s superiority over AI Economist (Zheng et al.
+), the Saez tax (Saez 2001), the 2022 U.S.
+Federal tax, and free market policy (Backhouse 2005) from perspectives of macroeconomic policy evaluation and dynamic responses to economic shocks.
+Further, through ablation studies, we verify the significant roles of mean field approximation and leader-follower update in terms of macroeconomic policy effects, convergence, and maximal social welfare.
+When applied in real-world scenarios, the SMFG approach shows that households refusing the policy have lower wealth and utility than adopters, which enhances the SMFG policy’s attractiveness.
+In summary, this paper’s contributions are threefold:
+1. We introduce the Dynamic Stackelberg Mean Field Games framework, a novel approach for modeling macroeconomic policymaking based on microfoundations, responding to the Lucas critique.
+2.
+We propose a model-free Stackelberg Mean Field Reinforcement Learning algorithm, designed to optimize macroeconomic policies by learning the dynamic responses of large-scale micro-agents within the Dynamic SMFG framework.
+3. Experimental results demonstrate the superiority of our method in modeling and solving macroeconomic policymaking, and further explore its effectiveness in real-world settings.
+Our code is available in an anonymous GitHub repository
+https://anonymous.4open.science/r/SMFG˙macro˙7740
+2 Related work
+Economic Methods
+In the field of economics, research on economic policy typically relies on theoretical model studies, empirical analysis, and econometric methods.
+Classic theoretical models, such as the IS-LM model (Hicks 1980; Gali 1992), are applicable for analyzing short-term policy effects but overlook long-term factors and price fluctuations.
+The aggregate demand-aggregate supply (AD-AS) model (Dutt 2006; Lee, Pesaran, and Smith 1997) focuses on the relationship between aggregate demand and supply, integrating short-term and long-term factors, but simplifies macroeconomic dynamics.
+The Solow growth model (Brock and Taylor 2010) provides a framework for economic growth but does not consider market imperfections and externalities.
+New Keynesian models (Blanchard and Galí 2007; Gabaix 2020) emphasize the stickiness of prices and wages, suitable for explaining economic fluctuations and policy interventions, but are based on strong assumptions.
+DSGE models (An and Schorfheide 2007; Smets and Wouters 2007), built on micro-foundations for macroeconomic forecasting, offer a consistent framework for policy analysis but are limited in handling nonlinearities and market imperfections.
+Empirical analysis (Ramesh et al.
+2010; Vedung 2017) primarily evaluates policy effects through historical data but faces limitations in data timeliness and accuracy.
+Econometric methods (Johnston and DiNardo 1963; Davidson, MacKinnon et al.
+2004), such as regression and time series analysis, can quantify policy effects but are constrained by model settings and data quality.
+Additionally, the Saez tax model (Saez 2001) uses earnings elasticities to optimize tax rates, enhancing social welfare without deterring economic activity.
+This model continues to guide current fiscal policy through effective taxation strategies.
+AI for Economics
+Artificial intelligence (AI), such as reinforcement learning, may offer new perspectives for solving complex economic problems (Tilbury 2022).
+In terms of macroeconomics, AI Economist (Zheng et al.
+) employs curriculum learning on tax policy design, (Trott et al.
+2021) studies the collaboration between central and local governments under COVID-19, as well as monetary policy (Hinterlang and Tänzer ; Chen et al.
+2023), international trade (Sch 2021), market pricing (Danassis et al.
+2023).
+However, these studies consider the government as a single agent, neglecting the dynamic response of the households to policies.
+(Koster et al.
+) studies democratic AI by human voting models, (Yaman et al.
+2023) examines the impact of social sanction rules on labor division.
+However, these works are based on simplified settings and remain a gap in real-world economic policy.
+In terms of microeconomics, (Shi 2021; Rui and Shi 2022; Atashbar and Aruhan Shi 2023) study the optimal saving and consumption problems of micro-agents, and other researchers explore rational expectation equilibrium (Kuriksha 2021; Hill, Bardoscia, and Turrell 2021), multiple equilibria under real-business-cycle (Curry et al.
+2022), the emergence of barter behavior (Johanson et al.
+2022), optimal asset allocation and savings strategies (Ozhamaratli and Barucca 2022).
+Stackelberg Mean Field Game
+Stackelberg Mean Field Games (SMFGs) model the dynamic interactions between a leader and numerous homogeneous followers, enabling the development of macroeconomic policies grounded in microfoundations.
+The methods for SMFGs are divided into model-based and model-free algorithms.
+Model-based methods, such as (Bensoussan, Chau, and Yam 2015; Fu and Horst 2020; Dayanikli and Lauriere 2023; Bergault, Cardaliaguet, and Rainer 2023) employ forward-backward stochastic functional differential equations to ascertain the followers’ mean field equilibrium prior to computing the optimal strategy for the leader.
+Additionally, studies such as (Bensoussan et al.
+2017; Lin, Jiang, and Zhang 2018; Moon and Başar 2018; Du, Wu et al.
+2019; Huang and Yang 2020) simplify scenarios through linear-quadratic formulations.
+(Guo, Hu, and Zhang 2022), convert SMFGs into minimax optimization problems.
+However, these model-based methods generally depend on simplified representations of SMFGs and explicit environmental transition information.
+In contrast, model-free approaches leverage interaction data with the environment to learn strategies without requiring detailed knowledge of the transition dynamics and reward functions.(Pawlick and Zhu 2017) explores the resolution of SMFGs in single-step decision scenarios.(Campbell et al.
+2021) utilizes a modified deep backward stochastic differential equation framework to calculate the followers’ equilibrium despite constraints on the leader’s decisions.(Miao et al.
+2024) investigates locally optimal defensive strategies given predefined attacker trajectories.
+Furthermore,(Li et al.
+2024) learns followers’ transition functions from empirical data, thus aiding the solution of SMFGs through the empirical Fokker-Planck equation.
+Nevertheless, the assumptions and constraints inherent in these algorithms often do not meet the complex needs of macroeconomic policymaking.
+3 Stackelberg Mean Field Game Modeling
+In this section, we provide details about macroeconomic policy-making problems and propose a modeling framework based on Dynamic Stackelberg Mean Field Games (Dynamic SMFGs), with the relevant mappings are shown in Figure 1.
+Macroeconomic Policy-Making
+To foster economic growth, enhance social welfare, or maintain social equity, governments implement economic policies such as fiscal and monetary policies.
+These policies directly influence micro-agents (i.e., households) and the markets they constitute, altering their expectations and behaviors.
+The aggregation of micro-level behaviors from numerous households manifests as macroeconomic phenomena.
+Therefore, in formulating policies, governments should model micro-agents’ behaviors and predict their responses to new policies, which is so-called microfoundations in the Lucas critique.
+However, in reality, micro-agents form a large and dynamically interacting group, each seeking to maximize personal benefits.
+This complexity significantly complicates predicting their responses to macroeconomic policy shifts.
+To address these challenges, we propose using Dynamic SMFGs to model macroeconomic policy-making problems:
 1.
-It innovatively proposes a general framework based on Dynamic Stackelberg Mean Field Games to model macroeconomic policy-making, considering key issues such as microfoundations, dynamic interactions and feedback, rational expectations, and randomness.
+For microfoundations, we model the government as the leader who first sets policies, with micro-agents as followers who adjust their behaviors according to these policies.
+The interaction between the leader and followers forms a Stackelberg game, where the leader considers the followers’ responses to make informed decisions.
 2.
-It introduces a data-driven Stackelberg Mean Field Reinforcement Learning algorithm to solve dynamic SMFGs without environmental transition information and prior knowledge, validated experimentally for its effectiveness.
-3. It empirically demonstrates the superiority of the SMFG method in modeling and solving optimal macroeconomic policy problems and discusses its robustness in real-world scenarios.
-The source code for the SMFG method is available in an anonymous GitHub repository  [^1] https://anonymous.4open.science/r/SMFG_macro_7740.
-2 Related work
-2.1 Economic Methods
-In the field of economics, research on economic policy typically relies on theoretical model studies, empirical analysis, and econometric methods.
-Classic theoretical models, such as the IS-LM model [24, 22], are applicable for analyzing short-term policy effects but overlook long-term factors and price fluctuations.
-The aggregate demand-aggregate supply (AD-AS) model [19, 34] focuses on the relationship between aggregate demand and supply, integrating short-term and long-term factors, but simplifies macroeconomic dynamics.
-The Solow growth model [10] provides a framework for economic growth but does not consider market imperfections and externalities.
-New Keynesian models [9, 21] emphasize the stickiness of prices and wages, suitable for explaining economic fluctuations and policy interventions, but are based on strong assumptions.
-DSGE models [1, 55], built on micro-foundations for macroeconomic forecasting, offer a consistent framework for policy analysis but are limited in handling nonlinearities and market imperfections.
-Empirical analysis [47, 59] primarily evaluates policy effects through historical data but faces limitations in data timeliness and accuracy.
-Econometric methods [30, 15], such as regression and time series analysis, can quantify policy effects but are constrained by model settings and data quality.
-2.2 AI for Economics
-Artificial intelligence (AI), such as reinforcement learning, may offer new perspectives for solving complex economic problems [56].
-In terms of macroeconomics, AI Economist [62] employs curriculum learning on tax policy design, [58] studies the collaboration between central and local governments under COVID-19, as well as monetary policy [26, 12], international trade [51], market pricing [14].
-However, these studies consider the government as a single agent, neglecting the dynamic response of the households to policies.
-[31] studies democratic AI by human voting models, [60] examines the impact of social sanction rules on labor division.
-However, these works are based on simplified settings and remain a gap in real-world economic policy.
-In terms of microeconomics, [53, 48, 3] study the optimal saving and consumption problems of micro-agents, and other researchers explore rational expectation equilibrium [32, 25], multiple equilibria under real-business-cycle [13], the emergence of barter behavior [29], optimal asset allocation and savings strategies [44].
-In summary, research on macroeconomic problems still lacks a bridge framework of macro policy and microfoundations, which has drawn our attention to the Stackelberg mean field game.
-Figure 1: The mapping from optimal macroeconomic policy problems (left) to the Dynamic Stackelberg Mean Field Games (right).
-The left represents the reality where the macro-level government implements fiscal or monetary policies to optimize GDP, welfare, or equity, and the micro-level consists of large-scale households engaging in economic activities.
-In dynamic SMFGs, the government acts as the leader, and households as followers, respond to the leader’s policy.
-2.3 Stackelberg Mean Field Game
-Stackelberg Mean Field Games (SMFG) model interactions between the leader and numerous homogeneous followers, employing both model-based and model-free approaches.
-Model-based methods, such as those presented in [23], treat SMFGs as minimax optimization problems, streamlining the leader’s strategy to a single optimal action rather than a comprehensive policy.
-Studies by [7, 20, 16, 8] utilize forward-backward stochastic functional differential equations to determine the followers’ mean field equilibrium before computing the leader’s optimal strategy.
-This includes addressing terminal state constraints [20] and simplifying scenarios via linear-quadratic formulations [6, 37, 43, 18, 27].
-Such methods, though theoretically robust, often rely on simplifying real-world complexities and require extensive knowledge of environmental transitions.
-Extending SMFG to model epidemic dynamics [4] showcases its broader applicability in public health management.
-Conversely, model-free approaches leverage complex, data-rich environments without the need for prior environmental knowledge.
-[11] employs a modified deep backward stochastic differential equation framework to establish followers’ equilibrium, despite constraints on leader decisions.
-[45] investigates applications in privacy protection through a single-step decision model.
-[42] delves into locally optimal defensive strategies given predefined attacker trajectories.
-[35] learns followers’ transition functions from empirical data, and facilitates SMFG solutions through the empirical Fokker-Planck equation.
-However, in real-world settings, the sequential decision-making between a leader and numerous followers to optimize long-term gains is common but inadequately studied.
-This paper proposes a data-driven reinforcement learning algorithm for solving the aforementioned problem, termed the Dynamic Stackelberg Mean Field Game, without reliance on environmental transition information or prior knowledge.
-3 Stackelberg Mean Field Game Modeling
-3.1 Macroeconomic Policy-Making
-Macroeconomic policy-making involves formulating effective policies to regulate and guide an economy toward achieving sustainable growth, low unemployment, and stable prices.
-In addressing these challenges, several key elements must be considered: ① Microfoundations: According to Lucas critique, basing macroeconomic models on microeconomic behaviors is crucial.
-This ensures that the actions of individual agents, such as households and firms, aggregate to produce macroeconomic outcomes.
-② Dynamic Interactions: Extensive interactions occur both between the macro-agent (the government) and a multitude of micro-agents (households), and within groups of households.
-These interactions can lead to complex economic dynamics and endogenous shocks.
-③ Rational Expectations: Agents are typically modeled as rational and self-interested, making decisions based on available information and their expectations about future policies and economic conditions.
-④ Stochastic Elements: Incorporating randomness and shocks, such as technological changes, policy shifts, and external economic shocks, is essential for understanding variability in economic outcomes.
-⑤ Feedback Mechanisms: Feedback loops exist within the economy where macro-policy impacts micro-agents, and the micro-agents’ behaviors, in turn, influence policy decisions.
-Given the aforementioned key points, Dynamic Stackelberg Mean Field Games (Dynamic SMFGs) are ideal for macroeconomic policy-making due to:
-1.
-Leader-Follower Dynamics (②⑤): The government, acting as the leader, sets policies first.
-Households, as followers, adjust their behaviors based on these policies, which captures the strategic interactions between policy-setters and economic agents.
-2.
-Mean Field Interaction (①⑤): The collective behavior of numerous micro-agents is efficiently modeled as a "mean field", which simplifies the analysis of large-scale interactions and their overall impact on the macroeconomy.
-3. Dynamic Strategy (②③④): The dynamic nature of these games reflects the evolving economic conditions and policy-making, enabling the model to adapt to new information and economic shocks.
-3.2 Dynamic Stackelberg Mean Field Games
-Following this, we consider a dynamic SMFG over discrete periods $t\in\{0,\ldots,T\}$ for macroeconomic policy-making issues: At each time $t$, the leader first selects an action $a_{t}^{l}\in\mathcal{A}^{l}$ based on the leader’s state $s_{t}^{l}\in\mathcal{S}^{l}$ and policy $\pi^{l}:\mathcal{S}^{l}\rightarrow\mathcal{A}^{l}$.
-Subsequently, followers simultaneously determine their actions based on the leader’s action $a^{l}_{t}$ and their private states $s^{f}_{t}\in\mathcal{S}^{f}$.
+For large scale micro-agents, we employ a Mean Field Game to model micro-interactions among a large number of households, treating them as a population and predicting their collective response to macroeconomic policy changes.
+3.
+For dynamic decision-making, the formulation of macroeconomic policies is inherently a multi-step decision-making process, where both the government and households make decisions at each step to optimize long-term benefits, thereby constituting a dynamic game.
+Dynamic Stackelberg Mean Field Games
+For macroeconomic policy-making issues, we consider a dynamic SMFG with a leader and $N$ homogeneous follower agents.
+At any time step $t\in\{0,\ldots,T\}$, the leader selects an action $a_{t}^{l}\in\mathcal{A}^{l}$ based on the leader’s state $s_{t}^{l}\in\mathcal{S}^{l}$ and policy $\pi^{l}:\mathcal{S}^{l}\rightarrow\mathcal{A}^{l}$.
+Subsequently, followers determine their actions based on the leader’s action $a^{l}_{t}$ and their private states $s^{f}_{t}\in\mathcal{S}^{f}$.
 The representative follower’s action $a_{t}^{f}\in\mathcal{A}^{f}$ is derived from a shared policy $\pi^{f}:\mathcal{S}^{f}\times\mathcal{A}^{l}\rightarrow\mathcal{A}^{f}$.
 The sequences $\{\pi_{t}^{l}\}_{t=0}^{T}$ and $\{\pi_{t}^{f}\}_{t=0}^{T}$ are denoted as $\pi^{l}$ and $\pi^{f}$, respectively.
+We denote the population state-action distribution of followers under leader action $a^{l}_{t}$ and follower policy $\pi^{f}$ as $L_{t}(s^{f}_{t},a^{f}_{t};\pi^{f},a^{l}_{t})$, and abbreviate it as $L_{t}(s^{f}_{t},a^{f}_{t})$.
+$L_{t}(s^{f}_{t},a^{f}_{t};\pi^{f},a^{l}_{t})\in\mathcal{P}(\mathcal{S}^{f}\times\mathcal{A}^{f}),\text{where }a^{f}_{t}=\pi^{f}(s^{f}_{t},a^{l}_{t})$
 Followers’ Mean Field Game
 Given the leader’s policy $\pi^{l}$, the dynamic SMFG is simplified into a mean field game for the followers.
-At time $t\in\{0,...,T-1\}$, given the leader’s action $a^{l}_{t}$, the representative follower’s state $s^{f}_{t}$ and action $a^{f}_{t}$, and the population state-action distribution of all followers denoted as $L^{f}_{t}\in\mathcal{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$, the representative follower receives a reward $r^{f}(s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$ and move to the next state $s_{t+1}^{f}\sim\mathbb{P}(\cdot|s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$.
+At time $t\in\{0,\ldots,T-1\}$, given $(a^{l}_{t},s^{f}_{t},a^{f}_{t},L_{t})$, the representative follower then receives a reward $r^{f}(\mathbf{s}_{t},a^{l}_{t},a^{f}_{t},L_{t})$ and transitions to the next state $s^{f}_{t+1}$ according to the transition probability $s^{f}_{t+1}\sim P(\cdot|\mathbf{s}_{t},a^{l}_{t},a^{f}_{t},L_{t})$, where joint state $\mathbf{s}_{t}=\{s^{l}_{t},\mathbf{s}_{t}^{f}\}$.
 Each follower aims to find the optimal policy $\pi^{f}$ that maximizes his cumulative reward over the time horizon:
-$J^{f}(\pi^{l},\pi^{f},L^{f})=\mathbb{E}^{f}\left[\sum_{t=0}^{T}r^{f}(s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})\right]$ (1)
-where $\mathbb{E}^{f}$ denotes the expectation with respect to initial state $s_{0}^{f}\in\mu_{0}^{f}$, $(s_{t+1}^{l},s_{t+1}^{f})\sim P(\cdot\mid s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$ for $t\in\{0,...,T-1\}$, and $a^{l}_{t}\sim\pi^{l}(\cdot\mid s_{t}^{l}),a^{f}_{t}\sim\pi^{f}(\cdot\mid s_{t}^{f},a^{l}_{t})$ for $t\in\{0,...,T\}$.
-Definition 3.1.
-[Followers’ Best Response for Leader’s Policy] Given any leader’s policy $\pi^{l}\in\Pi^{l}$ and the followers’ state-action distribution $L^{f}\in\mathcal{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$, the followers’ best response policy ${\pi^{f}}^{*}(\pi^{l},L^{f})$ is defined as
-${\pi^{f}}^{*}(\pi^{l},L^{f})\in\arg\max_{\pi^{\prime}}J^{f}(\pi^{l},\pi^{\prime},L^{f}).$
+$J^{f}(\pi^{l},\pi^{f},L)=\mathbb{E}_{s_{0}^{f}\sim\mu_{0}^{f},\mathbf{s}_{t+1}\sim P}\left[\sum_{t=0}^{T}r^{f}(\mathbf{s}_{t},a^{l}_{t},a^{f}_{t},L_{t})\right]$
+where $a^{l}_{t}=\pi^{l}(s^{l}_{t})$, $a^{f}_{t}=\pi^{f}(s^{f}_{t},a^{l}_{t})$ and $L_{t}=L_{t}(s^{f}_{t},a^{f}_{t})$.
+Definition 1 (Followers’ Best Response for Leader’s Policy).
+Given any leader’s policy $\pi^{l}\in\Pi^{l}$ and the followers’ state-action distributions $L=\{L_{t}\}_{t=0}^{T}$, the followers’ best response policy ${\pi^{f}}^{*}(\pi^{l},L)$ is defined as
+${\pi^{f}}^{*}(\pi^{l},L)\in\arg\max_{\pi^{\prime}}J^{f}(\pi^{l},\pi^{\prime},L).$
 Leader’s Game
-At time $t\in\{0,...,T-1\}$, given the leader’s state $s^{l}_{t}$, action $a^{l}_{t}$, and the followers’ state-action distribution $L_{t}^{f}$, the leader will receive a reward $r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})$, and move to a new state $s_{t+1}^{l}$ according to the transition probabilities $\mathbb{P}(\cdot|s_{t}^{l},a_{t}^{l},L^{f}_{t})$.
+At time $t\in\{0,...,T-1\}$, given the leader’s state $s^{l}_{t}$, action $a^{l}_{t}$, the followers’ policy $\pi^{f}$ and state-action distribution $L_{t}$, the leader will receive a reward $r^{l}(\mathbf{s}_{t},a_{t}^{l},L_{t})$, and move to the next state $s_{t+1}^{l}$ according to the transition probabilities $P(\cdot|\mathbf{s}_{t},a_{t}^{l},L_{t})$.
 The leader aims to find the optimal policy $\pi^{l}$ to maximize the total expected reward:
-$J^{l}(\pi^{l},\pi^{f},L^{f})=\mathbb{E}^{l}\left[\sum_{t=1}^{T}r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})\right],$ (2)
-where $\mathbb{E}^{l}$ denotes the expectation under this probability distribution for the leader when $s_{0}^{l}\sim\mu_{0}^{l},s_{t+1}^{l}\sim\mathbb{P}(\cdot|s_{t}^{l},a_{t}^{l},L^{f}_{t})$ for $t\in\{0,...,T-1\}$, and $a_{t}^{l}\sim\pi_{t}^{l}(\cdot\mid s_{t}^{l})$ for $t\in\{0,...,T\}$.
-Definition 3.2 (Leader’s Optimal Policy in Dynamic Stackelberg Mean Field Games).
-Considering the followers’ best response $(\pi^{f},L^{f})$ for the leader’s policy, which satisfies Definition 3.1, the leader’s optimal policy ${\pi^{l}}^{*}$ in dynamic Stackelberg mean field games is defined as follows:
-${\pi^{l}}^{*}$ $\in\mathop{\arg\max}_{{\pi^{l}}^{\prime}}J^{l}({\pi^{l}}^{\prime},\pi^{f},L^{f}),$
+$J^{l}(\pi^{l},\pi^{f},L)=\mathbb{E}_{s_{0}^{l}\sim\mu_{0}^{l},\mathbf{s}_{t+1}\sim P}\left[\sum_{t=1}^{T}r^{l}(\mathbf{s}_{t},a_{t}^{l},L_{t})\right],$ (1)
+where $a^{l}_{t}=\pi^{l}(s^{l}_{t})$, $a^{f}_{t}=\pi^{f}(s^{f}_{t},a^{l}_{t})$ and $L_{t}=L_{t}(s^{f}_{t},a^{f}_{t})$.
+Definition 2 (Leader’s Optimal Policy in Dynamic Stackelberg Mean Field Games).
+Considering the followers’ best response $(\pi^{f},L)$ to the leader’s policy, which satisfies Definition 1, learning the leader’s optimal policy ${\pi^{l}}^{*}$in dynamic Stackelberg mean field games is equivalent to solving the following fixed-point problem given initial condition $(\mu^{l}_{0},\mu^{f}_{0})$:
+${\pi^{l}}^{*}$ $\in\mathop{\arg\max}_{{\pi^{l}}^{\prime}}J^{l}({\pi^{l}}^{\prime},\pi^{f},L)$
 […]
-where the followers’ state-action distribution $L^{f}$ satisfies the following Mckean-Vlasov equation:
-$L^{f}_{t+1}(s_{t+1}^{f},a_{t+1}^{f})=\sum_{s_{t}^{l},a_{t}^{l},s_{t}^{f},a_{t}^{f}}L^{f}_{t}(s_{t}^{f},a_{t}^{f})\pi_{t}^{l}(a_{t}^{l}\mid s_{t}^{l})\mu_{t}^{l}(s_{t}^{l})P(s_{t+1}^{f}\mid s_{t}^{f},a_{t}^{f},a_{t}^{l})\pi_{t+1}^{f}(a_{t+1}^{f}\mid s_{t+1}^{f}),$
-$\mu_{t+1}^{l}(s_{t+1}^{l})=\sum_{s_{t}^{l},a_{t}^{l}}\mu_{t}^{l}(s_{t}^{l})\pi_{t}^{l}(a_{t}^{l}\mid s_{t}^{l})P(s_{t+1}^{l}\mid s_{t}^{l},a_{t}^{l}).$ (3)
-In conclusion, to study the issue of macroeconomic policy-making, we model it as a dynamic Stackelberg mean field game.
-This modeling approach aims to learn the optimal policy of the leader, namely the social planner, while considering the followers’ best response over discrete timesteps (Definition 3.2).
-3.3 Metrics for Dynamic SMFGs
-We will utilize the following three metrics to evaluate the effectiveness of the leader and follower policies: (1) The leader’s payoff, which indicates the performance of the leader’s policy in optimizing the leader’s objective; (2) Exploitability, which measures the deviation of the followers’ policy from the best response to the leader’s policy; (3) Social welfare, which assesses the deviation of the current state from the social optimum.
-Leader’s Payoff
-We define the leader’s payoff using the long-term expected rewards of the leader’s policy $\pi^{l}$ over discrete timesteps, as detailed in Equation 2.
-Exploitability
-Exploitability is a widely used metric for evaluating convergence and measuring the distance between the representative follower’s policy $\pi^{f}$ and the best response ${\pi^{f}}^{*}$ for the leader’s policy $\pi^{l}$.
-Formally, it quantifies the extent to which a single follower can improve their payoff by deviating from the population’s behavior to adopt a different policy.
-Given the leader’s policy $\pi^{l}$ and the followers’ state-action distribution $L^{f}$, the exploitability $\mathcal{E}(\pi^{f};\pi^{l})$ of the follower’s policy $\pi^{f}$ is defined by comparing the value functions induced by $\pi^{f}$ and its best response $\pi^{f*}$:
-$\mathcal{E}(\pi^{f};\pi^{l})=J^{f}(\pi^{l},{\pi^{f}}^{*},L^{f})-J^{f}(\pi^{l},{\pi^{f}},L^{f}).$
-Social Optimum and Social Welfare
-In economic theory, the Social Optimum describes a state in which the allocation of resources achieves maximum efficiency, as measured by social welfare [2, 40].
-Given the leader’s policy $\pi^{l}$ and the representative follower’s policy $\pi^{f}$ among large-scale followers, social welfare $\mathcal{SW}(\pi^{l},\pi^{f})$ is approximately calculated as the sum of the utility functions of the $N$ followers in Equation 1 when $N\to\infty$:
-$\mathcal{SW}(\pi^{l},\pi^{f})=\sum^{N}J^{f}\left(\pi^{l},\pi^{f},L^{f}\right).$
+where the followers’ state-action distribution $L_{t}$ satisfies the following McKean-Vlasov equation:
+$L_{t+1}(s_{t+1}^{f},a_{t+1}^{f})=\sum_{s_{t}^{l},a_{t}^{l},s_{t}^{f},a_{t}^{f}}L_{t}(s_{t}^{f},a_{t}^{f})\pi_{t}^{l}(a_{t}^{l}\mid s_{t}^{l})\mu_{t}^{l}(s_{t}^{l})$
+$P(s_{t+1}^{f}\mid s_{t}^{f},a_{t}^{f},a_{t}^{l},L_{t})\pi_{t+1}^{f}(a_{t+1}^{f}\mid s_{t+1}^{f}),$
+$\mu_{t+1}^{l}(s_{t+1}^{l})=\sum_{s_{t}^{l},a_{t}^{l}}\mu_{t}^{l}(s_{t}^{l})\pi_{t}^{l}(a_{t}^{l}\mid s_{t}^{l})P(s_{t+1}^{l}\mid s_{t}^{l},a_{t}^{l},L_{t}).$
+In conclusion, to study the issue of macroeconomic policy-making, we model it as a dynamic SMFG.
+Under this modeling framework, we can optimze the macroeconomic policy of the leader, namely the government, while considering the followers’ best response over discrete timesteps (Definition 2).
 4 Stackelberg Mean Field Reinforcement Learning
-However, solving for the leader’s optimal policy in dynamic SMFGs, as defined in Definition 3.2, presents several challenges, such as vast state and action spaces, strategic dependencies and information asymmetry between the leader and followers, and potentially conflicting objectives including the leader’s payoff, followers’ utilities, and social welfare.
-These complexities make devising a model-based algorithm with theoretical proof extremely difficult.
-Previous works have attempted simplifications using the myopic followers setting [63] or linear-quadratic MFGs [6], but this approach fails to align with macroeconomic policy-making issues.
-Consequently, we have designed the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm to address the above limitations.
-This is a model-free approach that operates independently of environmental transition information.
+To solve macroeconomic policymaking under Dynamic SMFG framework, we propose the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm, which is a centralized training with decentralized execution (CTDE) algorithm.
+We introduce the leader’s policy $\pi^{l}(s^{l})$ and a central critic Q-function $Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ for the leader agent, and the shared policy $\pi^{f}(s^{f},a^{l})$ and $Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ for each follower agent.
+SMFRL algorithm comprises two key components: mean field approximation and leader-follower update (see Figure 2).
 Figure 2: The architecture of SMFRL algorithm.
-Considering the dynamic SMFGs with a leader and $N$ homogeneous follower agents, at any time $t\in\{0,\ldots,T-1\}$, the joint action of the leader and follower agents $\mathbf{a}_{t}=\{a^{l}_{t},a^{f1}_{t},...,a^{fN}_{t}\},a^{fi}_{t}=\pi^{f}(s^{fi}_{t})$ impacts the environment and results in a reward $\mathbf{r}_{t}=\{r^{l}_{t},r^{f1}_{t},...,r^{fN}_{t}\}$ and the next state $\mathbf{s}_{t+1}=\{s^{l}_{t+1},s^{f1}_{t+1},...,s^{fN}_{t+1}\}$.
-The experience replay buffer $\mathcal{D}$ contains the tuples $(\mathbf{s}_{t},\mathbf{s}_{t+1},\mathbf{a}_{t},\mathbf{r}_{t},L^{f}_{t})$, recording the experiences of all agents.
-Specifically, we define the leader’s policy as $\pi_{\theta^{l}}$ parameterized by $\theta^{l}$ (abbreviated as $\pi^{l}$), and a shared policy for the followers denoted as $\pi_{\theta^{f}}$ with parameters $\theta^{f}$ (abbreviated as $\pi^{f}$).
-The SMFRL algorithm incorporates several settings (shown in Figure 2) to address the aforementioned challenges:
 Mean Field Approximation
-To mitigate the problem of non-stationarity, we consider a central critic Q function $Q(\mathbf{s}_{t},\mathbf{a}_{t}):\mathcal{S}^{l}\times\{\mathcal{S}^{f}\}^{N}\times\mathcal{A}^{l}\times\{\mathcal{A}^{f}\}^{N}\to\mathbb{R}$ for each agent.
-However, due to the high dimensionality of the joint state and action, the standard Q network is infeasible to learn.
-Thus, the SMFRL algorithm employs a mean field approximation, characterizing the followers’ state-action pairs $(\mathbf{s}^{f}_{t},\mathbf{a}^{f}_{t})$ by using the followers’ population distribution $L^{f}_{t}$, which can be presented through various metrics such as mean, variance, probability density functions, skewness, and kurtosis.
-The theoretical proof is detailed in Appendix F.
-This approach yields $Q^{l}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$ and $Q^{f}(s^{f}_{t},a^{l}_{t},a^{f}_{t},L^{f}_{t})$ for evaluating the leader and follower agents respectively.
-The mean field approximation effectively reduces the dimensionality of the inputs to the Q function, thereby facilitating the learning of the critic Q network and ensuring the Q function can evaluate leader and follower policies based on the overall follower distribution.
+Due to the dimensions of joint state $\mathbf{s}^{f}$ and action $\mathbf{a}^{f}$ scaling with the number of follower agents, the standard Q-function $Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ becomes infeasible to learn when dealing with a large number of followers.
+Therefore, we factorize the standard Q-function using pairwise local interactions between the leader and each follower (Yang et al.
+2018):
+$Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ $=\frac{1}{N}\sum_{i}^{N}\tilde{Q}^{l}(s^{l},a^{l},s^{fi},a^{fi})$ (2)
+$\approx\mathbb{E}_{(s^{f},a^{f})\sim L}\tilde{Q}^{l}(s^{l},a^{l},s^{f},a^{f})$
+Similarly, the $i$-th follower’s Q-function $Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ can be factorized using local interactions between the leader, the $i$-th follower, and each of the remaining followers:
+$Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})\approx\mathbb{E}_{(s^{f},a^{f})\sim L}\tilde{Q}^{f}(s^{l},a^{l},s^{fi},a^{fi},s^{f},a^{f})$
+where $(s^{fi},a^{fi})$ and $(s^{f},a^{f})$ are the state-action pair for the $i$-th follower and another follower agent, respectively.
+We use $\hat{L}(\pi^{f},a^{l})$, a finite set of state-action pairs sampled under the follower’s policy $\pi^{f}$ and the leader’s action $a^{l}$, to construct an empirical distribution that approximates the state-action distribution $L(s^{f},a^{f})$.
+$\hat{L}(\pi^{f},a^{l})=\left\{(s^{fi},a^{fi})\right\}_{i=1}^{N},\text{ where }a^{fi}=\pi^{f}(s^{fi},a^{l}).$
+Leader-follower Update
+To enhance the convergence and stability, we propose the leader-follower update: first, by fixing the leader’s policy, we train the followers’ shared policy and Q-networks towards the best response; subsequently, based on the followers’ policies, we optimize the leader’s policy and Q net, alternating these steps until convergence.
+We measure the distance between the agents’ policies and their best responses by using exploitability.
+Based on mean field approximation, we will train these networks $\pi_{\theta^{l}}$ and $\tilde{Q}_{\phi^{l}}$ for the leader agent, shared $\pi_{\theta^{f}}$ and $\tilde{Q}_{\phi^{f}}$ for the follower agents, with parameters $\theta^{l}$, $\phi^{l}$, $\theta^{f}$, and $\phi^{f}$.
+To ensure training stability, we introduce target networks with parameters $\theta^{l}_{-}$, $\phi^{l}_{-}$, $\theta^{f}_{-}$, and $\phi^{f}_{-}$.
+At any step $t\in\{0,\ldots,T\}$, the tuple $(s^{l}_{t},a^{l}_{t},\mathbf{s}^{f}_{t},\mathbf{a}^{f}_{t},s^{l}_{t+1},\mathbf{s}^{f}_{t+1},r^{l}_{t},\mathbf{r}^{f}_{t})$ is stored in the replay buffer $\mathcal{D}$ for training.
+When the number of followers is particularly large, a finite subset of state-action pairs $\{(s^{fi}_{t},a^{fi}_{t})\}_{i=1}^{N}$ can be selected for storage.
+Followers’ Update for Best Response
+Given leader’s policy $\pi_{\theta^{l}}$, the followers’ policy network $\pi_{\theta^{f}}$ is updated using a deterministic policy gradient approach (Silver et al.
+2014).
+The followers’ policy gradient is estimated:
+$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t}\sim\mathcal{D},(s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}}$
+$\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{fi}_{t},a^{l}_{t})\nabla_{a^{fi}_{-}}\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{fi}_{t},a^{fi}_{-},s^{f}_{t},a^{f}_{t})\Big{]}$
+where $a^{fi}_{-}=\pi_{\theta^{f}}(s^{fi}_{t},a^{l}_{t})$, $\hat{L}_{\mathcal{D}}$ denotes state-action pairs sampled from replay buffer $\mathcal{D}$.
+The action-value function $\tilde{Q}_{\phi^{f}}$ is updated by minimizing the mean squared error loss:
+$\mathcal{L}(\phi^{f})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\\ (s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}\end{subarray}}\left[\left(y^{fi}_{t}-\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{fi}_{t},a^{fi}_{t},s^{f}_{t},a^{f}_{t})\right)^{2}\right]$
+$y^{fi}_{t}=r^{fi}_{t}+$
+$\gamma\mathbb{E}_{(s^{f}_{t+1},a^{f}_{t+1})\sim\hat{L}(\pi_{\theta^{f}_{-}},a^{l}_{t+1})}\tilde{Q}_{\phi^{f}_{-}}(s^{l}_{t+1},a^{l}_{t+1},s^{fi}_{t+1},a^{fi}_{t+1},s^{f}_{t+1},a^{f}_{t+1})$
+where $a^{l}_{t+1}=\pi_{\theta^{l}}(s^{l}_{t+1})$, $a^{fi}_{t+1}=\pi_{\theta^{f}_{-}}(s^{fi}_{t+1},a^{l}_{t+1})$. The gradient of the loss function $\mathcal{L}(\phi^{f})$ is derived as:
+$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\\ (s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}\end{subarray}}$ $\Big{[}\left(y^{fi}_{t}-\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{fi}_{t},a^{fi}_{t},s^{f}_{t},a^{f}_{t})\right)$
+$\nabla_{\phi^{f}}\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{fi}_{t},a^{fi}_{t},s^{f}_{t},a^{f}_{t})\Big{]}.$
 Leader’s Update for Optimal Policy
-In SMFRL, the leader’s policy $\pi_{\theta^{l}}$, is optimized using a deterministic policy gradient approach [54]. The policy gradient for the leader can be estimated as:
-$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{-},L^{f}_{t})|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}\right],$
-where ${Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$, the leader’s action-value function, is approximated by a neural network parameterized by $\phi^{l}$.
+Given followers’ policy $\pi_{\theta^{f}}$, the leader’s policy $\pi_{\theta^{l}}$ is optimized by DPG approach, and the leader’s policy gradient is estimated:
+$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D},(s^{f},a^{f})\sim\hat{L}(\pi_{\theta^{f}},a^{l}_{-})}$
+$\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}\tilde{Q}_{\phi^{l}}(s^{l},a^{l}_{-},s^{f},a^{f})\right],\text{where }{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}.$
 This network is periodically updated to minimize the loss:
-$\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)^{2}\right],$
-where the target value $y^{l}_{t}$ is given by:
-$y^{l}_{t}=r_{t}^{l}+\gamma{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L^{f}_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})},$
-and $\phi^{l}_{-}$ represents the parameters of the target Q network, $L^{f}_{t+1}$ is determined based on Mckean-Vlasov equation (3), and $\gamma$ is the discount factor.
-Differentiating the loss function $\mathcal{L}(\phi^{l})$ yields the gradient utilized for training:
-$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{l}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right].$
-Followers’ Update for Best Response
-To enable follower agents to effectively learn the best response policy $\pi^{f*}$ to the leader’s policy $\pi^{l}$, we employ multiple updates for followers’ shared policy and Q networks in the SMFRL algorithm.
-Each update cycle refines the policy network $\pi_{\theta^{f}}$ for the followers using the policy gradient method:
-$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L^{f}_{t}\sim\mathcal{D}}\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{-},a^{l}_{t},L^{f}_{t})|_{a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})}\Big{]},$
-where the action-value function ${Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})$, parameterized by $\phi^{f}$, is periodically updated by minimizing the mean squared error (MSE) loss:
-$\mathcal{L}(\phi^{f})=$ $\mathbb{E}_{{s}_{t}^{f},{a}_{t}^{f},a^{l}_{t},L^{f}_{t},s_{t+1}^{f}\sim\mathcal{D}}\left[\left(y^{f}_{t}-{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)^{2}\right],$ (4)
-$y^{f}_{t}=$ $r^{f}_{t}+\gamma{Q}_{\phi^{f}_{-}}(s^{f}_{t+1},a^{f}_{t+1},a^{l}_{t+1},L^{f}_{t+1}),$
-with the next leader’s action $a^{l}_{t+1}=\mu_{\phi^{l}_{-}}(s^{l}_{t+1})$, and the next population distribution $L^{f}_{t+1}=L(\mathbf{s}_{t+1}^{f},\mathbf{a}_{t+1}^{f})$.
-The joint actions of followers $\mathbf{a}_{t+1}^{f}=\{a^{f1}_{t+1},\ldots,a^{fN}_{t+1}\}$ are determined with each action $a^{fi}_{t+1}=\mu(s^{fi}_{t+1},a^{l}_{t+1})$.
-The gradient of the loss function $\mathcal{L}(\phi^{f})$ is derived as:
-$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\mathcal{D}}\Big{[}\left(y^{f}_{t}-{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{f}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\Big{]}.$
-The pseudocode for the SMFRL algorithm (1) is presented in Appendix A.
-Table 1: Evaluation of SMFRL and five baselines on 3 metrics for dynamic Stackelberg mean field games with different followers’ numbers.
-Metrics Followers’ Optimal SMFRL MF-MARL MADDPG IDDPG DDPG U.S. Policy
-Number Value +Real +Real
-Leader’s Payoff N=100 \ 3387.36 -76.82 -282.62 -701.61 2840.07 1718.08
-N=1000 \ 3411.16 -668.03 121.18 -767.53 2874.19 -796.4
-Exploitability N=100 0. 0.000107 0.002190 0.007908 0.001333 0.009885 0.032490
-N=1000 0. 0.000158 0.003481 0.035613 0.000565 0.003002 0.010522
-Social Welfare N=100 100 96.87 82.02 55.74 83.09 94.5 79.23
-N=1000 1000 968.94 834.93 645.38 874.53 943.82 334.79
+$\mathcal{L}(\phi^{l})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\\ (s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}\end{subarray}}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{t})\right)^{2}\right]$
+The target value $y^{l}_{t}$ is given by:
+$y^{l}_{t}=r_{t}^{l}+\gamma\mathbb{E}_{(s^{f}_{t+1},a^{f}_{t+1})\sim\hat{L}(\pi_{\theta^{f}},a^{l}_{t+1})}\tilde{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},s^{f}_{t+1},a^{f}_{t+1})$
+where $a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})$, and $\gamma$ is the discount factor. Differentiating the loss function $\mathcal{L}(\phi^{l})$ yields the gradient utilized for training:
+$\nabla_{\phi^{l}}$ $\mathcal{L}(\phi^{l})=\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},\mathbf{s}_{t+1}\sim\mathcal{D},(s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}}$
+$\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{t})\right)\nabla_{\phi^{l}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{t})\right].$
+The pseudocode for the SMFRL algorithm 1 is presented in Appendix A.
 5 Experiment
-In this section, we conduct experiments within the TaxAI environment to answer the following three questions: (1) How effective is our SMFRL algorithm in solving dynamic SMFGs?
-(2) For macroeconomic policy-making issues, how does our method compare to real-world strategies, economic models, and existing AI-based methods?
-(3) Given the likely real-world scenario where some households do not adopt the SMFG household policy learned by our method, does our approach remain effective?
+In this section, we conduct experiments of the SMFG method, which is modeled by dynamic SMFGs and solved by the SMFRL algorithm, within the TaxAI (Mi et al.
+2023) environment to answer the following questions:
+1. For macroeconomic policy-making issues, how does our SMFG method compare to real-world policies, classic economic methods, and existing AI-based policy?
+2. Are both the leader-follower update and mean-field approximation modules of the SMFG method essential?
+3. Given the likely real-world scenario where some households do not adopt the SMFG policy, does the SMFG approach remain effective?
 TaxAI Environment
-TaxAI [41] is one of the rare reinforcement learning environments designed to study optimal macroeconomic policies.
-It simulates dynamic interactions between large-scale households and the government, making it particularly well-suited for experimenting with SMFGs.
-TaxAI models the optimal tax policy problems where the government optimizes tax policies to promote economic growth (i.e.
-GDP) and social equity (i.e.
-Gini).
-In response, households optimize their labor supply and savings strategies to maximize their long-term utility.
-5.1 Solving Dynamic SMFGs
-Metrics and Baselines
-We compare several classic multi-agent reinforcement learning (MARL) algorithms based on three metrics: leader’s payoff, exploitability, and social welfare, as defined in Section 3.3.
-Optimal Value: The ideal value of exploitability is 00, indicating that the follower’s policy achieves the best response to the leader’s policy.
-The upper bound of Social Welfare is $N$, given that household utility in TaxAI cannot exceed 1.
-For Leader’s Payoff, no maximum has been established.
-MF-MARL [61]: Utilizes the DDPG [36] algorithm for the leader and a mean-field MARL approach for the followers.
-MADDPG [38] and IDDPG [17]: Both feature shared actor and critic networks among followers.
-DDPG+Real: DDPG for the leader and behavior cloning [57] from real data for followers.
-U.S.
-policy+Real: U.S.
-federal tax policy for the leader with behavior cloning for the followers.
-Results Analysis
-Table 1 presents the performance metrics for SMFRL and five baselines, alongside their optimal values, averaged over 10 evaluations of models trained with at least five seeds.
-SMFRL notably outperforms the baselines in achieving higher leader’s payoff, demonstrating its effectiveness in optimizing the leader’s objective.
-The DDPG+Real method, using a real-data trained policy, attains a substantial payoff but exhibits a significant gap from the households’ optimal response, as indicated by its higher exploitability.
-In contrast, SMFRL significantly reduces exploitability for both $N=100$ and $N=1000$, suggesting that its policy for households closely approximates the best response to the leader’s strategy, which is crucial for solving Stackelberg mean field games.
-Moreover, SMFRL closely approaches the optimal social welfare value, indicating that its solutions are the most socially optimal among the considered baselines.
-In summary, the SMFRL algorithm effectively learns policies that maximize the leader’s objective while ensuring that followers attain the best response to the leader’s policy, closely approaching the social optimum.
-The training curves are presented in Figure 4, and the test results for macroeconomic indicators are detailed in Table 1 in Appendix D.
-5.2 Macroeconomic Policy-Making
-Table 2: Test results 6 tax policies across five macroeconomic indicators at two numbers of followers, $N=100$ and $N=1000$.
-Households’ Policies Macroeconomic Indicators
-Number Per Capita GDP Social Welfare Income Gini Wealth Gini Years
-$\mathbf{N=100}$ Free Market 1.37e+05 32.97 0.89 0.92 1.10
-Saez Tax 2.34e+12 73.82 0.21 0.38 300.00
-US Federal Tax 4.88e+11 94.19 0.40 0.40 289.55
-AI Economist-BC 4.24e+12 97.24 0.54 0.52 299.55
-AI Economist 1.26e+05 72.81 0.88 0.91 1.00
-SMFG 9.59e+12 96.87 0.52 0.51 300.0
-$\mathbf{N=1000}$ Free Market 1.41e+05 334.79 0.90 0.93 1.00
-Saez Tax 6.35e+11 498.88 0.68 0.73 100.578
-US Federal Tax 1.41e+05 351.17 0.89 0.93 1.00
-AI Economist-BC N/A N/A N/A N/A N/A
-AI Economist N/A N/A N/A N/A N/A
-SMFG 1.10e+13 968.94 0.54 0.53 300.00
-Metrics and Baselines
-In addressing macroeconomic policymaking, we utilize classical macroeconomic indicators from the real world to evaluate the performance of different policies.
-Per Capita GDP indicates the level of economic development, while Income Gini and Wealth Gini measure the inequality in households’ income and wealth, respectively; a lower Gini index suggests greater social equality.
-The Years metric represents the sustainable duration of an economy, capped at 300 years; economies may terminate earlier due to excessive inequality or inadequate productivity to meet the populace’s needs.
-We compare several methods for solving macroeconomic policies: the Free Market [5] policy entails no government intervention; the Saez Tax [50] represents a well-known economic approach, often advocated for specific tax reforms in practice.
-The U.S.
-Federal Tax is the real policy from the OECD data  [^2] https://data.oecd.org/.
-In these approaches, households employ strategies based on behavior cloning from real data.
-Additionally, the AI Economist [62] represents an AI-based policy using a two-level Proximal Policy Optimization (PPO) framework.
-The AI Economist-BC method leverages behavior cloning policy on real-world data for pretraining.
-Results Analysis
-Table 2 presents the average values of macroeconomic indicators across multiple test rounds for different algorithms under varying numbers of households.
-The results reveal that the Saez Tax method achieves the lowest income Gini and wealth Gini when the number of households is small ($N=100$).
-However, our SMFG method yields higher per capita GDP and social welfare.
-At $N=1000$, the advantages of the SMFG algorithm become even more pronounced compared to other methods.
-It secures the longest duration (Years), the highest GDP, and the lowest income Gini and wealth Gini.
-The social welfare derived from the SMFG approach closely approaches the optimal value of 1000.
-This demonstrates that our method can learn the macroeconomic policy that simultaneously optimizes economic growth, social equity, and social welfare, surpassing existing economic approaches like the Saez Tax and AI-based policies (i.e.
-AI Economist).
-We provide the corresponding training curves in Figure 5 in Appendix D.
-5.3 Effectiveness Analysis in Diverse Households
+TaxAI is a simulation environment based on real data designed to study optimal taxation policies.
+It simulates the dynamic interactions between the government and large-scale households, to promote economic growth (i.e., GDP) as the objective.
+Tax Policies Per Capita GDP Social Welfare Income Gini Wealth Gini Years
+$100$ $1000$ $100$ $1000$ $100$ $1000$ $100$ $1000$ $100$ $1000$
+Free Market 1.37e+05 1.41e+05 32.97 334.79 0.89 0.90 0.92 0.93 1.10 1.00
+Saez Tax 2.34e+12 6.35e+11 73.82 498.88 0.21 0.68 0.38 0.73 300.00 100.58
+US Federal Tax 4.88e+11 1.41e+05 94.19 351.17 0.40 0.89 0.40 0.93 289.55 1.00
+AI Economist-BC 4.24e+12 N/A 97.24 N/A 0.54 N/A 0.52 N/A 299.55 N/A
+AI Economist 1.26e+05 N/A 72.81 N/A 0.88 N/A 0.91 N/A 1.00 N/A
+SMFG 9.59e+12 1.10e+13 96.87 968.94 0.52 0.54 0.51 0.53 300.00 300.00
+Table 1: Assessment results of 6 macroeconomic policies for different numbers of households N=100 and N=1000.
+Figure 3: Dynamic response curves of 3 macroeconomic policies to economic shocks at step 100.
+The SMFG policy (dark blue line) exhibits the least fluctuation and the fastest recovery in key indicators, indicating superior dynamic response capabilities.
+Figure 4: Ablation studies of SMFG method based on macroeconomic indicators for households’ numbers N=100 and N=1000.
+Comparison of Macroeconomic Policies
+This section compares the tax policy derived from SMFG with existing taxation policies in two aspects: (1) Macroeconomic Policy Evaluation: We assess the performance of different policies using key macroeconomic indicators.
+(2) Dynamic Response: At step 100, we simulate an economic shock caused by a financial crisis to evaluate the dynamic response capabilities of the different tax policies.
+Baselines and Metrics
+We select four different types of tax policy baselines, including the Free Market (Backhouse 2005) policy, which entails no government intervention; the Saez Tax (Saez 2001), a well-known economic approach often advocated for specific tax reforms in practice; the U.S.
+Federal Tax, which is the actual policy derived from OECD data https://data.oecd.org/; and the AI Economist (Zheng et al.
+), a popular AI-based policy utilizing a two-level Proximal Policy Optimization (PPO) framework.
+Additionally, the AI Economist-BC method employs behavior cloning on real-world data for pretraining.
+In the TaxAI environment, we evaluate the performance of these policies using key macroeconomic indicators derived from real-world data.
+Per Capita GDP reflects the level of economic development, while Income Gini and Wealth Gini measure inequality in household income and wealth, respectively—a lower Gini index indicates greater social equality.
+The Years metric represents the sustainable duration of an economy, with a maximum cap of 300 years.
+Average Wealth, Income, and Consumption are crucial assessment metrics related to financial crises.
+Macroeconomic Policy Evaluation
+In the context of optimizing economic growth, our SMFG method achieves the highest per capita GDP, social welfare, and the longest duration of outcomes across both $N=100$ and $N=1000$ scenarios.
+Table 1 shows the average values of macroeconomic indicators across 10 test runs for different algorithms under varying household numbers.
+Although the Gini index for SMFG is higher than that of the Saez Tax when $N=100$, this is acceptable when optimizing GDP, as SMFG’s per capita GDP is nearly 4 times that of the Saez Tax.
+When $N=1000$, SMFG outperforms all other metrics, indicating that the advantages of our SMFG method increase with the number of households.
+Figure 7 in the Appendix provides the corresponding training curves.
+Dynamic Response
+We simulate an economic shock similar to a financial crisis in the TaxAI environment: at step 100, all households lose 50% of their wealth.
+Due to the short duration (Years) under the Free Market and AI economist strategies, we focus on comparing the Saez Tax, US Federal Tax, and SMFG policy.
+From Figure 3, we find that the SMFG policy (dark blue line) demonstrates the fastest dynamic response across all indicators: Social Welfare remains largely unaffected, while Average Income and Average Consumption recover to their original levels within about 5 steps.
+Wealth Gini and Average Wealth return to their previous levels within about 15 steps.
+This level of recovery is unmatched by any other tax policy.
 Figure 5: The left subfigure illustrates a scenario with heterogeneous households adopting various policies.
 The right subfigure presents test results for households’ decisions (consumption, labor), microeconomic indicators (utility, wealth, income), and macroeconomic indicators (GDP) when there are different proportions of SMFG households ($N=100$).
+Ablation Studies of SMFG
+We conduct ablation studies to analyze the impact of the leader-follower update and mean field approximation modules in the SMFG method on macroeconomic indicators, as well as on convergence and social welfare.
+Baselines
+We conduct comparisons based on the SMFG method by excluding the leader-follower update (denoted as S) and the mean field approximation (denoted as MF) individually:
+1.
+SMFG-S: SMFG without leader-follower update.
+The government agent uses the independent DDPG (de Witt et al.
+2020), while interactions among large-scale households are modeled as a mean field game and trained using the mean-field MARL algorithm (Yang et al.
+2018).
+2.
+SMFG-MF: SMFG without mean field approximation.
+Both the government and household agents are trained using a variant of MADDPG algorithm (Lowe et al.
+2017), with the government making decisions first, followed by households.
+3. SMFG-S-MF: SMFG without both leader-follower update and mean field approximation. In this case, both the government and household agents use independent DDPG.
+Macroeconomic Indicators Analysis
+Figure 4 presents the results of the ablation studies on various macroeconomic indicators, evaluated across five seeds.
+The height of each bar represents the mean of the macroeconomic indicators, while the error bars indicate the variance.
+Our findings reveal that: (1) Excluding either module leads to a significant performance drop compared to the full SMFG method, with this effect becoming more pronounced as the number of agents $N$ increases.
+This underscores the critical importance of both the S and MF modules.
+(2) The SMFG-S variant outperforms SMFG-MF and SMFG-S-MF in terms of higher Per Capita GDP, social welfare, and Years, as well as lower Gini index.
+This suggests the dimensionality reduction of the MF module is crucial for the SMFG method’s overall effectiveness.
+(3) The low variance of the SMFG method indicates its stability.
+Convergence and Social Welfare
+Table 2 presents the performance of different baselines across three game theory-related metrics: leader’s payoff, social welfare, and exploitability.
+Leader’s payoff reflects the performance of the leader agent (i.e., the government), social welfare corresponds to the total utility of all households, and exploitability measures the distance between current policy and Nash equilibrium.
+The methods for calculating the metrics and optimal values are provided in Appendix D.
+From Table 2, we find: (1) SMFG achieves the highest leader’s payoff, the highest social welfare, and near-zero exploitability, indicating that the SMFG policy is closest to both the equilibrium and the maximal social welfare.
+(2) The high exploitability exhibited by SMFG-S indicates that the leader-follower update module facilitates the algorithm’s convergence toward equilibrium.
+(3) SMFG-MF outperforms SMFG-S-MF in terms of the leader’s payoff, suggesting that the leader-follower update module aids in optimizing the leader’s objective.
+However, it is challenging for SMFG-MF to optimize the followers’ social welfare.
+Leader’s Exploitability Social
+Payoff Welfare
+[…]
+Optimal Value \ \ 0. 0. 100 1000
+SMFG-S-MF -774 -716 0.652 1.023 83 859
+SMFG-S -448 -856 3.161 1.213 82 782
+SMFG-MF -535 -441 0.782 0.725 54 499
+SMFG (ours) 3294 3376 0.002 0.023 98 971
+Table 2: Ablation studies of the SMFG method based on game theory metrics for N=100 and N=1000. Optimal values are provided for reference.
+Effectiveness Analysis of SMFG
 Problem Description
-The above experiments have already demonstrated the effectiveness of our method in addressing macroeconomic policy-making problems.
-However, when extending SMFG approach to real-world scenarios, where governments and households implement leader and follower policies trained via the SMFG method (referred to as SMFG government and SMFG households), there remain partial households that might not adopt the SMFG follower policy.
-In economic terms, this scenario can be interpreted as the government introducing new policies and providing recommendations to households, some of whom may choose not to follow these recommendations.
-When faced with heterogeneous households that adopt varying strategies, the question arises: is our SMFG method still effective?
-In experiments, to model households that do not adopt the SMFG policy, referred to as Real-data Households, we utilize behavior cloning policy trained by real data.
-Related experimental details are shown in Appendix D.4.
+The above experiments have shown the effectiveness of the SMFG method in macroeconomic policymaking.
+Assuming the government is willing to implement macroeconomic policies learned by the SMFG method, we aim to explore the impact of non-adoption by some households.
+In our experiments, we introduce Non-SMFG Households whose policies are trained using a behavior cloning algorithm based on real data, reflecting the decision-making styles of real-world households.
+Details of these behavior cloning experiments are detailed in Appendix E.
+Additionally, we denote governments and households that adopt the SMFG policy as SMFG Government and SMFG Households, respectively.
 We set different proportions of SMFG households (0%, 25%, 50%, 75%, and 100%) and evaluate our SMFG method using microeconomic and macroeconomic indicators.
-In Figure 3, each subplot displays average values for two types of households (bars on the left Y-axis) and their aggregate values (dots on the right Y-axis).
+The results is shown in Figure 5, each subplot displays average values for two types of households (bars on the left Y-axis) and their aggregate values (dots on the right Y-axis).
 GDP, a macroeconomic indicator, is shown only in aggregate, not for individual households.
-Result Analysis
-From Figure 3, we identify several key findings: (1) SMFG Households outperform Real-data Households across microeconomic indicators such as wealth, income, and utility, which should encourage more households to adopt the SMFG policy; (2) The presence of heterogeneous households does not prevent SMFG Households from achieving near-optimal utility, capped at 1 in the TaxAI environment; (3) The lower labor contribution from Real-data Households leads to a decrease in Per Capita GDP as their proportion increases.
-Thus, the more households adopt the SMFG follower policy, the more effective the SMFG method becomes.
+Results
+From Figure 5, we observe several intriguing findings: (1) SMFG Households outperform Non-SMFG Households across microeconomic indicators such as wealth, income, and utility, suggesting that the policy of SMFG households is more attractive and likely to encourage broader adoption; (2) The presence of Non-SMFG Households does not prevent SMFG Households from achieving near-optimal utility, which is capped at 1 in the TaxAI environment; (3) As the proportion of SMFG Households increases, the per capita GDP also rises, indicating that the government should encourage households to adopt the SMFG policy.
 6 Conclusion
-This paper innovatively models macroeconomic policymaking as a dynamic Stackelberg mean field game, with the government acting as the leader and large-scale households as followers.
-We introduce a data-driven Stackelberg mean field reinforcement learning algorithm that effectively solves SMFGs.
-When applied to macroeconomic policy formulation, our method significantly outperforms existing economic approaches and AI-based policies, especially in scenarios involving large numbers of households.
-Furthermore, we demonstrate the robustness of our approach, showing that it retains effectiveness even when some households do not adopt the SMFG policy.
-In conclusion, this paper contributes an effective approach to modeling and solving optimal macroeconomic policies in the field of AI for economics and AI for social good.
+Addressing issues in macroeconomic policymaking, this paper innovatively proposes the SMFG method, which includes dynamic SMFG for modeling and SMFRL algorithm for solving.
+The SMFG method optimizes government’s policy by considering the dynamic responses of large-scale households to policy shifts, ultimately deriving optimal policies for both government and households.
+Experimental results demonstrate the SMFG method’s superior performance, stability, and convergence over existing approaches.
+When extended to real-world scenarios, even if some households do not adopt the SMFG policy, their wealth and utility are not as high as those who do, encouraging more households to adopt the SMFG policy.
+In conclusion, this paper contributes an effective approach to modeling and solving optimal macroeconomic policies in the fields of AI for economics and AI for social impact.
 Appendix A SMFRL algorithm pseudocode
-Algorithm 1 Stackelberg Mean Field Reinforcement Learning (SMFRL)
-  Initialize ${Q}_{\phi^{l}},{Q}_{\phi^{l}_{-}},{Q}_{\phi^{f}},{Q}_{\phi^{f}_{-}},\pi_{\theta^{l}},\pi_{\theta^{l}_{-}},\pi_{\theta^{f}},\pi_{\theta^{f}_{-}}$.
-  for epoch = 1 to M do
-     Receive initial state $\mathbf{s}_{t}=\{s^{l}_{t},s^{f1}_{t},...,s^{fN}_{t}\}$.
-     for  $t=1$ to max-epoch-length do
-        For the leader agent, select action $a^{l}_{t}=\pi_{\theta^{l}}(s^{l}_{t})+\mathcal{N}_{t}$; for each follower agent $i$, select action $a^{fi}=\pi_{\theta^{f}}(s^{fi}_{t},a^{l}_{t})+\mathcal{N}_{t}$ w.r.t.
-the current policy and exploration.
-        Execute actions $\mathbf{a}_{t}=\{a^{l}_{t},a^{f1}_{t},...,a^{fN}_{t}\}$ and observe reward $\mathbf{r}_{t}=\{r_{t}^{l},r^{f1}_{t},...,r^{fN}_{t}\}$ and next state $\mathbf{s}_{t+1}$.
-        Store $(\mathbf{s}_{t},\mathbf{s}_{t+1},\mathbf{a}_{t},\mathbf{r}_{t},L^{f}_{t})$ in the replay buffer $\mathcal{D}$.
-        $\mathbf{s}_{t}\leftarrow\mathbf{s}_{t+1}$.
-        for j =1 to update-cycles do
-           Sample a random minibatch of $N_{b}$ samples from $\mathcal{D}$.
-           For each follower agent, perform inner updates:
-           for k = 1 to inner-update-cycles do
-              Set
-$y^{f}_{t}=r^{f}_{t}+\gamma{Q}_{\phi^{f}_{-}}(s^{f}_{t+1},a^{f}_{t+1},a^{l}_{t+1},L^{f}_{t+1}).$
-              Update follower’s critic by minimizing the loss:
-$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\mathcal{D}}\Big{[}\left(y^{f}_{t}-{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{f}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\Big{]}.$
-              Update follower’s actor by policy gradient:
-$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L^{f}_{t}\sim\mathcal{D}}\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{-},a^{l}_{t},L^{f})|_{a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})}\Big{]}.$
-           end for
-           After follower updates, update the leader’s networks:
-           Set
-$y^{l}_{t}=r_{t}^{l}+\gamma{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L^{f}_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})}.$
-           Update leader’s critic net by minimizing the loss:
-$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{l}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right]$
-           Update leader’s actor by sampled policy gradient:
-$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{-},L^{f}_{t})|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}\right].$
-        end for
-        Update target network parameters:
-$\begin{gathered}\phi_{-}^{l}\leftarrow\tau_{\phi}\phi^{l}+\left(1-\tau_{\phi}\right)\phi_{-}^{l}\\ \theta_{-}^{l}\leftarrow\tau_{\theta}\theta^{l}+\left(1-\tau_{\theta}\right)\theta_{-}^{l}\\ \phi_{-}^{f}\leftarrow\tau_{\phi}\phi^{f}+\left(1-\tau_{\phi}\right)\phi_{-}^{f}\\ \theta_{-}^{f}\leftarrow\tau_{\theta}\theta^{f}+\left(1-\tau_{\theta}\right)\theta_{-}^{f}\end{gathered}$
-     end for
-  end for
 Appendix B Assumptions and Limitations
@@
 Addressing these limitations will provide further insights applicable to real-world scenarios.
-Appendix C Impact Statements
-This paper presents work whose goal is to advance the fields of AI for social science and AI for economics.
-Our work aims to offer suggestions and references for governments and the people, yet it must not be rashly applied to the real world.
-There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
-Appendix D Additional Results
-Compute Resources
-All experiments are run on 2 workstations: A 64-bit server with dual AMD EPYC 7742 64-Core Processors @2.25 GHz, 256 cores, 512 threads, 503GB RAM, and 2 NVIDIA A100-PCIE-40GB GPU.
-A 64-bit workstation with Intel Core i9-10920X CPU @ 3.50GHz, 24 cores, 48 threads, 125 GB RAM, and 2 NVIDIA RTX2080 Ti GPUs.
-The following Table 3 shows the approximate training times for several algorithms.
-Table 3: the approximate training times for several algorithms for solving dynamic SMFGs.
-Algorithm N=100 (hours) N=1000 (hours)
-SMFRL 4 14
-MADDPG 3 16
-MF-MARL 4 9
-IDDPG 2 6
-DDPG+Real 1 5
-Rule-based+Real 0.25 2
-D.1 Training Curves for Dynamic SMFGs
-Due to the challenging nature of the task, the baselines we compared did not perform well in solving dynamic SMFGs.
-Therefore, we utilized behavior cloning as a pre-training strategy for the follower agents.
-Therefore, we compared SMFRL with 5 baselines in 4 different experimental setups: without behavior cloning as pre-training for follower agents at N=100 and N=1000 (marked as N=100 without BC and N=1000 without BC); with BC-based pre-training for follower agents at N=100 and N=1000 (N=100-BC; N=1000-BC).
-Figure 4 illustrates the training curves of 4 key macroeconomic indicators under these three settings.
-The solid line represents the average value of the metrics across the 5 random seeds, while the shaded area represents the standard deviation.
-Each row corresponds to one setting, and each column to a macroeconomic indicator, including per capita GDP, social welfare, income Gini, and wealth Gini.
-A rise in per capita GDP indicates economic growth, an increase in social welfare implies happier households and a lower Gini index indicates a fairer society.
-Each subplot’s Y-axis represents the indicators’ values, and the X-axis represents the training steps.
-Table 4 displays the test results of the 6 algorithms across 5 indicators, with each column corresponding to an experimental setting.
-Figure 4 and Table 4 present two experimental findings: (1) Using BC as a pre-training method for the follower’s policy enhances the algorithms’ stability and performance.
-Comparing settings with and without BC (the first two rows), our method, SMFRL, shows similar convergence outcomes; however, the performance of other algorithms significantly improves across all four indicators with BC-based pre-training.
-Furthermore, the training curves of each algorithm are more stable.
-(2) The SMFRL algorithm substantially outperforms other algorithms in solving SMFGs, both in large-scale followers and without pre-training scenarios.
-In the setting of N=100-BC, SMFRL achieved a significantly higher per capita GDP compared to other algorithms, while its social welfare and Gini index are similar to others, essentially reaching the upper limit.
-Besides, in N=100 without BC and N=1000-BC, SMFRL consistently obtains the most optimal solutions across all indicators.
-Table 4: Performance metrics of six algorithms across different settings on key economic indicators.
-Settings Algorithm GDP per Capita Social Welfare Income Gini Wealth Gini Years
-N=100 without BC Rule-based+Random 1.41e+05 69.27 0.89 0.92 1.00
-DDPG+Random 1.41e+05 70.91 0.88 0.92 1.00
-MADDPG 4.93e+04 55.74 0.91 0.92 1.00
-IDDPG 1.21e+05 83.09 0.88 0.92 1.00
-MF-MARL 8.66e+07 82.02 0.82 0.83 75.75
-SMFRL 9.59e+12 96.87 0.52 0.51 300.00
-N=1000 without BC Rule-based+Random 1.42e+05 705.85 0.90 0.93 1.00
-DDPG+Random 1.42e+05 654.37 0.90 0.93 1.00
-MADDPG 3.37e+09 645.38 0.83 0.85 63.58
-IDDPG 1.33e+05 874.53 0.89 0.92 1.00
-MF-MARL 1.31e+05 834.93 0.88 0.92 1.50
-SMFRL 1.10e+13 968.94 0.54 0.53 300.00
-N=100 with BC Rule-based+Real 3.66e+11 79.23 0.52 0.53 217.45
-DDPG+Real 2.03e+12 94.50 0.46 0.48 299.85
-MADDPG 6.38e+12 93.89 0.57 0.58 268.53
-IDDPG 7.41e+12 98.16 0.53 0.55 300.00
-MF-MARL 5.44e+12 98.21 0.50 0.52 300.00
-SMFRL 1.01e+13 96.90 0.51 0.53 299.89
-N=1000 with BC Rule-based+Real 1.41e+05 334.79 0.90 0.93 1.00
-DDPG+Real 4.92e+11 527.09 0.75 0.79 100.68
-MADDPG 6.82e+12 954.88 0.56 0.62 278.50
-IDDPG 2.79e+12 512.19 0.77 0.81 100.68
-MF-MARL 1.13e+05 440.00 0.90 0.93 1.00
-SMFRL 9.68e+12 975.15 0.52 0.51 300.00
-Figure 4: The training curves for 6 algorithms on 4 macroeconomic indicators, comparing settings without behavior cloning as pre-train (N=100 & N=1000) and with behavior cloning (N=100-BC & N=1000-BC).
-D.2 Training Curves for Various Tax Policies
-We compare the performance of 6 policies across four economic indicators under two settings: with N=100 and N=1000 households.
-Figure 5 displays the training curves and Table 2 shows the test results in Appendix.
-Both Figure 5 and Table 2 indicate that the SMFG method significantly surpasses other policies in the task of optimizing GDP, and achieves the highest social welfare.
-When N=100, the Saez tax achieves the lowest income and wealth Gini coefficients, suggesting greater fairness.
-However, at N=1000, SMFG performs optimally across all economic indicators, while the effectiveness of other policies noticeably diminishes as the number of households increases.
-The Saez tax also reduces the Gini index, but not as effectively or stably as the SMFG.
-Figure 5: The training curves for 6 tax policies on 5 macroeconomic indicators (N=100 & N=1000 with BC)
-D.3 Efficiency-Equity Tradeoff of Policies
-Figure 6: Comparative performance of various policies under multi-objective assessment (Efficiency-Equity).
-In economics, the Efficiency-Equity Tradeoff is a highly debated issue.
-We find that our SMFG method is optimal in balancing efficiency-equity, except in cases of extreme concern for social fairness.
-In our study, we depict the economic efficiency (Per capita GDP) on the Y-axis and equity (wealth Gini) on the X-axis of Figure 6(a) for various policies.
-Different policies are represented by circles of different colors, with their sizes proportional to social welfare.
-Different circles of the same color correspond to different seeds.
-Figure 6 (a) shows that the wealth Gini indices for SMFG and AI Economist-BC are similar, but SMFG has a higher GDP, suggesting its superiority over AI Economist-BC.
-SMFG significantly outperforms the free market policy and AI Economist due to its higher GDPs and lower wealth Ginis.
-However, comparing SMFG with the Saez tax and the U.S.
-Federal tax policy in terms of both economic efficiency (GDP) and social equity (Gini) is challenging.
-Therefore, we introduce Figure 6 (b) to demonstrate the performance of different policies under various weights in a multi-objective assessment.
-In Figure 6 (b), the Y-axis shows the weighted values of the multi-objective function $Y=log(\text{per capita GDP})+\alpha(\text{wealth Gini})$, and the X-axis represents the weight of the wealth Gini index.
-For each weight $\alpha$, we compute the multi-objective weighted values for those policies, represented as circles of different colors.
-Due to the logarithmic treatment of GDP in (b), when $\alpha=10$, the overall objective focuses solely on social fairness; when $\alpha=0$, the overall objective is concerned only with efficiency.
-Our findings in (b) reveal that only when $\alpha\geq 8$, which indicates a substantial emphasis on social equity, does the Saez tax outperform SMFG.
-However, SMFG consistently proves to be the most effective policy under a wide range of preference settings.
-D.4 Behavior Cloning Experiments
-We conduct behavior cloning based on real data to simulate the behavior strategies of households in realistic scenarios, which are then used in Experiment 5.3 to compare with SMFG Households.
-We use the statistical data from the 2022 Survey of Consumer Finances (SCF)  [^3] https://www.federalreserve.gov/econres/scfindex.htm.
-Details about the behavior cloning experiments are provided below.
-This experiment conducts behavior cloning on networks for four different household policies: Multilayer Perceptron (MLP), AI economist’s network (MLP+LSTM+MLP), MF-MARL, and MADDPG network.
-The first two, as their network outputs, are probability distributions, use negative log-likelihood loss (Figure 7 left); the latter two’s networks employ deterministic policies, hence they use mean square error loss against real data (Figure 7 right).
-The loss convergence curve of behavior cloning is shown in Figure 7.
-It can be observed that the AI economist’s network, due to its complexity, struggles to converge to near -1 like MLP.
-The losses corresponding to MFRL and MADDPG can converge to below 0.1.
-Figure 7: The behavior cloning loss for 4 networks in two loss types.
-Appendix E Saez tax
+Appendix C Saez tax
 The Saex tax policy is often considered a suggestion for specific tax reforms in the real world.
-The specific calculation method is as follows [50].
+The specific calculation method is as follows (Saez 2001).
 The Saez tax utilizes income distribution $f(z)$ and cumulative distribution $F(z)$ to get the tax rates.
@@
 where $Z=\sum_{i}z_{i}$ when tax rates is $\tau$.
-Appendix F Theoretical Proof
-Proposition F.1.
-The action value functions $Q_{i}$ and ${Q}_{t}$ are equivalent, i.e.
-$Q_{i}(\mathbf{s}_{t},\mathbf{a}_{t})={Q}_{t}(s_{t}^{l},a_{t}^{l},L_{t}^{f})$, $\forall\mathbf{s}_{t}\in\mathcal{S}^{l}\times\{\mathcal{S}^{f}\}^{N}$, $\forall\mathbf{a}_{t}\in\mathcal{A}^{l}\times\{\mathcal{A}^{f}\}^{N}$, $\forall L_{t}^{f}\in\mathbb{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$.
-Proof.
-We commence by defining a partial order relation $\prec$ on $\mathcal{S}^{f}\times\mathcal{A}^{f}$.
-For any given vector $\mathbf{v}$, we construct a sorted counterpart $\tilde{\mathbf{v}}$, which maintains the same elements as $\mathbf{v}$ but orders them in ascending sequence based on $\prec$.
-For example, if $\mathbf{v}=\{(s_{1},a_{1}),(s_{2},a_{2}),(s_{3},a_{3})\}$, then $\tilde{\mathbf{v}}$ is arranged as $\{(s_{(1)},a_{(1)}),(s_{(2)},a_{(2)}),(s_{(3)},a_{(3)})\}$, where $(s_{(1)},a_{(1)})\prec(s_{(2)},a_{(2)})\prec(s_{(3)},a_{(3)})$).
-Exploiting the homogeneity of the follower agents, we can equate $Q(\mathbf{s}_{t},\mathbf{a}_{t})$ with $Q(s^{l}_{t},a^{l}_{t},\tilde{\mathbf{v}})$, where $\mathbf{v}$ encapsulates the joint state-action pairs of the followers, and $\tilde{\mathbf{v}}$ represents its sorted form.
-We further define $\mathcal{\mathbf{v}}$ as the collection of all such sorted vectors, and $\mathcal{L}$ as the ensemble of mean-field distributions.
-The empirical distribution of the followers’ state-action pairs, denoted as $L:\mathcal{V}\rightarrow\Delta(\mathcal{V})$, is expressed as $L^{f}_{t}(s,a)=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}_{\{(\tilde{s}_{i},\tilde{a}_{i})=(\tilde{s},\tilde{a})\}}$.
-The bijectiveness of the mapping $L$, ensured by the sorting process, leads to the conclusion that $Q_{t}(\mathbf{s}_{t},\mathbf{a}_{t})=Q_{t}(s^{l}_{t},a^{l}_{t},\mathbf{v})=Q_{t}(s^{l}_{t},a^{l}_{t},\tilde{\mathbf{v}})={Q}_{t}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$.
-∎
-Appendix G Economic Model Details
+Appendix D Game Theory Metrics
+We will utilize the following metrics related to game theory to evaluate the effectiveness of the leader and follower policies: (1) The leader’s payoff, which indicates the performance of the leader’s policy in optimizing the leader’s objective; (2) Exploitability, which measures the deviation of the agent’s policy from the best response; (3) Social welfare, which assesses the deviation of the current state from the social optimum.
+Leader’s Payoff
+We define the leader’s payoff using the long-term expected rewards of the leader’s policy $\pi^{l}$ over discrete timesteps, as detailed in Equation 1.
+Exploitability
+Exploitability is a critical metric in evaluating the convergence of policies and quantifying the divergence from the best response strategy in game theory.
+For a follower, exploitability $\mathcal{E}^{f}(\pi^{f};\pi^{l})$ is defined as the difference in payoffs between the follower’s actual policy $\pi^{f}$ and its optimal response ${\pi^{f}}^{*}$, given the leader’s policy $\pi^{l}$.
+Formally, it is represented as:
+$\mathcal{E}^{f}(\pi^{f};\pi^{l})=J^{f}(\pi^{l},{\pi^{f}}^{*},L)-J^{f}(\pi^{l},\pi^{f},L),$
+where $J^{f}$ denotes the cumulative reward for the follower, defined in Section 3.
+Similarly, the leader’s exploitability $\mathcal{E}^{l}(\pi^{l};\pi^{f})$ measures the payoff difference between the leader’s policy $\pi^{l}$ and its best response ${\pi^{l}}^{*}$, given the followers’ response policy $\pi^{f}$ and state-action distribution $L$.
+This is given by:
+$\mathcal{E}^{l}(\pi^{l};\pi^{f})=J^{l}(\pi^{l*},\pi^{f},L)-J^{l}(\pi^{l},\pi^{f},L),$
+with $J^{l}$ representing the cumulative reward for the leader, and $L=\{L_{t}\}_{t=0}^{T}$ detailing the state-action distribution for followers over time (see Section 3).
+The overall exploitability, which measures the discrepancy from Nash equilibrium for both the leader and the followers, is defined as:
+$\mathcal{E}(\pi^{l},\pi^{f})=\mathcal{E}^{f}(\pi^{f};\pi^{l})+\mathcal{E}^{l}(\pi^{l};\pi^{f}),$
+A near-zero value of $\mathcal{E}(\pi^{l},\pi^{f})$ indicates that the policies of both the leader and the followers are approaching their respective optimal strategies ${\pi^{l}}^{*}$ and ${\pi^{f}}^{*}$, signifying an equilibrium state.
+Social Optimum and Social Welfare
+In economic theory, the Social Optimum describes a state in which the allocation of resources achieves maximum efficiency, as measured by social welfare (Arrow and Debreu 1954; Malul and Bar-El 2009).
+Given the leader’s policy $\pi^{l}$ and the representative follower’s policy $\pi^{f}$ among large-scale followers, social welfare $\mathcal{SW}(\pi^{l},\pi^{f})$ is approximately calculated as the sum of the utility functions defined in Section 3 of the $N$ followers:
+$\mathcal{SW}(\pi^{l},\pi^{f})$ $=\sum^{N}_{i=1}J^{fi}\left(\pi^{l},\pi^{f},L\right)$
+$=\mathbb{E}_{s_{0}^{f}\sim\mu_{0}^{f},\mathbf{s}_{t+1}\sim P}\left[\sum_{t=0}^{T}\sum^{N}_{i=1}r^{fi}(\mathbf{s}_{t},a^{l}_{t},a^{fi}_{t},L_{t})\right]$
+Appendix E Additional Results
+Compute Resources
+All experiments are run on 2 workstations: A 64-bit server with dual AMD EPYC 7742 64-Core Processors @2.25 GHz, 256 cores, 512 threads, 503GB RAM, and 2 NVIDIA A100-PCIE-40GB GPU.
+A 64-bit workstation with Intel Core i9-10920X CPU @ 3.50GHz, 24 cores, 48 threads, 125 GB RAM, and 2 NVIDIA RTX2080 Ti GPUs.
+The following Table 3 shows the approximate training times for several algorithms.
+Algorithm N=100 (hours) N=1000 (hours)
+SMFG 4 14
+SMFG-MF 3 16
+SMFG-S 4 9
+SMFG-S-MF 2 6
+Free Market 0.25 2
+Saez Tax 4 23
+AI Economist 6.5 N/A
+Table 3: The approximate training times for baselines in our experiments.
+Experimental Results of Ablation Studies
+In this section, we present additional experimental results from ablation studies, including training curves 6 and a test table 4, as well as experiments incorporating the use of behavior cloning as a pre-training strategy for follower agents.
+We find that the SMFG method without behavior cloning as pre-training still surpasses other baselines that utilize behavior cloning.
+More specifically, we compared SMFG with 5 baselines across 4 different experimental setups: without behavior cloning as pre-training for follower agents at N=100 and N=1000 (marked as N=100 without BC and N=1000 without BC); with BC-based pre-training for follower agents at N=100 and N=1000 (N=100-BC; N=1000-BC).
+Figure 6 illustrates the training curves of 4 key macroeconomic indicators under these four settings.
+The solid line represents the average value of the metrics across the 5 random seeds, while the shaded area represents the standard deviation.
+Each row corresponds to one setting, and each column to a macroeconomic indicator, including per capita GDP, social welfare, income Gini, and wealth Gini.
+A rise in per capita GDP indicates economic growth, an increase in social welfare implies happier households and a lower Gini index indicates a fairer society.
+Each subplot’s Y-axis represents the indicators’ values, and the X-axis represents the training steps.
+Table 4 displays the test results of the 6 algorithms across 5 indicators, with each column corresponding to an experimental setting.
+Figure 6 and Table 4 present two experimental findings: (1) Using BC as a pre-training method for the follower’s policy enhances the algorithms’ stability and performance.
+Comparing settings with and without BC (the first two rows), our method, SMFG, shows similar convergence outcomes; however, the performance of other algorithms significantly improves across all four indicators with BC-based pre-training.
+Furthermore, the training curves of each algorithm are more stable.
+(2) The SMFG method substantially outperforms other algorithms in solving SMFGs, both in large-scale followers and without pre-training scenarios.
+In the setting of N=100-BC, SMFG achieved a significantly higher per capita GDP compared to other algorithms, while its social welfare and Gini index are similar to others, essentially reaching the upper limit.
+Besides, in N=100 without BC and N=1000-BC, SMFG consistently obtains the most optimal solutions across all indicators.
+Settings Algorithm Per Capita GDP Social Welfare Income Gini Wealth Gini Years
+N=100 without BC Rule-based+Random 1.41e+05 69.27 0.89 0.92 1.00
+DDPG+Random 1.41e+05 70.91 0.88 0.92 1.00
+SMFG-MF 1.23e+05 48.17 0.89 0.93 1.00
+SMFG-S-MF 1.21e+05 83.09 0.88 0.92 1.00
+SMFG-S 8.66e+07 82.02 0.82 0.83 75.75
+SMFG 9.59e+12 96.87 0.52 0.51 300.00
+N=1000 without BC Rule-based+Random 1.42e+05 705.85 0.90 0.93 1.00
+DDPG+Random 1.42e+05 654.37 0.90 0.93 1.00
+SMFG-MF 1.48e+05 499.01 0.89 0.93 1.02
+SMFG-S-MF 1.33e+05 874.53 0.89 0.92 1.00
+SMFG-S 1.31e+05 834.93 0.88 0.92 1.50
+SMFG 1.10e+13 968.94 0.54 0.53 300.00
+N=100 with BC Rule-based+Real 3.66e+11 79.23 0.52 0.53 217.45
+DDPG+Real 2.03e+12 94.50 0.46 0.48 299.85
+SMFG-MF 6.38e+12 93.89 0.57 0.58 268.53
+SMFG-S-MF 7.41e+12 98.16 0.53 0.55 300.00
+SMFG-S 5.44e+12 98.21 0.50 0.52 300.00
+SMFG 1.01e+13 96.90 0.51 0.53 299.89
+N=1000 with BC Rule-based+Real 1.41e+05 334.79 0.90 0.93 1.00
+DDPG+Real 4.92e+11 527.09 0.75 0.79 100.68
+SMFG-MF 6.82e+12 954.88 0.56 0.62 278.50
+SMFG-S-MF 2.79e+12 512.19 0.77 0.81 100.68
+SMFG-S 1.13e+05 440.00 0.90 0.93 1.00
+SMFG 9.68e+12 975.15 0.52 0.51 300.00
+Table 4: Performance metrics of six algorithms across different settings on key macroeconomic indicators.
+Figure 6: The training curves for 6 algorithms on 4 macroeconomic indicators, comparing settings without behavior cloning as pre-train (N=100 & N=1000) and with behavior cloning (N=100-BC & N=1000-BC).
+Training Curves for Various Tax Policies
+We compare the performance of 6 policies across four economic indicators under two settings: with N=100 and N=1000 households.
+Figure 7 displays the training curves and Table 1 shows the test results.
+Both Figure 7 and Table 1 indicate that the SMFG method significantly surpasses other policies in the task of optimizing GDP, and achieves the highest social welfare.
+When N=100, the Saez tax achieves the lowest income and wealth Gini coefficients, suggesting greater fairness.
+However, at N=1000, SMFG performs optimally across all economic indicators, while the effectiveness of other policies noticeably diminishes as the number of households increases.
+The Saez tax also reduces the Gini index, but not as effectively or stably as the SMFG.
+Figure 7: The training curves for 6 tax policies on 5 macroeconomic indicators (N=100 & N=1000 with BC)
+Efficiency-Equity Tradeoff of Policies
+Figure 8: Comparative performance of various policies under multi-objective assessment (Efficiency-Equity).
+In economics, the Efficiency-Equity Tradeoff is a highly debated issue.
+We find that our SMFG method is optimal in balancing efficiency-equity, except in cases of extreme concern for social fairness.
+In our study, we depict the economic efficiency (Per capita GDP) on the Y-axis and equity (wealth Gini) on the X-axis of Figure 8(a) for various policies.
+Different policies are represented by circles of different colors, with their sizes proportional to social welfare.
+Different circles of the same color correspond to different seeds.
+Figure 8 (a) shows that the wealth Gini indices for SMFG and AI Economist-BC are similar, but SMFG has a higher GDP, suggesting its superiority over AI Economist-BC.
+SMFG significantly outperforms the free market policy and AI Economist due to its higher GDPs and lower wealth Ginis.
+However, comparing SMFG with the Saez tax and the U.S.
+Federal tax policy in terms of both economic efficiency (GDP) and social equity (Gini) is challenging.
+Therefore, we introduce Figure 8 (b) to demonstrate the performance of different policies under various weights in a multi-objective assessment.
+In Figure 8 (b), the Y-axis shows the weighted values of the multi-objective function $Y=log(\text{per capita GDP})+\alpha(\text{wealth Gini})$, and the X-axis represents the weight of the wealth Gini index.
+For each weight $\alpha$, we compute the multi-objective weighted values for those policies, represented as circles of different colors.
+Due to the logarithmic treatment of GDP in (b), when $\alpha=10$, the overall objective focuses solely on social fairness; when $\alpha=0$, the overall objective is concerned only with efficiency.
+Our findings in (b) reveal that only when $\alpha\geq 8$, which indicates a substantial emphasis on social equity, does the Saez tax outperform SMFG.
+However, SMFG consistently proves to be the most effective policy under a wide range of preference settings.
+Behavior Cloning Experiments
+We conduct behavior cloning based on real data to simulate the behavior strategies of households in realistic scenarios, which are then used in Experiment 5 to compare with SMFG Households.
+We collect the statistical data from the 2022 Survey of Consumer Finances (SCF) (https://www.federalreserve.gov/econres/scfindex.htm) as the real data buffer $\mathcal{D}_{real}$.
+Based on real data, we fetch a large number of followers’ state-action pairs $\{s^{f},a^{f}\}$ from a real-data buffer $\mathcal{D}_{real}$ for behavior cloning.
+For different settings of network structures, we have chosen two types of loss: when the neural network outputs a probability distribution of actions, we use the negative log-likelihood loss (NLL loss); when the neural network outputs action values, we employ the mean square error loss (MSE loss).
+Our goal is to find the optimal parameters $\theta$ as the follower’s policy network $\pi_{\theta}$ initialization, thereby minimizing the loss to its lowest convergence.
+$\min_{\theta}\mathcal{L}_{NLL}$ $=-\mathbb{E}_{s^{f},a^{f}\sim\mathcal{D}}\log{\pi_{\theta}(a^{f}\mid s^{f})},$
+$\min_{\theta}\mathcal{L}_{MSE}$ $=\mathbb{E}_{s^{f},a^{f}\sim\mathcal{D}}\left(a^{f}-a\right)^{2}|_{a=\pi_{\theta}(s^{f})}.$
+This experiment conducts behavior cloning on networks for four different household policies: Multilayer Perceptron (MLP), AI economist’s network (MLP+LSTM+MLP), SMFG-S, and SMFG-MF network.
+The first two, as their network outputs, are probability distributions, use negative log-likelihood loss (Figure 9 left); the latter two’s networks employ deterministic policies, hence they use mean square error loss against real data (Figure 9 right).
+The loss convergence curve of behavior cloning is shown in Figure 9.
+It can be observed that the AI economist’s network, due to its complexity, struggles to converge to near -1 like MLP.
+The losses corresponding to MFRL and SMFG-MF can converge to below 0.1.
+Figure 9: The behavior cloning loss for 4 networks in two loss types.
+Appendix F Economic Model Details
 Economic activities among households aggregate into labor markets, capital markets, goods markets, etc.
@@
 where $C_{t}=\sum_{i}^{N}c^{i}_{t}$ represents the total consumption of consumers, and $G_{t}$ is government spending. The supply, demand, and price represent the states of the market.
+Economic Shocks
+In Experiment 5, we simulate economic shocks analogous to a financial crisis: at the 100-th step, the wealth of all households is reduced by 50%.
+In our economic model, this scenario is mathematically represented as follows: for each household member, the wealth $a_{t}^{i}$ at time $t$ is updated according to the rule
+$a_{t}^{i}=0.5a_{t-1}^{i},\quad\forall i\in\{1,\ldots,N\}$
+where $N$ denotes the total number of household members.
+Appendix G Hyperparameters
+Hyperparameter Value
+Discount factor $\gamma$ 0.975
+Replay buffer size 1e6
+Num of epochs 1000
+Epoch length 300
+Batch size 128
+Adam epsilon 1e-5
+Update cycles 100
+Evaluation epochs 10
+Hidden size 128
+Tau 0.95
+Critic initial learning rate 3e-4
+Actor initial learning rate 3e-4
+Learning rate adjustment $0.95^{(\text{epoch}//35)}$
+Table 5: Hyperparameters of SMFG methods and its variants.
+Hyperparameter Value
+Noise rate 0.01
+Epsilon start 0.1
+Epsilon end 0.05
+Epsilon decay 1e-5
+Table 6: Hyperparameters of SMFG-MF algorithm different from SMFG method.
+Hyperparameter Value
+Tau $\tau$ 5e-3
+Gamma $\gamma$ 0.95
+Eps $\epsilon$ 1e-5
+Clip 0.1
+Vloss coef 0.5
+Ent coef 0.01
+Government’s initial learning rate 3e-4
+Learning rate adjustment 0, epoch $<$ 10
+$0.97^{(\text{epoch}//35)}$, epoch $\geq$ 10
+Households’ initial learning rate 1e-6
+Learning rate adjustment $0.97^{(\text{epoch}//35)}$
+Table 7: Hyperparameters of AI Economist Algorithm different from SMFG approach.
+Ethical Statement
+This research introduces a novel SMFG method, designed to optimize macroeconomic policies by modeling complex interactions at the micro level.
+The potential impact of this work extends across several domains:
+Academic Contributions
+The framework and algorithm proposed represent significant advancements in AI for economics and AI for social impact field, potentially serving as foundational tools for future research in macroeconomic policy making.
+By addressing the Lucas critique through dynamic modeling of individual agents within a mean field game, this work encourages more accurate and robust economic predictions and policy evaluations.
+Policy Making and Societal Impact
+By enabling the optimization of macroeconomic policies through real-time, dynamic responses of micro-agents, this model provides policymakers with a powerful tool for assessing the impact of different economic strategies, leading to more informed decisions that maximize social welfare and economic stability, particularly in response to economic shocks.
+The application of this model can have profound implications for wealth distribution and social equity, helping ensure that economic policies are beneficial to a broader section of the population, potentially reducing inequality and enhancing societal well-being.
+Ethical Considerations
+While the model aims to improve economic outcomes, the manipulation of macroeconomic policies must be approached with caution to avoid unintended negative consequences such as increased inequality or destabilization of economic sectors.
+Further, the reliance on AI-based decisions necessitates continuous scrutiny to ensure that the model accurately represents all population segments.
+Limitations and Risks
+The complexity of the models also introduces risks related to the oversimplification of real-world dynamics and potential biases in the simulation of economic responses.
+Continuous validation against empirical data and diverse economic scenarios is essential to ensure the reliability and ethical application of the proposed methods.
+In summary, the proposed SMFG framework and SMFRL algorithm hold the potential to significantly impact both academic research and practical policy making, offering a new perspective on dynamic economic modeling that prioritizes realistic, individual-level responses within large-scale economic systems.

### v4

@@
-Learning Macroeconomic Policies based on Microfoundations: A Stackelberg Mean Field Game Approach
+Learning Macroeconomic Policies through Dynamic Stackelberg Mean-Field Games
 Abstract
-The Lucas critique emphasizes the importance of considering microfoundations—how micro-agents (i.e., households) respond to policy changes—in macroeconomic policymaking.
-However, due to the vast scale and complex dynamics among micro-agents, predicting microfoundations is challenging.
-Consequently, this paper introduces a Stackelberg Mean Field Game (SMFG) approach that models macroeconomic policymaking based on microfoundations, with the government as the leader and micro-agents as dynamic followers.
-This approach treats large-scale micro-agents as a population, to optimize macroeconomic policies by learning the dynamic response of this micro-population.
-Our experimental results indicate that the SMFG approach outperforms real-world macroeconomic policies, existing AI-based and economic methods, enabling the learned macroeconomic policy to achieve the highest performance while guiding large-scale micro-agents toward maximal social welfare.
-Additionally, when extended to real-world scenarios, households that do not adopt the SMFG policy experience lower utility and wealth than adopters, thereby increasing the attractiveness of our policy.
-In summary, this paper contributes to the field of AI for economics by offering an effective tool for modeling and solving macroeconomic policymaking issues.
-Figure 1: The mapping from macroeconomic policymaking (left) to Dynamic Stackelberg Mean Field Games (right).
-On the left, macro-level government actions, including fiscal or monetary policies, aim to optimize GDP, welfare, or equity.
-The microfoundations reflect the economic behaviors of large-scale micro-agents responding to shifts in macroeconomic policy.
-In Dynamic SMFGs, the government acts as the leader, and micro-agents as followers, respond to the leader’s policy.
+Macroeconomic outcomes emerge from individuals’ decisions, making it essential to model how agents interact with macro policy via consumption, investment, and labor choices.
+We formulate this as a dynamic Stackelberg game: the government (leader) sets policies, and agents (followers) respond by optimizing their behavior over time.
+Unlike static models, this dynamic formulation captures temporal dependencies and strategic feedback critical to policy design.
+However, as the number of agents increases, explicitly simulating all agent–agent and agent–government interactions becomes computationally infeasible.
+To address this, we propose the Dynamic Stackelberg Mean Field Game (DSMFG) framework, which approximates these complex interactions via agent–population and government–population couplings.
+This approximation preserves individual-level feedback while ensuring scalability, enabling DSMFG to jointly model three core features of real-world policymaking: dynamic feedback, asymmetry, and large scale.
+We further introduce Stackelberg Mean Field Reinforcement Learning (SMFRL), a data-driven algorithm that learns the leader’s optimal policies while maintaining personalized responses for individual agents.
+Empirically, we validate our approach in a large-scale simulated economy, where it scales to 1,000 agents (vs.
+100 in prior work) and achieves a 4×4$\times$4 × GDP gain over classical economic methods and a 19×19$\times$19 × improvement over the static 2022 U.S.
+federal income tax policy.
+[…]
+3449
 1 Introduction
-Macroeconomic policymaking is fundamental to the sustained development of an economy (Schneider and Frey 1988; Persson and Tabellini 1999).
-Governments influence economic production, wealth distribution, social stability, and welfare through economic policies such as interest rates, taxation, and fiscal spending (Sachs and Warner 1995).
-Therefore, accurately modeling and solving optimal macroeconomic policies, and evaluating the effects of their implementation are critical issues.
-According to the Lucas critique (Lucas Jr 1976), predicting the impact of macroeconomic policies effectively requires modeling individual micro-level behaviors—termed “microfoundations”—to forecast micro-level responses to policy shifts and aggregate these into macroeconomic effects.
-Nonetheless, the vast number of micro-agents (i.e., households) and the complex dynamic interactions among them render predictions based on microfoundations exceedingly challenging.
-To address the challenges outlined, this paper introduces a Stackelberg Mean Field Game (SMFG) approach: initially proposing a Dynamic Stackelberg Mean Field Game (Dynamic SMFG) framework to model macroeconomic policymaking based on microfoundations (see Figure 1).
-Within this framework, we design a model-free reinforcement learning algorithm, Stackelberg Mean-Field Reinforcement Learning (SMFRL), aimed at optimizing macroeconomic policies by learning the dynamic responses of a large-scale population of micro-agents.
-In the Dynamic SMFG framework, the government acts as the leader, iteratively adjusting macroeconomic policies, while a vast array of micro-agents, serving as followers, react accordingly.
-This process entails sequential decision-making, with both the leader and the followers striving to optimize their cumulative long-term benefits.
-Given the leader’s policy, the SMFG is reduced to a mean field game, treating the large-scale micro-agents as a population to study their dynamic response to policy shifts.
-This dynamic SMFG framework responds to the Lucas critique.
-Existing models often simplify Dynamic SMFGs into single-step decisions (Pawlick and Zhu 2017), repeated games (Alcantara-Jimenez and Clempner 2020), or scenarios with myopic followers (Zhong et al.
-2021), which fail to capture the complexities of macroeconomic policymaking adequately.
-Therefore, we introduce the SMFRL algorithm to optimize long-term utility for both the macro-agent and numerous micro-agents under the dynamic SMFG framework.
-In the SMFRL algorithm, we utilize mean field approximation to reduce the dimensionality of the micro-agents’ high-dimensional joint states and actions and employ a leader-follower update to learn the micro-population’s best response to macroeconomic policy adjustments, as well as to optimize macroeconomic policies considering dynamic feedback from the micro-population.
-This learning method enhances the stability and convergence of policies without requiring transition information or prior knowledge of the environment.
-Our experiments, conducted on the TaxAI (Mi et al.
-2023) environment, firstly demonstrate SMFG approach’s superiority over AI Economist (Zheng et al.
-), the Saez tax (Saez 2001), the 2022 U.S.
-Federal tax, and free market policy (Backhouse 2005) from perspectives of macroeconomic policy evaluation and dynamic responses to economic shocks.
-Further, through ablation studies, we verify the significant roles of mean field approximation and leader-follower update in terms of macroeconomic policy effects, convergence, and maximal social welfare.
-When applied in real-world scenarios, the SMFG approach shows that households refusing the policy have lower wealth and utility than adopters, which enhances the SMFG policy’s attractiveness.
-In summary, this paper’s contributions are threefold:
-1. We introduce the Dynamic Stackelberg Mean Field Games framework, a novel approach for modeling macroeconomic policymaking based on microfoundations, responding to the Lucas critique.
-2.
-We propose a model-free Stackelberg Mean Field Reinforcement Learning algorithm, designed to optimize macroeconomic policies by learning the dynamic responses of large-scale micro-agents within the Dynamic SMFG framework.
-3. Experimental results demonstrate the superiority of our method in modeling and solving macroeconomic policymaking, and further explore its effectiveness in real-world settings.
-Our code is available in an anonymous GitHub repository
-https://anonymous.4open.science/r/SMFG˙macro˙7740
+Macroeconomic policy formulation is fundamental to achieving sustainable economic growth [52, 46].
+The effectiveness of these policies depends crucially on the behaviors of micro-level individuals, such as labor supply, consumption, and investment decisions [49].
+Nobel laureate Lucas has emphasized that individuals adapt their decision-making in response to changes in macroeconomic policies [37].
+Thus, systematically modeling the interactions between individuals and the government is crucial for designing effective macroeconomic policies.
+The Stackelberg game naturally captures the asymmetric interactions [19, 30] where the government (leader) sets a tax policy, and individuals (followers) adjust their labor supply and consumption in response.
+Unlike static models, this dynamic formulation accounts for the time-dependent strategic feedback essential for effective policy design [43, 5, 15].
+However, scaling this approach to large populations is computationally intractable: with $N$ agents, there are $O(N^{2})$ pairwise agent-agent interactions plus $O(N)$ leader–agent interactions.
+To address the scalability challenge inherent in macroeconomic policy modeling, we propose the Dynamic Stackelberg Mean Field Game (DSMFG) framework, which unifies dynamic feedback, asymmetry, and large scale into a coherent formulation.
+Unlike standard SMFG methods—e.g., single-step models [19] or approaches with fixed dynamics [8]—DSMFG captures the multi-period feedback loops essential to the co-evolution of policy and individual response.
+By embedding a multi-step Stackelberg game within a mean-field approximation [12, 1], DSMFG reduces the $O(N^{2})$ complexity of agent-agent interactions to $O(N)$ agent–population interactions.
+At each timestep, the government optimizes policy based on the current mean field—i.e., the population’s state–action distribution—while agents adapt to both the policy and the mean field.
+This iterative structure preserves individual feedback, reduces computational cost, and enables scalable optimization in complex macroeconomic environments.
+To learn optimal policies under DSMFG, we propose the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm.
+SMFRL introduces a Stackelberg Mean Field Q-function that enables the leader to evaluate its interactions with the aggregate population.
+Another Q-function for followers evaluates their interactions with both the leader and the population.
+Followers share a common policy that takes heterogeneous features as input, enabling personalized actions while maintaining population-level consistency.
+The central Q and shared policy are designed to ensure scalability in large populations.
+Moreover, SMFRL employs alternating updates between the leader’s and followers’ policies to ensure a stable training.
+We empirically evaluate DSMFG and SMFRL in a large-scale macroeconomic simulation environment, TaxAI, which models dynamic interactions between the government and large scale agents.
+Comparing against static policies (e.g., the 2022 U.S.
+federal tax) and dynamic rule-based methods (e.g., the Saez tax), DSMFG yields substantially better outcomes—achieving a 4×4$\times$4 × gain in per capita GDP over the Saez tax and a 19×19$\times$19 × improvement over the 2022 U.S.
+baseline.
+Unlike static or independent-agent baselines, DSMFG ensures sustainability and rapidly stabilizes income and consumption after shocks.
+Ablation studies confirm the necessity of both Stackelberg hierarchy and mean-field approximation: removing either leads to lower welfare, instability, and poor convergence.
+Despite sharing a policy, followers maintain robust performance under population heterogeneity, enabling scalable training and consistently high utility.
+These results demonstrate DSMFG’s effectiveness in solving dynamic, large-scale government–agent problems within a controlled, reproducible simulation framework, extending beyond the scope of traditional approaches.
+In summary, Our key contributions are:
+Figure 1: Macroeconomic policy-making involves both intensive individual–individual and individual–government interactions, whose pairwise complexity grows as $O(N^{2})$ with the agent number $N$—rendering direct simulation intractable.
+We address this by modeling the process as a Dynamic Stackelberg Mean Field Game (DSMFG), which approximates these complex interactions via agent–population and leader–population couplings.
+This DSMFG retains key personal behaviors while enabling scalability and capturing three defining features of real-world policymaking: dynamics, asymmetry, and large-scale.
+  • A scalable DSMFG framework that integrates three core features—dynamic feedback, asymmetry, and large scale—into a unified model for macroeconomic policymaking. (Section 3)
+  • A SMFRL algorithm that efficiently learns macro policies under DSMFG while preserving personalized decision-making at the individual level. (Section 4)
+  • Extensive empirical validation showing that SMFG scales to 1,000 agents and outperforms both economic and AI-based baselines. (Section 5)
 2 Related Work
-Economic Methods
-In the field of economics, research on economic policy typically relies on theoretical model studies, empirical analysis, and econometric methods.
-Classic theoretical models, such as the IS-LM model (Hicks 1980; Gali 1992), are applicable for analyzing short-term policy effects but overlook long-term factors and price fluctuations.
-The aggregate demand-aggregate supply (AD-AS) model (Dutt 2006; Lee, Pesaran, and Smith 1997) focuses on the relationship between aggregate demand and supply, integrating short-term and long-term factors, but simplifies macroeconomic dynamics.
-The Solow growth model (Brock and Taylor 2010) provides a framework for economic growth but does not consider market imperfections and externalities.
-New Keynesian models (Blanchard and Galí 2007; Gabaix 2020) emphasize the stickiness of prices and wages, suitable for explaining economic fluctuations and policy interventions, but are based on strong assumptions.
-DSGE models (An and Schorfheide 2007; Smets and Wouters 2007), built on micro-foundations for macroeconomic forecasting, offer a consistent framework for policy analysis but are limited in handling nonlinearities and market imperfections.
-Empirical analysis (Ramesh et al.
-2010; Vedung 2017) primarily evaluates policy effects through historical data but faces limitations in data timeliness and accuracy.
-Econometric methods (Johnston and DiNardo 1963; Davidson, MacKinnon et al.
-2004), such as regression and time series analysis, can quantify policy effects but are constrained by model settings and data quality.
-Additionally, the Saez tax model (Saez 2001) uses earnings elasticities to optimize tax rates, enhancing social welfare without deterring economic activity.
-This model continues to guide current fiscal policy through effective taxation strategies.
-AI for Economics
-Artificial intelligence (AI), such as reinforcement learning, may offer new perspectives for solving complex economic problems (Tilbury 2022).
-In terms of macroeconomics, AI Economist (Zheng et al.
-) employs curriculum learning on tax policy design, (Trott et al.
-2021) studies the collaboration between central and local governments under COVID-19, as well as monetary policy (Hinterlang and Tänzer ; Chen et al.
-2023), international trade (Sch 2021), market pricing (Danassis et al.
-2023).
-However, these studies consider the government as a single agent, neglecting the dynamic response of the households to policies.
-(Koster et al.
-) studies democratic AI by human voting models, (Yaman et al.
-2023) examines the impact of social sanction rules on labor division.
-However, these works are based on simplified settings and remain a gap in real-world economic policy.
-In terms of microeconomics, (Shi 2021; Rui and Shi 2022; Atashbar and Aruhan Shi 2023) study the optimal saving and consumption problems of micro-agents, and other researchers explore rational expectation equilibrium (Kuriksha 2021; Hill, Bardoscia, and Turrell 2021), multiple equilibria under real-business-cycle (Curry et al.
-2022), the emergence of barter behavior (Johanson et al.
-2022), optimal asset allocation and savings strategies (Ozhamaratli and Barucca 2022).
-Stackelberg Mean Field Game
-Stackelberg Mean Field Games (SMFGs) model the dynamic interactions between a leader and numerous homogeneous followers, enabling the development of macroeconomic policies grounded in microfoundations.
-The methods for SMFGs are divided into model-based and model-free algorithms.
-Model-based methods, such as (Bensoussan, Chau, and Yam 2015; Fu and Horst 2020; Dayanikli and Lauriere 2023; Bergault, Cardaliaguet, and Rainer 2023) employ forward-backward stochastic functional differential equations to ascertain the followers’ mean field equilibrium prior to computing the optimal strategy for the leader.
-Additionally, studies such as (Bensoussan et al.
-2017; Lin, Jiang, and Zhang 2018; Moon and Başar 2018; Du, Wu et al.
-2019; Huang and Yang 2020) simplify scenarios through linear-quadratic formulations.
-(Guo, Hu, and Zhang 2022), convert SMFGs into minimax optimization problems.
-However, these model-based methods generally depend on simplified representations of SMFGs and explicit environmental transition information.
-In contrast, model-free approaches leverage interaction data with the environment to learn strategies without requiring detailed knowledge of the transition dynamics and reward functions.(Pawlick and Zhu 2017) explores the resolution of SMFGs in single-step decision scenarios.(Campbell et al.
-2021) utilizes a modified deep backward stochastic differential equation framework to calculate the followers’ equilibrium despite constraints on the leader’s decisions.(Miao et al.
-2024) investigates locally optimal defensive strategies given predefined attacker trajectories.
-Furthermore,(Li et al.
-2024) learns followers’ transition functions from empirical data, thus aiding the solution of SMFGs through the empirical Fokker-Planck equation.
-Nevertheless, the assumptions and constraints inherent in these algorithms often do not meet the complex needs of macroeconomic policymaking.
-3 Stackelberg Mean Field Game Modeling
-In this section, we provide details about macroeconomic policy-making problems and propose a modeling framework based on Dynamic Stackelberg Mean Field Games (Dynamic SMFGs), with the relevant mappings are shown in Figure 1.
-Macroeconomic Policy-Making
-To foster economic growth, enhance social welfare, or maintain social equity, governments implement economic policies such as fiscal and monetary policies.
-These policies directly influence micro-agents (i.e., households) and the markets they constitute, altering their expectations and behaviors.
-The aggregation of micro-level behaviors from numerous households manifests as macroeconomic phenomena.
-Therefore, in formulating policies, governments should model micro-agents’ behaviors and predict their responses to new policies, which is so-called microfoundations in the Lucas critique.
-However, in reality, micro-agents form a large and dynamically interacting group, each seeking to maximize personal benefits.
-This complexity significantly complicates predicting their responses to macroeconomic policy shifts.
-To address these challenges, we propose using Dynamic SMFGs to model macroeconomic policy-making problems:
-1.
-For microfoundations, we model the government as the leader who first sets policies, with micro-agents as followers who adjust their behaviors according to these policies.
-The interaction between the leader and followers forms a Stackelberg game, where the leader considers the followers’ responses to make informed decisions.
-2.
-For large scale micro-agents, we employ a Mean Field Game to model micro-interactions among a large number of households, treating them as a population and predicting their collective response to macroeconomic policy changes.
-3.
-For dynamic decision-making, the formulation of macroeconomic policies is inherently a multi-step decision-making process, where both the government and households make decisions at each step to optimize long-term benefits, thereby constituting a dynamic game.
-Dynamic Stackelberg Mean Field Games
-For macroeconomic policy-making issues, we consider a dynamic SMFG with a leader and $N$ homogeneous follower agents.
-At any time step $t\in\{0,\ldots,T\}$, the leader selects an action $a_{t}^{l}\in\mathcal{A}^{l}$ based on the leader’s state $s_{t}^{l}\in\mathcal{S}^{l}$ and policy $\pi^{l}:\mathcal{S}^{l}\rightarrow\mathcal{A}^{l}$.
-Subsequently, followers determine their actions based on the leader’s action $a^{l}_{t}$ and their private states $s^{f}_{t}\in\mathcal{S}^{f}$.
-The representative follower’s action $a_{t}^{f}\in\mathcal{A}^{f}$ is derived from a shared policy $\pi^{f}:\mathcal{S}^{f}\times\mathcal{A}^{l}\rightarrow\mathcal{A}^{f}$.
+2.1 Macroeconomic Models
+Classical macroeconomic frameworks—such as IS–LM [26, 22], AD–AS [20, 34], and the Solow growth model [10]—have elucidated short- and long-run policy effects.
+New Keynesian DSGE models [9, 56] then introduced micro individuals and stochastic shocks, enabling rigorous analysis under rational expectations.
+The Saez tax model [50] further offered a practical, elasticity-based tool for setting optimal tax rates.
+However, these approaches commonly rely on linearization, representative-agent assumptions, or fixed price-stickiness parameters, which prevent them from capturing individual responses, nonlinear feedback loops, and aggregate dynamics in large populations.
+Empirical and econometric methods [47, 17] quantify policy impacts from historical data but struggle with sparsity, identification, and out-of-sample validity.
+These gaps highlight the necessity of a framework that models dynamic interactions between the government and large-scale individuals.
+2.2 RL for Economic Policy
+Recent advances have applied reinforcement learning (RL) into economic modeling.
+In macroeconomic settings, frameworks like AI Economist [60] use curriculum learning to optimize tax schedules, and others apply RL to crisis management [57], monetary policy design [28, 14], international trade dynamics [51], and market pricing with externalities [16].
+These studies typically treat the government as an independent decision-maker, overlooking how heterogeneous households adapt over time.
+At the micro level, RL has been used to study optimal savings and consumption [53, 48, 3], solve heterogeneous general-equilibrium problems [32, 27], and model agent behaviors in barter [31] and asset allocation [44].
+While these works showcase RL’s promise, they rely on simplified environments or small agent populations, limiting their applicability to dynamic, large-scale macroeconomic policy design.
+2.3 Stackelberg Mean Field Games
+Stackelberg Mean Field Games (SMFGs) integrate Stackelberg leader–follower dynamics with mean-field approximations to model interactions in large populations.
+Early model-based SMFG methods solve forward–backward stochastic differential equations to compute follower equilibria before optimizing the leader’s policy [21, 18, 8].
+Linear–quadratic formulations [7, 42, 29] and minimax rewritings [23] enhance analytical tractability but impose restrictive assumptions on dynamics and transitions, limiting applicability to complex economic settings.
+Model-free SMFG approaches dispense with explicit transition models by learning directly from interaction data.
+For example, Pawlick and Zhu [45] handle single-step SMFGs, Campbell et al.
+[11] apply deep BSDE solvers for equilibrium computation, and Miao et al.
+[40] explore defensive follower strategies under fixed attacker trajectories.
+More recently, Li et al.
+[35] estimate empirical transition kernels and solve the resulting Fokker–Planck equations.
+However, these methods remain confined to simplified benchmarks—single-period decisions, low-dimensional state spaces, or predefined follower classes—and do not capture the multi-period feedback loops and scale required for realistic macroeconomic policy design.
+Thus, there is a clear need for a stable, model-free algorithm that can solve SMFGs under complex, dynamic economic interactions without requiring knowledge of true transition dynamics.
+3 Dynamic Stackelberg Mean Field Game Framework
+In this section, we first identify the core features of the macroeconomic policy-making problem, and then propose a dynamic Stackelberg mean-field game framework to model them effectively.
+3.1 Macroeconomic Policy-Making Problem
+Macroeconomic policy comprises government actions—such as monetary and fiscal interventions—that stabilize growth, reduce unemployment, and control inflation [6].
+These interventions shape individual decisions (e.g.
+labor supply, consumption, investment), which in turn generate aggregate outcomes that inform subsequent policy adjustments [41].
+In the left panel of Figure 1, for instance, a change in the central bank’s interest-rate rule shifts households’ portfolio allocations, while each household’s choice also depends on the behavior of others.
+Such large-scale interactions induce complex feedback loops that static or small-scale models fail to capture.
+We identify three salient, interdependent features of the macroeconomic policy-making problem:
+  1. 1.
+    Dynamic feedback.
+A policy change triggers micro-level behavioral adjustments; the aggregate of these adjustments produces new macro indicators, which then feed back into the next policy decision.
+Modeling this continuous loop is essential, yet beyond the scope of static modeling methods.
+  2. 2.
+    Asymmetry.
+The government (leader) first commits to a policy rule; individual agents (followers) observe this policy and then optimize their private objectives.
+This sequential leader-follower structure underlies the inherent asymmetry dynamics between policymaker and population.
+  3. 3.
+    Large scale. Effective macro policy influences large scale micro-agents, thereby rendering the dynamic, asymmetric interactions described above more complex.
+3.2 Dynamic Stackelberg Mean Field Game
+To design effective macroeconomic policies, we must first model three core features—dynamic feedback, asymmetry, and large scale—introduced above.
+To capture dynamic feedback, we model macroeconomic policymaking as a dynamic game in which the government and individuals iteratively update their strategies based on evolving economic conditions.
+These strategies influence both immediate outcomes and the future decisions of other players [25, 58].
+To capture asymmetry, we adopt a Stackelberg leader–follower framework [55], modeling the government as the leader that sets policies first, followed by individuals’ responses.
+To capture the large-scale nature of macroeconomic systems, an agent-based model with numerous micro-agents could be considered.
+However, scaling agent-based models to large populations is computationally challenging.
+With $N$ agents, the model requires $O(N^{2})$ pairwise agent–agent interactions and $O(N)$ government–agent interactions, rendering it computationally intractable for large-scale systems.
+To address this challenge, we employ a mean-field approximation [33], where interactions between individual agents are replaced by those between a representative agent and the aggregate population, and government–agent interactions are modeled as interactions with the population’s mean field.
+This approach reduces the computational complexity from $O(N^{2})$ to $O(N)$, significantly streamlining policy optimization.
+In summary, our Dynamic Stackelberg Mean-Field Game (DSMFG) framework effectively integrates the three core features—dynamic feedback, asymmetry, and large scale—into a cohesive model for macroeconomic policymaking.
+The precise mathematical formulation of DSMFG is as follows:
+Framework Overview
+In the DSMFG framework, we consider one leader and $N$ follower agents.
+At each time step $t\in\{0,\ldots,T\}$, the leader selects an action $a_{t}^{l}\in\mathcal{A}^{l}$ based on its state $s_{t}^{l}\in\mathcal{S}^{l}$ and a policy $\pi^{l}:\mathcal{S}^{l}\to\mathcal{A}^{l}$.
+Subsequently, the followers determine their actions based on the leader’s action $a_{t}^{l}$ and their private states $s_{t}^{f}\in\mathcal{S}^{f}$.
+A representative follower’s action $a_{t}^{f}\in\mathcal{A}^{f}$ is derived from a shared policy $\pi^{f}:\mathcal{S}^{f}\times\mathcal{A}^{l}\to\mathcal{A}^{f}$.
 The sequences $\{\pi_{t}^{l}\}_{t=0}^{T}$ and $\{\pi_{t}^{f}\}_{t=0}^{T}$ are denoted as $\pi^{l}$ and $\pi^{f}$, respectively.
-We denote the population state-action distribution of followers under leader action $a^{l}_{t}$ and follower policy $\pi^{f}$ as $L_{t}(s^{f}_{t},a^{f}_{t};\pi^{f},a^{l}_{t})$, and abbreviate it as $L_{t}(s^{f}_{t},a^{f}_{t})$.
-$L_{t}(s^{f}_{t},a^{f}_{t};\pi^{f},a^{l}_{t})\in\mathcal{P}(\mathcal{S}^{f}\times\mathcal{A}^{f}),\text{where }a^{f}_{t}=\pi^{f}(s^{f}_{t},a^{l}_{t})$
-Followers’ Mean Field Game
-Given the leader’s policy $\pi^{l}$, the dynamic SMFG is simplified into a mean field game for the followers.
-At time $t\in\{0,\ldots,T-1\}$, given $(a^{l}_{t},s^{f}_{t},a^{f}_{t},L_{t})$, the representative follower then receives a reward $r^{f}(\mathbf{s}_{t},a^{l}_{t},a^{f}_{t},L_{t})$ and transitions to the next state $s^{f}_{t+1}$ according to the transition probability $s^{f}_{t+1}\sim P(\cdot|\mathbf{s}_{t},a^{l}_{t},a^{f}_{t},L_{t})$, where joint state $\mathbf{s}_{t}=\{s^{l}_{t},\mathbf{s}_{t}^{f}\}$.
-Each follower aims to find the optimal policy $\pi^{f}$ that maximizes his cumulative reward over the time horizon:
-$J^{f}(\pi^{l},\pi^{f},L)=\mathbb{E}_{s_{0}^{f}\sim\mu_{0}^{f},\mathbf{s}_{t+1}\sim P}\left[\sum_{t=0}^{T}r^{f}(\mathbf{s}_{t},a^{l}_{t},a^{f}_{t},L_{t})\right]$
-where $a^{l}_{t}=\pi^{l}(s^{l}_{t})$, $a^{f}_{t}=\pi^{f}(s^{f}_{t},a^{l}_{t})$ and $L_{t}=L_{t}(s^{f}_{t},a^{f}_{t})$.
+The mean field $L_{t}(s_{t}^{f},a_{t}^{f};\pi^{f},a_{t}^{l})$, abbreviated as $L_{t}(s_{t}^{f},a_{t}^{f})$, represents the population state-action distribution of followers, defined as:
+$L_{t}(s_{t}^{f},a_{t}^{f};\pi^{f},a_{t}^{l})\in\mathcal{P}(\mathcal{S}^{f}\times\mathcal{A}^{f}),\text{ where }a_{t}^{f}=\pi^{f}(s_{t}^{f},a_{t}^{l}).$
+Followers
+At each time step $t\in\{0,\ldots,T-1\}$, given the joint state $\mathbf{s}_{t}=\{s_{t}^{l},\mathbf{s}_{t}^{f}\}$, a representative follower receives a reward $r^{f}(\mathbf{s}_{t},a_{t}^{l},a_{t}^{f},L_{t})$ and transitions to the next state $s_{t+1}^{f}\sim P(\cdot\mid\mathbf{s}_{t},a_{t}^{l},a_{t}^{f},L_{t})$.
+The follower’s objective is to optimize their policy $\pi^{f}$ to maximize cumulative rewards over the time horizon:
+$J^{f}(\pi^{l},\pi^{f},L)=\mathbb{E}_{s_{0}^{f}\sim\mu_{0}^{f},\mathbf{s}_{t+1}\sim P}\left[\sum_{t=0}^{T}r^{f}(\mathbf{s}_{t},a_{t}^{l},a_{t}^{f},L_{t})\right],$
+where the leader’s action $a_{t}^{l}=\pi^{l}(s_{t}^{l})$, the follower’s action $a_{t}^{f}=\pi^{f}(s_{t}^{f},a_{t}^{l})$, and the mean field $L_{t}=L_{t}(s_{t}^{f},a_{t}^{f})$.
 Definition 1 (Followers’ Best Response for Leader’s Policy).
-Given any leader’s policy $\pi^{l}\in\Pi^{l}$ and the followers’ state-action distributions $L=\{L_{t}\}_{t=0}^{T}$, the followers’ best response policy ${\pi^{f}}^{*}(\pi^{l},L)$ is defined as
+Given a leader’s policy $\pi^{l}\in\Pi^{l}$ and followers’ state-action distributions $L=\{L_{t}\}_{t=0}^{T}$, the followers’ best response policy ${\pi^{f}}^{*}(\pi^{l},L)$ is defined as:
 ${\pi^{f}}^{*}(\pi^{l},L)\in\arg\max_{\pi^{\prime}}J^{f}(\pi^{l},\pi^{\prime},L).$
-Leader’s Game
-At time $t\in\{0,...,T-1\}$, given the leader’s state $s^{l}_{t}$, action $a^{l}_{t}$, the followers’ policy $\pi^{f}$ and state-action distribution $L_{t}$, the leader will receive a reward $r^{l}(\mathbf{s}_{t},a_{t}^{l},L_{t})$, and move to the next state $s_{t+1}^{l}$ according to the transition probabilities $P(\cdot|\mathbf{s}_{t},a_{t}^{l},L_{t})$.
-The leader aims to find the optimal policy $\pi^{l}$ to maximize the total expected reward:
+Leader
+At each time step $t\in\{0,\ldots,T-1\}$, the leader receives a reward $r^{l}(\mathbf{s}_{t},a_{t}^{l},L_{t})$ based on its state $s_{t}^{l}$, action $a_{t}^{l}$, and the population mean field $L_{t}$, and transitions to the next state $s_{t+1}^{l}\sim P(\cdot\mid\mathbf{s}_{t},a_{t}^{l},L_{t})$.
+The leader aims to optimize its policy $\pi^{l}$ to maximize the expected cumulative reward:
 $J^{l}(\pi^{l},\pi^{f},L)=\mathbb{E}_{s_{0}^{l}\sim\mu_{0}^{l},\mathbf{s}_{t+1}\sim P}\left[\sum_{t=1}^{T}r^{l}(\mathbf{s}_{t},a_{t}^{l},L_{t})\right],$ (1)
@@
 $\mu_{t+1}^{l}(s_{t+1}^{l})=\sum_{s_{t}^{l},a_{t}^{l}}\mu_{t}^{l}(s_{t}^{l})\pi_{t}^{l}(a_{t}^{l}\mid s_{t}^{l})P(s_{t+1}^{l}\mid s_{t}^{l},a_{t}^{l},L_{t}).$
-In conclusion, to study the issue of macroeconomic policy-making, we model it as a dynamic SMFG.
-Under this modeling framework, we can optimze the macroeconomic policy of the leader, namely the government, while considering the followers’ best response over discrete timesteps (Definition 2).
+In conclusion, we model the problem of macroeconomic policy-making as a DSMFG, capturing three core features.
+Based on this model, we optimize the government’s policy by taking into account the followers’ best responses over discrete timesteps (Definition 2).
 4 Stackelberg Mean Field Reinforcement Learning
-To solve macroeconomic policymaking under Dynamic SMFG framework, we propose the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm, which is a centralized training with decentralized execution (CTDE) algorithm.
-We introduce the leader’s policy $\pi^{l}(s^{l})$ and a central critic Q-function $Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ for the leader agent, and the shared policy $\pi^{f}(s^{f},a^{l})$ and $Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ for each follower agent.
-SMFRL algorithm comprises two key components: mean field approximation and leader-follower update (see Figure 2).
+Within the DSMFG framework, we propose the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm (Figure 2), which adopts a centralized training with decentralized execution (CTDE) paradigm [36].
+In standard CTDE settings, the leader agent is equipped with a policy $\pi^{l}(s^{l})$ and a centralized critic $Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$, while each follower is equipped with a policy $\pi^{f}(s^{f},a^{l})$ and a corresponding Q-function $Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$.
+However, the joint state $\mathbf{s}^{f}$ and action $\mathbf{a}^{f}$ of all followers scale linearly with the population size, rendering direct learning of $Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ computationally infeasible in large-scale environments.
+This directly reflects the computational complexity challenge induced by large-scale agent interactions, as discussed in Section 3.
+4.1 Stackelberg Mean-Field Q and Policy
+Stackelberg Mean-Field Q
+Within the DSMFG framework, we abstract the interactions between the leader and all followers into an interaction between the leader and the population mean field, thereby the leader’s centralized Q-function can be reformulated as:
+$Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})\approx\tilde{Q}^{l}(s^{l},a^{l},L).$
+Similarly, based on the mean field approximation, we simplify the interaction of a follower with the leader and other followers into an interaction with the leader and the population mean field.
+Accordingly, the original Q-function $Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ can be approximated as:
+$Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ $=Q^{f}(s^{l},a^{l},s^{fi},a^{fi},\mathbf{s}^{f,-i},\mathbf{a}^{f,-i})$
+$\approx\tilde{Q}^{f}(s^{l},a^{l},s^{fi},a^{fi},L),$
+where $(s^{fi},a^{fi})$ denotes the individual state-action pair of the $i$-th follower, and $(\mathbf{s}^{f,-i},\mathbf{a}^{f,-i})$ represent the rest of the population.
+Given that each individual has a negligible impact on the collective, the population mean field $L$ can be used to approximate the rest of the population, following the theory of mean field games [33].
+In experiments, the mean field $L$ is constructed from the empirical distribution over the followers’ state-action pairs. At timestep $t$,
+$L_{t}=P(s^{f},a^{f}),\quad(s^{f},a^{f})\sim\{(s^{f_{i}}_{t},a^{f_{i}}_{t})\}_{i=1}^{N},a^{f_{i}}_{t}\sim\pi^{f}(\cdot\mid s^{f_{i}}_{t},a^{l}_{t}),$
+where $N$ denotes the number of followers.
+In prior works on mean-field methods, a widely used simplification is to approximate $L_{t}$ using population averages:
+$L_{t}\approx(\bar{s}^{f},\bar{a}^{f}),\quad\bar{s}^{f}=\frac{1}{N}\sum_{i=1}^{N}s^{f_{i}},\quad\bar{a}^{f}=\frac{1}{N}\sum_{i=1}^{N}a^{f_{i}}.$ (2)
+Alternative representations include neighborhood-based action averages [59], empirical distributions [13], and graph-based weighted mean fields [24].
+In our experiments, we find that using the average-based mean field achieves strong empirical performance while maintaining tractable computational cost.
+In reinforcement learning, the local Q-functions $\tilde{Q}^{l}$ and $\tilde{Q}^{f}$ are defined as the expected cumulative rewards under discount factor $\gamma\in[0,1]$, starting from given states and actions:
+$\tilde{Q}^{l}(s^{l},a^{l},s^{f},a^{f})=\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}^{l}\right],$ (3)
+$\tilde{Q}^{f}(s^{l},a^{l},s^{fi},a^{fi},s^{f},a^{f})=\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}r^{fi}_{t}\right].$
+Here, $r_{t}^{l}$ and $r_{t}^{fi}$ are the rewards for the leader and $i$-th follower, respectively.
+The discount factor $\gamma$ determines the planning horizon: $\gamma=0$ models myopic agents focusing on immediate rewards, while $\gamma>0$ models non-myopic agents with long-term effects, increasing computational complexity.
+Followers’ policy
+Under the standard CTDE paradigm, the leader’s policy takes the individual state as input, while the follower’s policy conditions on both the individual’s state and the leader’s action.
+These input dimensions are tractable.
+However, training a separate policy for each follower is computationally infeasible.
+To address this, we adopt a shared policy $\pi^{f}(s^{f},a^{l})$ for all followers, as required by the mean-field setting.
+While this introduces inherent homogeneity—a known limitation of mean-field methods—we preserve individual heterogeneity by encoding personalized information in the state input.
+In reinforcement learning, a policy maps states to actions.
+Through training, the model learns how state features influence decisions.
+For instance, individual states include attributes such as age, education, and wealth, while actions cover economic choices like investment.
+This enables the shared policy to generate personalized behaviors, e.g., individuals of different ages and wealth levels exhibit distinct investment patterns.
 Figure 2: The architecture of SMFRL algorithm.
-Mean Field Approximation
-Due to the dimensions of joint state $\mathbf{s}^{f}$ and action $\mathbf{a}^{f}$ scaling with the number of follower agents, the standard Q-function $Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ becomes infeasible to learn when dealing with a large number of followers.
-Therefore, we factorize the standard Q-function using pairwise local interactions between the leader and each follower (Yang et al.
-2018):
-$Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ $=\frac{1}{N}\sum_{i}^{N}\tilde{Q}^{l}(s^{l},a^{l},s^{fi},a^{fi})$ (2)
-$\approx\mathbb{E}_{(s^{f},a^{f})\sim L}\tilde{Q}^{l}(s^{l},a^{l},s^{f},a^{f})$
-Similarly, the $i$-th follower’s Q-function $Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ can be factorized using local interactions between the leader, the $i$-th follower, and each of the remaining followers:
-$Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})\approx\mathbb{E}_{(s^{f},a^{f})\sim L}\tilde{Q}^{f}(s^{l},a^{l},s^{fi},a^{fi},s^{f},a^{f})$
-where $(s^{fi},a^{fi})$ and $(s^{f},a^{f})$ are the state-action pair for the $i$-th follower and another follower agent, respectively.
-We use $\hat{L}(\pi^{f},a^{l})$, a finite set of state-action pairs sampled under the follower’s policy $\pi^{f}$ and the leader’s action $a^{l}$, to construct an empirical distribution that approximates the state-action distribution $L(s^{f},a^{f})$.
-$\hat{L}(\pi^{f},a^{l})=\left\{(s^{fi},a^{fi})\right\}_{i=1}^{N},\text{ where }a^{fi}=\pi^{f}(s^{fi},a^{l}).$
-Leader-follower Update
-To enhance the convergence and stability, we propose the leader-follower update: first, by fixing the leader’s policy, we train the followers’ shared policy and Q-networks towards the best response; subsequently, based on the followers’ policies, we optimize the leader’s policy and Q net, alternating these steps until convergence.
+4.2 Leader-follower Update
+To enhance the convergence and stability, we propose the leader-follower update (shown in Figure 2): first, by fixing the leader’s policy, we train the followers’ shared policy and Q-networks towards the best response; subsequently, based on the followers’ policies, we optimize the leader’s policy and Q net, alternating these steps until convergence.
 We measure the distance between the agents’ policies and their best responses by using exploitability.
@@
 At any step $t\in\{0,\ldots,T\}$, the tuple $(s^{l}_{t},a^{l}_{t},\mathbf{s}^{f}_{t},\mathbf{a}^{f}_{t},s^{l}_{t+1},\mathbf{s}^{f}_{t+1},r^{l}_{t},\mathbf{r}^{f}_{t})$ is stored in the replay buffer $\mathcal{D}$ for training.
-When the number of followers is particularly large, a finite subset of state-action pairs $\{(s^{fi}_{t},a^{fi}_{t})\}_{i=1}^{N}$ can be selected for storage.
 Followers’ Update for Best Response
-Given leader’s policy $\pi_{\theta^{l}}$, the followers’ policy network $\pi_{\theta^{f}}$ is updated using a deterministic policy gradient approach (Silver et al.
-2014).
-The followers’ policy gradient is estimated:
-$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t}\sim\mathcal{D},(s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}}$
-$\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{fi}_{t},a^{l}_{t})\nabla_{a^{fi}_{-}}\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{fi}_{t},a^{fi}_{-},s^{f}_{t},a^{f}_{t})\Big{]}$
-where $a^{fi}_{-}=\pi_{\theta^{f}}(s^{fi}_{t},a^{l}_{t})$, $\hat{L}_{\mathcal{D}}$ denotes state-action pairs sampled from replay buffer $\mathcal{D}$.
+Given leader’s policy $\pi_{\theta^{l}}$, the followers’ policy network $\pi_{\theta^{f}}$ is updated using a deterministic policy gradient [54]. The followers’ policy gradient is estimated:
+$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L_{t}\sim\mathcal{D}}\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{-},L_{t})\Big{]}$
+where $a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})$, $L_{t}$ is computed by state-action pairs sampled from replay buffer $\mathcal{D}$ by Eq.
+(2).
 The action-value function $\tilde{Q}_{\phi^{f}}$ is updated by minimizing the mean squared error loss:
-$\mathcal{L}(\phi^{f})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\\ (s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}\end{subarray}}\left[\left(y^{fi}_{t}-\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{fi}_{t},a^{fi}_{t},s^{f}_{t},a^{f}_{t})\right)^{2}\right]$
-$y^{fi}_{t}=r^{fi}_{t}+$
-$\gamma\mathbb{E}_{(s^{f}_{t+1},a^{f}_{t+1})\sim\hat{L}(\pi_{\theta^{f}_{-}},a^{l}_{t+1})}\tilde{Q}_{\phi^{f}_{-}}(s^{l}_{t+1},a^{l}_{t+1},s^{fi}_{t+1},a^{fi}_{t+1},s^{f}_{t+1},a^{f}_{t+1})$
-where $a^{l}_{t+1}=\pi_{\theta^{l}}(s^{l}_{t+1})$, $a^{fi}_{t+1}=\pi_{\theta^{f}_{-}}(s^{fi}_{t+1},a^{l}_{t+1})$. The gradient of the loss function $\mathcal{L}(\phi^{f})$ is derived as:
-$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\\ (s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}\end{subarray}}$ $\Big{[}\left(y^{fi}_{t}-\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{fi}_{t},a^{fi}_{t},s^{f}_{t},a^{f}_{t})\right)$
-$\nabla_{\phi^{f}}\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{fi}_{t},a^{fi}_{t},s^{f}_{t},a^{f}_{t})\Big{]}.$
+$\mathcal{L}(\phi^{f})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},L_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\end{subarray}}\left[\left(y^{f}_{t}-\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{-},L_{t})\right)^{2}\right]$
+$y^{f}_{t}=r^{f}_{t}+\gamma\tilde{Q}_{\phi^{f}_{-}}(s^{l}_{t+1},a^{l}_{t+1},s^{f}_{t+1},a^{f}_{t+1},L_{t+1})$
+where $a^{l}_{t+1}=\pi_{\theta^{l}}(s^{l}_{t+1})$, $a^{f}_{t+1}=\pi_{\theta^{f}_{-}}(s^{f}_{t+1},a^{l}_{t+1})$. The gradient of the loss function $\mathcal{L}(\phi^{f})$ is derived as:
+$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},L_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\end{subarray}}$ $\Big{[}\left(y^{f}_{t}-\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{t},L_{t})\right)$
+$\nabla_{\phi^{f}}\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{t},L_{t})\Big{]}.$
 Leader’s Update for Optimal Policy
 Given followers’ policy $\pi_{\theta^{f}}$, the leader’s policy $\pi_{\theta^{l}}$ is optimized by DPG approach, and the leader’s policy gradient is estimated:
-$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D},(s^{f},a^{f})\sim\hat{L}(\pi_{\theta^{f}},a^{l}_{-})}$
-$\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}\tilde{Q}_{\phi^{l}}(s^{l},a^{l}_{-},s^{f},a^{f})\right],\text{where }{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}.$
+$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t},L_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}\tilde{Q}_{\phi^{l}}(s_{t}^{l},a^{l}_{-},L_{t})\right]|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}.$
 This network is periodically updated to minimize the loss:
-$\mathcal{L}(\phi^{l})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\\ (s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}\end{subarray}}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{t})\right)^{2}\right]$
+$\mathcal{L}(\phi^{l})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},L_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\end{subarray}}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L_{t})\right)^{2}\right]$
 The target value $y^{l}_{t}$ is given by:
-$y^{l}_{t}=r_{t}^{l}+\gamma\mathbb{E}_{(s^{f}_{t+1},a^{f}_{t+1})\sim\hat{L}(\pi_{\theta^{f}},a^{l}_{t+1})}\tilde{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},s^{f}_{t+1},a^{f}_{t+1})$
-where $a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})$, and $\gamma$ is the discount factor. Differentiating the loss function $\mathcal{L}(\phi^{l})$ yields the gradient utilized for training:
-$\nabla_{\phi^{l}}$ $\mathcal{L}(\phi^{l})=\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},\mathbf{s}_{t+1}\sim\mathcal{D},(s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}}$
-$\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{t})\right)\nabla_{\phi^{l}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{t})\right].$
-The pseudocode for the SMFRL algorithm 1 is presented in Appendix A.
+$y^{l}_{t}=r_{t}^{l}+\gamma\tilde{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})}$
+where $\gamma$ is the discount factor. Differentiating the loss function $\mathcal{L}(\phi^{l})$ yields the gradient utilized for training:
+$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=\mathbb{E}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L_{t})\right)\nabla_{\phi^{l}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L_{t})\right].$
+where the expectation $\mathbb{E}$ is taken over $(\mathbf{s}_{t},a^{l}_{t},L_{t},\mathbf{s}_{t+1})\sim\mathcal{D}$. The pseudocode for the SMFRL algorithm 1 in Appendix A.
 5 Experiment
-In this section, we conduct experiments of the SMFG method, which is modeled by dynamic SMFGs and solved by the SMFRL algorithm, within the TaxAI (Mi et al.
-2023) environment to answer the following questions:
-1. For macroeconomic policy-making issues, how does our SMFG method compare to real-world policies, classic economic methods, and existing AI-based policy?
-2. Are both the leader-follower update and mean-field approximation modules of the SMFG method essential?
-3. Given the likely real-world scenario where some households do not adopt the SMFG policy, does the SMFG approach remain effective?
-TaxAI Environment
-TaxAI is a simulation environment based on real data designed to study optimal taxation policies.
-It simulates the dynamic interactions between the government and large-scale households, to promote economic growth (i.e., GDP) as the objective.
-Tax Policies Per Capita GDP Social Welfare Income Gini Wealth Gini Years
-$100$ $1000$ $100$ $1000$ $100$ $1000$ $100$ $1000$ $100$ $1000$
-Free Market 1.37e+05 1.41e+05 32.97 334.79 0.89 0.90 0.92 0.93 1.10 1.00
-Saez Tax 2.34e+12 6.35e+11 73.82 498.88 0.21 0.68 0.38 0.73 300.00 100.58
-US Federal Tax 4.88e+11 1.41e+05 94.19 351.17 0.40 0.89 0.40 0.93 289.55 1.00
-AI Economist-BC 4.24e+12 N/A 97.24 N/A 0.54 N/A 0.52 N/A 299.55 N/A
-AI Economist 1.26e+05 N/A 72.81 N/A 0.88 N/A 0.91 N/A 1.00 N/A
-SMFG 9.59e+12 1.10e+13 96.87 968.94 0.52 0.54 0.51 0.53 300.00 300.00
-Table 1: Assessment results of 6 macroeconomic policies for different numbers of households N=100 and N=1000.
-Figure 3: Dynamic response curves of 3 macroeconomic policies to economic shocks at step 100.
-The SMFG policy (dark blue line) exhibits the least fluctuation and the fastest recovery in key indicators, indicating superior dynamic response capabilities.
-Figure 4: Ablation studies of SMFG method based on macroeconomic indicators for households’ numbers N=100 and N=1000.
-Comparison of Macroeconomic Policies
-This section compares the tax policy derived from SMFG with existing taxation policies in two aspects: (1) Macroeconomic Policy Evaluation: We assess the performance of different policies using key macroeconomic indicators.
-(2) Dynamic Response: At step 100, we simulate an economic shock caused by a financial crisis to evaluate the dynamic response capabilities of the different tax policies.
-Baselines and Metrics
-We select four different types of tax policy baselines, including the Free Market (Backhouse 2005) policy, which entails no government intervention; the Saez Tax (Saez 2001), a well-known economic approach often advocated for specific tax reforms in practice; the U.S.
-Federal Tax, which is the actual policy derived from OECD data https://data.oecd.org/; and the AI Economist (Zheng et al.
-), a popular AI-based policy utilizing a two-level Proximal Policy Optimization (PPO) framework.
-Additionally, the AI Economist-BC method employs behavior cloning on real-world data for pretraining.
-In the TaxAI environment, we evaluate the performance of these policies using key macroeconomic indicators derived from real-world data.
+Table 1: Performance of multiple policies on key macroeconomic indicators for $N=100$ and $N=1000$ households. The best values are highlighted in bold, and the second-best values are underlined.
+Category Subcategory Policies Per Capita GDP $\uparrow$ Social Welfare $\uparrow$ Wealth Gini $\downarrow$ Years $\uparrow$
+$100$ $1000$ $100$ $1000$ $100$ $1000$ $100$ $1000$
+Static Non-intervention Free Market 1.37e+05 1.41e+05 32.97 334.79 0.92 0.93 1.10 1.00
+Policies Real-data US Federal Tax 4.88e+11 1.41e+05 94.19 351.17 0.40 0.93 289.55 1.00
+Dynamic Rule-based Saez Tax 2.34e+12 6.35e+11 73.82 498.88 0.38 0.73 300.00 100.58
+Policies Independent-based AI Economist 1.26e+05 N/A 72.81 N/A 0.91 N/A 1.00 N/A
+Game-based DSMFG (ours) 9.59e+12 1.10e+13 96.87 968.94 0.51 0.53 300.00 300.00
+DSMFG w/o S 8.66e+07 1.31e+05 82.02 834.93 0.83 0.92 75.75 1.50
+DSMFG w/o MF 1.23e+05 1.48e+05 48.17 499.01 0.93 0.93 1.00 1.02
+DSMFG w/o MF & S 1.21e+05 1.33e+05 83.09 874.53 0.92 0.92 1.00 1.00
+To validate the effectiveness of our DSMFG framework and SMFRL algorithm for macroeconomic policymaking, our experiments are designed to answer the following key questions:
+  1. 1.
+    Effectiveness of dynamic modeling. Does the DSMFG framework outperform static macroeconomic policies in optimizing critical economic indicators? (§ 5.2)
+  2. 2.
+    Necessity of Stackelberg structure and mean field approximation. Are the Stackelberg structure and mean field approximation essential for the scalability and performance of DSMFG? (§ 5.3)
+  3. 3.
+    Impact of mean field homogeneity. How does the homogeneity assumption inherent in mean field approximations affect decision personalization and policy robustness? (§ 5.4)
+In the appendix, we include details on computational resources and efficiency (E.1), full training curves and result tables (E.2,E.3), and discussions on the efficiency-equity of different policies (E.4).
+5.1 Experimental Setting
+Environment
+We conduct experiments in TaxAI [39], a simulation platform for optimal tax policy.
+TaxAI enables dynamic interactions between governments and large-scale households using real-world datasets.
+Details are in Appendix F.1.
+Evaluation Metrics
 Per Capita GDP reflects the level of economic development, while Income Gini and Wealth Gini measure inequality in household income and wealth, respectively—a lower Gini index indicates greater social equality.
@@
 Average Wealth, Income, and Consumption are crucial assessment metrics related to financial crises.
-Macroeconomic Policy Evaluation
-In the context of optimizing economic growth, our SMFG method achieves the highest per capita GDP, social welfare, and the longest duration of outcomes across both $N=100$ and $N=1000$ scenarios.
-Table 1 shows the average values of macroeconomic indicators across 10 test runs for different algorithms under varying household numbers.
-Although the Gini index for SMFG is higher than that of the Saez Tax when $N=100$, this is acceptable when optimizing GDP, as SMFG’s per capita GDP is nearly 4 times that of the Saez Tax.
-When $N=1000$, SMFG outperforms all other metrics, indicating that the advantages of our SMFG method increase with the number of households.
-Figure 7 in the Appendix provides the corresponding training curves.
-Dynamic Response
-We simulate an economic shock similar to a financial crisis in the TaxAI environment: at step 100, all households lose 50% of their wealth.
-Due to the short duration (Years) under the Free Market and AI economist strategies, we focus on comparing the Saez Tax, US Federal Tax, and SMFG policy.
-From Figure 3, we find that the SMFG policy (dark blue line) demonstrates the fastest dynamic response across all indicators: Social Welfare remains largely unaffected, while Average Income and Average Consumption recover to their original levels within about 5 steps.
-Wealth Gini and Average Wealth return to their previous levels within about 15 steps.
-This level of recovery is unmatched by any other tax policy.
+Baselines
+We compare our method against static, dynamic, and game-based policies to validate the necessity of the proposed DSMFG framework (see Table 1).
+The parameters of the baselines are provided in Appendix G.
+  • Static Policies:
+    • Free Market [4]: No government intervention.
+    • U.S. Federal Tax: The actual progressive personal income tax policy implemented by the U.S. federal government in 2022. This serves as a strong static policy baseline.
+  • Dynamic Policies:
+    • Saez Tax [50]: A rule-based economic method widely recommended for tax reforms in real world (details in Appendix C).
+    • AI Economist [60]: An independent-based policy employing independent Proximal Policy Optimization (PPO), which does not consider multi-agent interactions.
+  • Game-based Policies:
+    • DSMFG (Ours): Incorporates dynamic Stackelberg Mean Field Games.
+    • DSMFG w/o S: DSMFG without Stackelberg structure.
+    • DSMFG w/o MF: DSMFG without mean field approximation.
+    • DSMFG w/o MF & S: DSMFG excluding both Stackelberg structure and mean field approximation.
+5.2 Effectiveness of Dynamic Modeling
+Figure 3: Dynamic response curves of 3 macroeconomic policies to economic shocks at step 100.
+The DSMFG policy (dark blue line) exhibits the least fluctuation and the fastest recovery in key indicators, indicating superior dynamic response capabilities.
+Our DSMFG framework, a dynamic game-based policy, outperforms static, rule-based, and independent policies across key economic metrics (Table 1) and responds more rapidly to economic shocks (Figure 3).
+Superior Performance of Dynamic Policies
+Table 1 ranks policy performance: DSMFG (dynamic game-based) $\succ$ Saez Tax (dynamic rule-based) $\succ$ U.S.
+Federal Tax (static) $\succ$ others.
+In the TaxAI environment, which models complex economic interactions, ineffective policies often trigger termination conditions (e.g., extreme inequality, Gini $>0.9$).
+Free Market and AI Economist policies, lacking government leadership or multi-agent modeling, fail to achieve sustainable outcomes, as evidenced by the Years metric.
+Among sustainable policies (U.S.
+Federal Tax, Saez Tax, DSMFG), dynamic policies outperform static ones: DSMFG achieves a Per Capita GDP 19 times higher than U.S.
+Federal Tax ($9.59\times 10^{12}$ vs.
+$4.88\times 10^{11}$) and 4 times higher than Saez Tax ($9.59\times 10^{12}$ vs.
+$2.34\times 10^{12}$).
+At $N=1000$, U.S.
+Federal Tax, based on 2022 static data, fails to sustain economic development, while Saez Tax shows declining Per Capita GDP, Social Welfare, and Gini metrics.
+In contrast, DSMFG maintains higher Per Capita GDP with comparable welfare and Gini, demonstrating superior scalability and performance over static, rule-based, and AI-based policies.
+Rapid Response to Economic Shocks
+To assess dynamic responsiveness, we simulate a financial crisis in the TaxAI environment, where all households lose 50% of their wealth at step 100 (see Appendix F.3 for details).
+As Free Market and AI Economist policies are limited to one-year simulations (Table 1), we compare DSMFG against U.S.
+Federal Tax and Saez Tax.
+Figure 3 shows DSMFG (dark blue line) recovering fastest across all metrics: Social Welfare remains stable, Average Income and Consumption recover within 5 steps, and Wealth Gini and Average Wealth stabilize within 15 steps.
+This unmatched resilience highlights DSMFG’s precise modeling of follower decisions and superior adaptability.
 Figure 4: The left subfigure illustrates a scenario with heterogeneous households adopting various policies.
-The right subfigure presents test results for households’ decisions (consumption, labor), microeconomic indicators (utility, wealth, income), and macroeconomic indicators (GDP) when there are different proportions of SMFG households ($N=100$).
-Ablation Studies of SMFG
-We conduct ablation studies to analyze the impact of the leader-follower update and mean field approximation modules in the SMFG method on macroeconomic indicators, as well as on convergence and social welfare.
-Baselines
-We conduct comparisons based on the SMFG method by excluding the leader-follower update (denoted as S) and the mean field approximation (denoted as MF) individually:
-1.
-SMFG-S: SMFG without leader-follower update.
-The government agent uses the independent DDPG (de Witt et al.
-2020), while interactions among large-scale households are modeled as a mean field game and trained using the mean-field MARL algorithm (Yang et al.
-2018).
-2.
-SMFG-MF: SMFG without mean field approximation.
-Both the government and household agents are trained using a variant of MADDPG algorithm (Lowe et al.
-2017), with the government making decisions first, followed by households.
-3. SMFG-S-MF: SMFG without both leader-follower update and mean field approximation. In this case, both the government and household agents use independent DDPG.
-Macroeconomic Indicators Analysis
-Figure 4 presents the results of the ablation studies on various macroeconomic indicators, evaluated across five seeds.
-The height of each bar represents the mean of the macroeconomic indicators, while the error bars indicate the variance.
-Our findings reveal that: (1) Excluding either module leads to a significant performance drop compared to the full SMFG method, with this effect becoming more pronounced as the number of agents $N$ increases.
-This underscores the critical importance of both the S and MF modules.
-(2) The SMFG-S variant outperforms SMFG-MF and SMFG-S-MF in terms of higher Per Capita GDP, social welfare, and Years, as well as lower Gini index.
-This suggests the dimensionality reduction of the MF module is crucial for the SMFG method’s overall effectiveness.
-(3) The low variance of the SMFG method indicates its stability.
-Convergence and Social Welfare
-Table 2 presents the performance of different baselines across three game theory-related metrics: leader’s payoff, social welfare, and exploitability.
-Leader’s payoff reflects the performance of the leader agent (i.e., the government), social welfare corresponds to the total utility of all households, and exploitability measures the distance between current policy and Nash equilibrium.
-The methods for calculating the metrics and optimal values are provided in Appendix D.
-From Table 2, we find: (1) SMFG achieves the highest leader’s payoff, the highest social welfare, and near-zero exploitability, indicating that the SMFG policy is closest to both the equilibrium and the maximal social welfare.
-(2) The high exploitability exhibited by SMFG-S indicates that the leader-follower update module facilitates the algorithm’s convergence toward equilibrium.
-(3) SMFG-MF outperforms SMFG-S-MF in terms of the leader’s payoff, suggesting that the leader-follower update module aids in optimizing the leader’s objective.
-However, it is challenging for SMFG-MF to optimize the followers’ social welfare.
+The right subfigure presents test results for households’ decisions (consumption, labor), microeconomic indicators (utility, wealth, income), and macroeconomic indicators (GDP) when there are different proportions of DSMFG followers ($N=100$).
+Figure 5: The training curves for 6 algorithms on 4 macroeconomic indicators, comparing settings without behavior cloning as pre-train (N=100 & N=1000) and with behavior cloning (N=100-BC & N=1000-BC).
+5.3 Necessity of Stackelberg and Mean-Field Components
+To evaluate the importance of the Stackelberg and mean-field components in DSMFG, we compare DSMFG against its ablated variants (DSMFG w/o S, DSMFG w/o MF, DSMFG w/o MF & S) using macroeconomic metrics (Table 1), training dynamics (Figure 5), and game-theoretic indicators (Table 2).
+These indicators-leader payoff, exploitability, and social welfare-quantify the leader’s policy optimization, convergence to equilibrium, and follower policy quality, respectively (see Appendix D for details).
+Economic performance and training stability
+Table 1 demonstrates DSMFG’s superior performance across all economic metrics.
+Removing either the Stackelberg or mean-field component causes substantial performance degradation.
+For instance, DSMFG w/o S yields a social welfare of $834.93$ for $N=1000$, a $14\%$ drop from DSMFG, while DSMFG w/o MF reduces social welfare to $499.01$, nearly halving DSMFG’s value.
+Training curves in Figure 5 further confirm these findings: DSMFG converges stably, while its variants exhibit slower convergence and higher variance (shaded regions, based on five seeds), particularly for DSMFG w/o S and DSMFG w/o MF.
+These results highlight the indispensable role of both components in achieving robust and optimal policy outcomes.
+Table 2: Ablation studies of the DSMFG method based on game theory metrics for N=100 and N=1000. Optimal values are provided for reference.
 Leader’s Exploitability Social
 Payoff Welfare
-[…]
+Methods 100 1000 100 1000 100 1000
 Optimal Value \ \ 0. 0. 100 1000
-SMFG-S-MF -774 -716 0.652 1.023 83 859
-SMFG-S -448 -856 3.161 1.213 82 782
-SMFG-MF -535 -441 0.782 0.725 54 499
-SMFG (ours) 3294 3376 0.002 0.023 98 971
-Table 2: Ablation studies of the SMFG method based on game theory metrics for N=100 and N=1000. Optimal values are provided for reference.
-Effectiveness Analysis of SMFG
-Problem Description
-The above experiments have shown the effectiveness of the SMFG method in macroeconomic policymaking.
-Assuming the government is willing to implement macroeconomic policies learned by the SMFG method, we aim to explore the impact of non-adoption by some households.
-In our experiments, we introduce Non-SMFG Households whose policies are trained using a behavior cloning algorithm based on real data, reflecting the decision-making styles of real-world households.
-Details of these behavior cloning experiments are detailed in Appendix E.
-Additionally, we denote governments and households that adopt the SMFG policy as SMFG Government and SMFG Households, respectively.
-We set different proportions of SMFG households (0%, 25%, 50%, 75%, and 100%) and evaluate our SMFG method using microeconomic and macroeconomic indicators.
-The results is shown in Figure 5, each subplot displays average values for two types of households (bars on the left Y-axis) and their aggregate values (dots on the right Y-axis).
-GDP, a macroeconomic indicator, is shown only in aggregate, not for individual households.
-Results
-From Figure 5, we observe several intriguing findings: (1) SMFG Households outperform Non-SMFG Households across microeconomic indicators such as wealth, income, and utility, suggesting that the policy of SMFG households is more attractive and likely to encourage broader adoption; (2) The presence of Non-SMFG Households does not prevent SMFG Households from achieving near-optimal utility, which is capped at 1 in the TaxAI environment; (3) As the proportion of SMFG Households increases, the per capita GDP also rises, indicating that the government should encourage households to adopt the SMFG policy.
+DSMFG (ours) 3294 3376 0.002 0.023 98 971
+DSMFG w/o S -448 -856 3.161 1.213 82 782
+DSMFG w/o MF -535 -441 0.782 0.725 54 499
+DSMFG w/o MF & S -774 -716 0.652 1.023 83 859
+Game-Theoretic Analysis
+Table 2 provides deeper insights into the contributions of each component.
+DSMFG achieves a leader payoff of $3294$ for $N=100$ and an exploitability of $0.002$, closely approaching the optimal value of 0.
+(i) Removing the Stackelberg module (DSMFG w/o S) drastically reduces the leader payoff to $-448$ and increases exploitability to $3.161$, indicating the Stackelberg structure’s critical role in optimizing leader policies and ensuring equilibrium convergence.
+Similarly, (ii) removing the mean-field component (DSMFG w/o MF) lowers social welfare to $54$ for $N=100$, a $45\%$ reduction from DSMFG’s $98$, underscoring its necessity for effective follower policy optimization.
+The combined ablation (DSMFG w/o MF & S) reduces the framework to standard multi-agent reinforcement learning, performing comparably to DSMFG w/o MF but with increased exploitability ($1.023$ vs.
+$0.725$ for $N=1000$).
+This further emphasizes the Stackelberg module’s importance in large-scale settings, where it facilitates convergence to a stable equilibrium.
+5.4 Impact of Mean-Field Homogeneity
+DSMFG leverages a mean-field approximation to ensure scalability in modeling and learning, but this introduces an inherent homogeneity assumption.
+In DSMFG, individual heterogeneity is preserved through state representations, allowing the shared policy to generate personalized actions.
+This section investigates the impact of shared follower policies by examining: (i) the effects of shared policies on training scalability and efficiency; and (ii) the robustness of DSMFG policies when interacting with heterogeneous follower behaviors.
+Scalability and efficiency of shared policy
+The shared-policy design in DSMFG significantly improves both scalability and training efficiency compared to heterogeneous-policy variants, such as DSMFG w/o MF and DSMFG w/o MF & S.
+(i) Scalability: As shown in Table 1, heterogeneous-policy variants suffer from severe training instability, achieving sustainability for only $1$ year and producing lower GDP ($1.48\times 10^{5}$ vs.
+$1.10\times 10^{13}$ for DSMFG).
+Additionally, inequality worsens with a higher wealth Gini (0.93 vs.
+0.53 at $N=1000$).
+(ii) Efficiency: Table 3 (Appendix E.1) shows that DSMFG reduces training time by 30% to reach equivalent reward levels.
+These results highlight the advantages of shared-policy training in large-scale environments, enabling efficient convergence and better economic performance.
+Robustness of DSMFG policies
+We test DSMFG policy robustness by introducing heterogeneous followers (named Non-DSMFG followers) using behavior-cloned policies derived from real-world data (Appendix E.5).
+The leader and a subset of followers retain DSMFG-trained policies, while the rest adopt Non-DSMFG followers.
+We vary the proportion of DSMFG followers (0%, 25%, 50%, 75%, 100%) and track both micro-level (wealth, income, utility) and macro-level (GDP) indicators.
+Figure 4 reports average values (left Y-axis, bars) and total values (right Y-axis, points).
+Results show three key findings: (1) DSMFG followers consistently maintain high utility (96–97) across all proportions, indicating robustness against policy heterogeneity.
+(2) DSMFG followers outperform Non-DSMFG followers in wealth, income, and utility—often by more than ×$\times 2$—demonstrating the superior effectiveness of DSMFG policies.
+(3) Per capita GDP increases monotonically with the proportion of DSMFG followers, suggesting that widespread adoption of DSMFG policies can yield substantial macroeconomic gains.
 6 Conclusion
-Addressing issues in macroeconomic policymaking, this paper innovatively proposes the SMFG method, which includes dynamic SMFG for modeling and SMFRL algorithm for solving.
-The SMFG method optimizes government’s policy by considering the dynamic responses of large-scale households to policy shifts, ultimately deriving optimal policies for both government and households.
-Experimental results demonstrate the SMFG method’s superior performance, stability, and convergence over existing approaches.
-When extended to real-world scenarios, even if some households do not adopt the SMFG policy, their wealth and utility are not as high as those who do, encouraging more households to adopt the SMFG policy.
-In conclusion, this paper contributes an effective approach to modeling and solving optimal macroeconomic policies in the fields of AI for economics and AI for social impact.
+We introduce the Dynamic Stackelberg Mean Field Game (DSMFG) framework, which captures the dynamic, asymmetric, and large-scale nature of government–individual interactions in macroeconomic settings.
+To solve DSMFG, we develop the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm, which combines Stackelberg game theory with mean-field approximation to enable scalable and efficient policy learning in large populations.
+This approach provides a principled and scalable solution to policy optimization problems that are otherwise computationally intractable.
+Our results underscore the value of integrating game-theoretic modeling with data-driven learning for large-scale economic decision-making.
+Future directions include extending DSMFG to multi-policy settings, modeling richer behavioral heterogeneity, and calibrating with real-world economic data.
 Appendix A SMFRL algorithm pseudocode
@@
 Limitations
-The limitations of our SMFG method will be thoroughly investigated in future work: (1) We plan to consider dynamic games involving multiple leaders and large-scale followers to explore policy coordination across various macroeconomic sectors.
+The limitations of our DSMFG method will be thoroughly investigated in future work: (1) We plan to consider dynamic games involving multiple leaders and large-scale followers to explore policy coordination across various macroeconomic sectors.
 (2) We will continue to develop theoretical proofs for the equilibrium solutions in Stackelberg mean field games.
@@
 The Saex tax policy is often considered a suggestion for specific tax reforms in the real world.
-The specific calculation method is as follows (Saez 2001).
+The specific calculation method is as follows [50].
 The Saez tax utilizes income distribution $f(z)$ and cumulative distribution $F(z)$ to get the tax rates.
@@
 $\mathcal{E}^{f}(\pi^{f};\pi^{l})=J^{f}(\pi^{l},{\pi^{f}}^{*},L)-J^{f}(\pi^{l},\pi^{f},L),$
-where $J^{f}$ denotes the cumulative reward for the follower, defined in Section 3.
+where $J^{f}$ denotes the cumulative reward for the follower, defined in Section 3.2.
 Similarly, the leader’s exploitability $\mathcal{E}^{l}(\pi^{l};\pi^{f})$ measures the payoff difference between the leader’s policy $\pi^{l}$ and its best response ${\pi^{l}}^{*}$, given the followers’ response policy $\pi^{f}$ and state-action distribution $L$.
@@
 $\mathcal{E}^{l}(\pi^{l};\pi^{f})=J^{l}(\pi^{l*},\pi^{f},L)-J^{l}(\pi^{l},\pi^{f},L),$
-with $J^{l}$ representing the cumulative reward for the leader, and $L=\{L_{t}\}_{t=0}^{T}$ detailing the state-action distribution for followers over time (see Section 3).
+with $J^{l}$ representing the cumulative reward for the leader, and $L=\{L_{t}\}_{t=0}^{T}$ detailing the state-action distribution for followers over time (see Section 3.2).
 The overall exploitability, which measures the discrepancy from Nash equilibrium for both the leader and the followers, is defined as:
@@
 Social Optimum and Social Welfare
-In economic theory, the Social Optimum describes a state in which the allocation of resources achieves maximum efficiency, as measured by social welfare (Arrow and Debreu 1954; Malul and Bar-El 2009).
+In economic theory, the Social Optimum describes a state in which the allocation of resources achieves maximum efficiency, as measured by social welfare [2, 38].
 Given the leader’s policy $\pi^{l}$ and the representative follower’s policy $\pi^{f}$ among large-scale followers, social welfare $\mathcal{SW}(\pi^{l},\pi^{f})$ is approximately calculated as the sum of the utility functions defined in Section 3 of the $N$ followers:
@@
 Appendix E Additional Results
-Compute Resources
+E.1 Compute Resources
 All experiments are run on 2 workstations: A 64-bit server with dual AMD EPYC 7742 64-Core Processors @2.25 GHz, 256 cores, 512 threads, 503GB RAM, and 2 NVIDIA A100-PCIE-40GB GPU.
@@
 The following Table 3 shows the approximate training times for several algorithms.
-Algorithm N=100 (hours) N=1000 (hours)
-SMFG 4 14
-SMFG-MF 3 16
-SMFG-S 4 9
-SMFG-S-MF 2 6
-Free Market 0.25 2
-Saez Tax 4 23
-AI Economist 6.5 N/A
-Table 3: The approximate training times for baselines in our experiments.
-Experimental Results of Ablation Studies
-In this section, we present additional experimental results from ablation studies, including training curves 6 and a test table 4, as well as experiments incorporating the use of behavior cloning as a pre-training strategy for follower agents.
-We find that the SMFG method without behavior cloning as pre-training still surpasses other baselines that utilize behavior cloning.
-More specifically, we compared SMFG with 5 baselines across 4 different experimental setups: without behavior cloning as pre-training for follower agents at N=100 and N=1000 (marked as N=100 without BC and N=1000 without BC); with BC-based pre-training for follower agents at N=100 and N=1000 (N=100-BC; N=1000-BC).
+Algorithm Training Time (hours) Utility (years) Utility per training time
+N=100 N=1000 N=100 N=1000 N=100 N=1000
+DSMFG 4 14 300 300 75.00 21.43
+DSMFG w/o MF 3.5 16 1 1.02 0.29 0.06
+DSMFG w/o S 4 9 75.75 1.5 18.94 0.17
+DSMFG w/o MF & S 2 6 1 1 0.50 0.17
+Free Market 0.25 2 1 1 4.00 0.50
+Saez Tax 4 23 300 100.58 75.00 4.37
+AI Economist 6.5 N/A 1 N/A 0.15 N/A
+Table 3: The average training times, utility (Years), and utility per training time for baselines in our experiments.
+The best values are highlighted in bold, and the second-best values are underlined.
+Utility is measured using the "Years" metric, which represents the number of simulation steps achievable under a given policy.
+A higher number of simulation steps indicates better policy performance but also corresponds to increased computational complexity.
+E.2 Further Experiments on the Necessity of SMFG
+In this section, we present additional experimental results for validating the necessity of SMFG, including training curves Figure 6 and Table 4 and 5, as well as experiments incorporating the use of behavior cloning as a pre-training strategy for follower agents.
+We find that the DSMFG method without behavior cloning as pre-training still surpasses other baselines that utilize behavior cloning.
+More specifically, we compared DSMFG with 5 baselines across 4 different experimental setups: without behavior cloning as pre-training for follower agents at N=100 and N=1000 (marked as N=100 without BC and N=1000 without BC); with BC-based pre-training for follower agents at N=100 and N=1000 (N=100-BC; N=1000-BC).
 Figure 6 illustrates the training curves of 4 key macroeconomic indicators under these four settings.
@@
 Each subplot’s Y-axis represents the indicators’ values, and the X-axis represents the training steps.
-Table 4 displays the test results of the 6 algorithms across 5 indicators, with each column corresponding to an experimental setting.
-Figure 6 and Table 4 present two experimental findings: (1) Using BC as a pre-training method for the follower’s policy enhances the algorithms’ stability and performance.
-Comparing settings with and without BC (the first two rows), our method, SMFG, shows similar convergence outcomes; however, the performance of other algorithms significantly improves across all four indicators with BC-based pre-training.
+Table 4 and 5 displays the test results of the 7 algorithms across 4 indicators, with each column corresponding to an experimental setting.
+Figure 6 and Table 4 and 5 present two experimental findings: (1) Using BC as a pre-training method for the follower’s policy enhances the algorithms’ stability and performance.
+Comparing settings with and without BC (the first two rows), our method, DSMFG, shows similar convergence outcomes; however, the performance of other algorithms significantly improves across all four indicators with BC-based pre-training.
 Furthermore, the training curves of each algorithm are more stable.
-(2) The SMFG method substantially outperforms other algorithms in solving SMFGs, both in large-scale followers and without pre-training scenarios.
-In the setting of N=100-BC, SMFG achieved a significantly higher per capita GDP compared to other algorithms, while its social welfare and Gini index are similar to others, essentially reaching the upper limit.
-Besides, in N=100 without BC and N=1000-BC, SMFG consistently obtains the most optimal solutions across all indicators.
-Settings Algorithm Per Capita GDP Social Welfare Income Gini Wealth Gini Years
-N=100 without BC Rule-based+Random 1.41e+05 69.27 0.89 0.92 1.00
-DDPG+Random 1.41e+05 70.91 0.88 0.92 1.00
-SMFG-MF 1.23e+05 48.17 0.89 0.93 1.00
-SMFG-S-MF 1.21e+05 83.09 0.88 0.92 1.00
-SMFG-S 8.66e+07 82.02 0.82 0.83 75.75
-SMFG 9.59e+12 96.87 0.52 0.51 300.00
-N=1000 without BC Rule-based+Random 1.42e+05 705.85 0.90 0.93 1.00
-DDPG+Random 1.42e+05 654.37 0.90 0.93 1.00
-SMFG-MF 1.48e+05 499.01 0.89 0.93 1.02
-SMFG-S-MF 1.33e+05 874.53 0.89 0.92 1.00
-SMFG-S 1.31e+05 834.93 0.88 0.92 1.50
-SMFG 1.10e+13 968.94 0.54 0.53 300.00
-N=100 with BC Rule-based+Real 3.66e+11 79.23 0.52 0.53 217.45
-DDPG+Real 2.03e+12 94.50 0.46 0.48 299.85
-SMFG-MF 6.38e+12 93.89 0.57 0.58 268.53
-SMFG-S-MF 7.41e+12 98.16 0.53 0.55 300.00
-SMFG-S 5.44e+12 98.21 0.50 0.52 300.00
-SMFG 1.01e+13 96.90 0.51 0.53 299.89
-N=1000 with BC Rule-based+Real 1.41e+05 334.79 0.90 0.93 1.00
-DDPG+Real 4.92e+11 527.09 0.75 0.79 100.68
-SMFG-MF 6.82e+12 954.88 0.56 0.62 278.50
-SMFG-S-MF 2.79e+12 512.19 0.77 0.81 100.68
-SMFG-S 1.13e+05 440.00 0.90 0.93 1.00
-SMFG 9.68e+12 975.15 0.52 0.51 300.00
-Table 4: Performance metrics of six algorithms across different settings on key macroeconomic indicators.
+(2) The DSMFG method substantially outperforms other algorithms in solving DSMFGs, both in large-scale followers and without pre-training scenarios.
+In the setting of N=100-BC, DSMFG achieved a significantly higher per capita GDP compared to other algorithms, while its social welfare and Gini index are similar to others, essentially reaching the upper limit.
+Besides, in N=100 without BC and N=1000-BC, DSMFG consistently obtains the most optimal solutions across all indicators.
+Table 4: Performance of multiple policies on key macroeconomic indicators for $N=100$ households. The best values are highlighted in bold, and the second-best values are underlined.
+Category Subcategory Policies Per Capita Social Income Wealth Years $\uparrow$
+GDP $\uparrow$ Welfare $\uparrow$ Gini $\downarrow$ Gini $\downarrow$
+Static Policy Non-intervention Free Market 1.37e+05 32.97 0.89 0.92 1.10
+Real-data US Federal Tax 4.88e+11 94.19 0.40 0.40 289.55
+Rule-based Saez Tax 2.34e+12 73.82 0.21 0.38 300.00
+Dynamic Policy Independent-based AI Economist 1.26e+05 72.81 0.88 0.91 1.00
+without BC Game-based Markov Game 1.21e+05 83.09 0.88 0.92 1.00
+Stackelberg Game 1.23e+05 48.17 0.89 0.93 1.00
+Mean Field Game 8.66e+07 82.02 0.82 0.83 75.75
+DSMFG (ours) 9.59e+12 96.87 0.52 0.51 300.00
+Dynamic Policy Independent-based AI Economist 2.03e+12 94.50 0.46 0.48 299.85
+with BC Game-based Markov Game 7.41e+12 98.16 0.53 0.55 300.00
+Stackelberg Game 6.38e+12 93.89 0.57 0.58 268.53
+Mean Field Game 5.44e+12 98.21 0.50 0.52 300.00
+DSMFG (ours) 1.01e+13 96.90 0.51 0.53 299.89
+Table 5: Performance of multiple policies on key macroeconomic indicators for $N=1000$ households. The best values are highlighted in bold, and the second-best values are underlined.
+Category Subcategory Policies Per Capita Social Income Wealth Years $\uparrow$
+GDP $\uparrow$ Welfare $\uparrow$ Gini $\downarrow$ Gini $\downarrow$
+Static Policy Non-intervention Free Market 1.41e+05 334.79 0.90 0.93 1.00
+Real-data US Federal Tax 1.41e+05 351.17 0.89 0.93 1.00
+Rule-based Saez Tax 6.35e+11 498.88 0.68 0.73 100.58
+Dynamic Policy Independent-based AI Economist N/A N/A N/A N/A N/A
+without BC Game-based Markov Game 1.33e+05 874.53 0.89 0.92 1.00
+Stackelberg Game 1.48e+05 499.01 0.89 0.93 1.02
+Mean Field Game 1.31e+05 834.93 0.88 0.92 1.50
+DSMFG (ours) 1.10e+13 968.94 0.54 0.53 300.00
+Dynamic Policy Independent-based AI Economist N/A N/A N/A N/A N/A
+with BC Game-based Markov Game 2.79e+12 512.19 0.77 0.81 100.68
+Stackelberg Game 6.82e+12 954.88 0.56 0.62 278.50
+Mean Field Game 1.13e+05 440.00 0.90 0.93 1.00
+DSMFG (ours) 9.68e+12 975.15 0.52 0.51 300.00
 Figure 6: The training curves for 6 algorithms on 4 macroeconomic indicators, comparing settings without behavior cloning as pre-train (N=100 & N=1000) and with behavior cloning (N=100-BC & N=1000-BC).
-Training Curves for Various Tax Policies
+E.3 Training Curves for Various Tax Policies
 We compare the performance of 6 policies across four economic indicators under two settings: with N=100 and N=1000 households.
-Figure 7 displays the training curves and Table 1 shows the test results.
-Both Figure 7 and Table 1 indicate that the SMFG method significantly surpasses other policies in the task of optimizing GDP, and achieves the highest social welfare.
+Figure 7 displays the training curves and Table 4 and 5 shows the test results.
+Both Figure 7 and Table 4 and 5 indicate that the DSMFG method significantly surpasses other policies in the task of optimizing GDP, and achieves the highest social welfare.
 When N=100, the Saez tax achieves the lowest income and wealth Gini coefficients, suggesting greater fairness.
-However, at N=1000, SMFG performs optimally across all economic indicators, while the effectiveness of other policies noticeably diminishes as the number of households increases.
-The Saez tax also reduces the Gini index, but not as effectively or stably as the SMFG.
+However, at N=1000, DSMFG performs optimally across all economic indicators, while the effectiveness of other policies noticeably diminishes as the number of households increases.
+The Saez tax also reduces the Gini index, but not as effectively or stably as the DSMFG.
 Figure 7: The training curves for 6 tax policies on 5 macroeconomic indicators (N=100 & N=1000 with BC)
-Efficiency-Equity Tradeoff of Policies
+E.4 Efficiency-Equity Tradeoff of Policies
 Figure 8: Comparative performance of various policies under multi-objective assessment (Efficiency-Equity).
 In economics, the Efficiency-Equity Tradeoff is a highly debated issue.
-We find that our SMFG method is optimal in balancing efficiency-equity, except in cases of extreme concern for social fairness.
+We find that our DSMFG method is optimal in balancing efficiency-equity, except in cases of extreme concern for social fairness.
 In our study, we depict the economic efficiency (Per capita GDP) on the Y-axis and equity (wealth Gini) on the X-axis of Figure 8(a) for various policies.
@@
 Different circles of the same color correspond to different seeds.
-Figure 8 (a) shows that the wealth Gini indices for SMFG and AI Economist-BC are similar, but SMFG has a higher GDP, suggesting its superiority over AI Economist-BC.
-SMFG significantly outperforms the free market policy and AI Economist due to its higher GDPs and lower wealth Ginis.
-However, comparing SMFG with the Saez tax and the U.S.
+Figure 8 (a) shows that the wealth Gini indices for DSMFG and AI Economist-BC are similar, but DSMFG has a higher GDP, suggesting its superiority over AI Economist-BC.
+DSMFG significantly outperforms the free market policy and AI Economist due to its higher GDPs and lower wealth Ginis.
+However, comparing DSMFG with the Saez tax and the U.S.
 Federal tax policy in terms of both economic efficiency (GDP) and social equity (Gini) is challenging.
@@
 Due to the logarithmic treatment of GDP in (b), when $\alpha=10$, the overall objective focuses solely on social fairness; when $\alpha=0$, the overall objective is concerned only with efficiency.
-Our findings in (b) reveal that only when $\alpha\geq 8$, which indicates a substantial emphasis on social equity, does the Saez tax outperform SMFG.
-However, SMFG consistently proves to be the most effective policy under a wide range of preference settings.
-Behavior Cloning Experiments
-We conduct behavior cloning based on real data to simulate the behavior strategies of households in realistic scenarios, which are then used in Experiment 5 to compare with SMFG Households.
+Our findings in (b) reveal that only when $\alpha\geq 8$, which indicates a substantial emphasis on social equity, does the Saez tax outperform DSMFG.
+However, DSMFG consistently proves to be the most effective policy under a wide range of preference settings.
+E.5 Behavior Cloning Experiments
+We conduct behavior cloning based on real data to simulate the behavior strategies of households in realistic scenarios, which are then used in Experiment 5.4 to compare with DSMFG followers.
 We collect the statistical data from the 2022 Survey of Consumer Finances (SCF) (https://www.federalreserve.gov/econres/scfindex.htm) as the real data buffer $\mathcal{D}_{real}$.
@@
 $\min_{\theta}\mathcal{L}_{MSE}$ $=\mathbb{E}_{s^{f},a^{f}\sim\mathcal{D}}\left(a^{f}-a\right)^{2}|_{a=\pi_{\theta}(s^{f})}.$
-This experiment conducts behavior cloning on networks for four different household policies: Multilayer Perceptron (MLP), AI economist’s network (MLP+LSTM+MLP), SMFG-S, and SMFG-MF network.
+This experiment conducts behavior cloning on networks for four different household policies: Multilayer Perceptron (MLP), AI economist’s network (MLP+LSTM+MLP), DSMFG w/o S, and DSMFG w/o MF network.
 The first two, as their network outputs, are probability distributions, use negative log-likelihood loss (Figure 9 left); the latter two’s networks employ deterministic policies, hence they use mean square error loss against real data (Figure 9 right).
@@
 It can be observed that the AI economist’s network, due to its complexity, struggles to converge to near -1 like MLP.
-The losses corresponding to MFRL and SMFG-MF can converge to below 0.1.
+The losses corresponding to MFRL and DSMFG w/o MF can converge to below 0.1.
 Figure 9: The behavior cloning loss for 4 networks in two loss types.
-Appendix F Economic Model Details
+Appendix F TaxAI
+F.1 Introduction of TaxAI
+TaxAI is a novel Multi-Agent Reinforcement Learning (MARL) environment designed to model dynamic interactions among governments, households, firms, and financial intermediaries.
+Built on the Bewley-Aiyagari economic model, TaxAI addresses critical limitations of existing economic simulators by offering enhanced scalability, realism, and benchmarking capabilities.
+  • Scalability: TaxAI simulates dynamic interactions involving up to 10,000 households, significantly surpassing the scale of prior simulators and enabling large-scale analysis.
+  • Realism: Calibrated with real-world data from the 2022 Survey of Consumer Finances (SCF), TaxAI ensures its simulations reflect realistic economic scenarios, improving the relevance of its outcomes for policymaking.
+  • Benchmarking: TaxAI evaluates 7 MARL algorithms against 2 traditional economic approaches (e.g., genetic algorithms, dynamic programming), demonstrating the superiority of MARL in addressing dynamic, partially observable economic environments.
+  • Policy Optimization: TaxAI leverages MARL’s adaptive learning capabilities to model complex government-household interactions and discover optimal tax policies that promote growth and equity.
+With its ability to integrate scalability, real-world calibration, and MARL-based adaptive optimization, TaxAI sets a new benchmark for realistic and effective economic simulators, providing actionable insights for policy design and implementation.
+Therefore, TaxAI is highly suitable as the experimental environment for this paper, particularly due to its scalability and realism.
+F.2 Economic model details
 Economic activities among households aggregate into labor markets, capital markets, goods markets, etc.
@@
 $Y_{t}=C_{t}+G_{t}+X_{t}$
-where $C_{t}=\sum_{i}^{N}c^{i}_{t}$ represents the total consumption of consumers, and $G_{t}$ is government spending. The supply, demand, and price represent the states of the market.
-Economic Shocks
+where $C_{t}=\sum_{i}^{N}c^{i}_{t}$ represents the total consumption of households, and $G_{t}$ is government spending. The supply, demand, and price represent the states of the market.
+F.3 Economic Shocks
 In Experiment 5, we simulate economic shocks analogous to a financial crisis: at the 100-th step, the wealth of all households is reduced by 50%.
@@
 Learning rate adjustment $0.95^{(\text{epoch}//35)}$
-Table 5: Hyperparameters of SMFG methods and its variants.
+Table 6: Hyperparameters of DSMFG methods and its variants.
 Hyperparameter Value
@@
 Epsilon decay 1e-5
-Table 6: Hyperparameters of SMFG-MF algorithm different from SMFG method.
+Table 7: Hyperparameters of DSMFG w/o MF algorithm different from DSMFG method.
 Hyperparameter Value
@@
 Learning rate adjustment $0.97^{(\text{epoch}//35)}$
-Table 7: Hyperparameters of AI Economist Algorithm different from SMFG approach.
+Table 8: Hyperparameters of AI Economist Algorithm different from DSMFG approach.
 Ethical Statement
-This research introduces a novel SMFG method, designed to optimize macroeconomic policies by modeling complex interactions at the micro level.
+This research introduces a novel DSMFG method, designed to optimize macroeconomic policies by modeling complex interactions at the micro level.
 The potential impact of this work extends across several domains:
@@
 Continuous validation against empirical data and diverse economic scenarios is essential to ensure the reliability and ethical application of the proposed methods.
-In summary, the proposed SMFG framework and SMFRL algorithm hold the potential to significantly impact both academic research and practical policy making, offering a new perspective on dynamic economic modeling that prioritizes realistic, individual-level responses within large-scale economic systems.
+In summary, the proposed DSMFG framework and SMFRL algorithm hold the potential to significantly impact both academic research and practical policy making, offering a new perspective on dynamic economic modeling that prioritizes realistic, individual-level responses within large-scale economic systems.